### v1

# PlanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

Abstract
We introduce PlanQA, a diagnostic benchmark for evaluating geometric and spatial reasoning in large-language models (LLMs). PlanQA is grounded in structured representations of indoor scenes, such as kitchens, living rooms, and bedrooms, encoded in a symbolic format (e.g., JSON, XML layouts). The benchmark includes diverse question types that test not only metric and topological reasoning (e.g., distance, visibility, shortest paths) but also interior design constraints such as affordance, clearance, balance, and usability. Our results across a variety of frontier open-source and commercial LLMs show that while models may succeed in shallow queries, they often fail to simulate physical constraints, preserve spatial coherence, or generalize under layout perturbation. PlanQA uncovers a clear blind spot in today’s LLMs: they don’t consistently reason about real-world layouts. We hope that this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
1 Introduction
Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over physically grounded environments remains poorly understood. In numerous real-world applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images or natural language. Reasoning in these contexts requires geometric inference over symbolic layouts, not pixel-level perception.
We introduce PlanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floor plans represented in structured text-based formats. Each instance consists of a JSON-encoded layout paired with natural language questions that require the model to compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints. PlanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments.
Although LLMs can increasingly be used in tool-assisted pipelines, for example to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities. PlanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding.
Specifically, our contributions are as follows:
  • We introduce a dataset of 1,800 structured 2D layouts, 600 each from kitchens, living rooms, and bedrooms, represented in JSON and paired with spatial reasoning questions.
  • We provide a diverse suite of 18,000 questions spanning geometric, functional, and design-aware reasoning, with one question per layout across multiple categories.
  • We establish structured evaluation protocols and scoring metrics that enable a fine-grained diagnosis of reasoning performance by task type and error mode.
  • We conduct a comparative analysis of 11 LLMs, including GPT-4.1 OpenAI et al. (2024) and open models such as DeepSeek-R1 DeepSeek-AI (2025) and Llama-3.3-70B-Instruct Grattafiori et al. (2024), revealing consistent failure patterns in spatial inference from symbolic input.
PlanQA provides a diagnostic substrate for measuring spatial competence in language models, with applications in layout validation, agent planning, and symbolic grounding.
2 Related Work
Prior benchmarks have explored spatial reasoning across vision and language domains. CLEVR Johnson et al. (2017) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relationships. In real-world settings, SpatialSense Yang et al. (2019) focuses on recognizing spatial relations in images through adversarially mined examples. Benchmarks like BabyAI Chevalier-Boisvert et al. (2019), ALFRED Shridhar et al. (2020), and Room-to-Room (R2R) Anderson et al. (2018) integrate spatial understanding into embodied tasks, requiring agents to follow instructions involving navigation and object manipulation in simulated environments. Recent datasets such as ScanQA Azuma et al. (2022) and 3DSRBench Ma et al. (2024a) extend spatial reasoning evaluation into 3D environments, emphasizing the need for models to comprehend and reason about spatial relationships in three dimensions.
Vision-language models have advanced spatial reasoning but often handle it qualitatively. The VQA dataset Antol et al. (2015) challenges models to answer questions about images, while VL-T5 Cho et al. (2021) unifies vision-and-language tasks via text generation. Recent work on 3D scene graphs Armeni et al. (2019) introduces structured representations of environments, facilitating spatial reasoning. However, these approaches may miss fine-grained geometric details necessary for precise spatial inference. Efforts like SpatialVLM Chen et al. (2024) aim to endow vision-language models with enhanced spatial reasoning capabilities, addressing some of these limitations.
Advancements in generative models have also contributed to spatial reasoning tasks. LayoutGPT Feng et al. (2023) leverages large language models for compositional visual planning and layout generation, while Holodeck Yang et al. (2024) enables language-guided generation of 3D embodied AI environments. Similarly, AnyHome Fu et al. (2024) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts. Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but suffers from poor object placement, often yielding physically implausible or random layouts due to non-convergent simulated annealing. LayoutVLM Sun et al. (2024) and FirePlace Huang et al. (2025) improve layout generation via optimization and constraint solving, respectively. But they assess spatial reasoning through output realism, not direct inference. In contrast, our benchmark tests symbolic reasoning without tool-assisted refinement.
Evaluations of large language models’ spatial understanding have been conducted in studies like Evaluating Spatial Understanding of Large Language Models Yamada et al. (2024), which assesses the spatial reasoning capabilities of LLMs through structured tasks. Additionally, benchmarks such as BALROG Paglieri et al. (2025) test agentic reasoning in game environments, further exploring the spatial and decision-making abilities of language and vision-language models. While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning in structured layouts, providing fine-grained and task-specific insights into models’ spatial competence. Recent 3D-LLM surveys such as Ma et al. (2024b) cover tasks like navigation and interaction, but not symbolic spatial reasoning. PlanQA fills this gap by testing raw spatial competence from structured layouts without multimodal input.
PlanQA addresses the gap in existing benchmarks by directly evaluating structured spatial inference from symbolic room layouts. Unlike prior benchmarks relying on raw images or focusing on commonsense spatial language, PlanQA provides explicit spatial representations (object coordinates and dimensions) and tests models’ abilities to perform precise spatial reasoning tasks, such as calculating distances, assessing visibility, and verifying object fit within a controlled setting.
3 Method
3.1 Synthetic Layout Generation
Our initial aim was to utilize publicly available real-world floorplan datasets. However, a comprehensive review revealed critical limitations. Prominent datasets such as SUNCG (Song et al., 2017) and HouseExpo (Li et al., 2019) were withdrawn due to unresolved copyright claims, rendering them unusable. Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2020), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse. Datasets like CubiCasa5K (Kalervo et al., 2019) and Rent3D (Liu et al., 2015) offer fixed architectural plans from real environments but lack furnishing annotations. RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al. (2020), while large, appears procedurally generated albeit with realtor supervision and imposes restrictions on redistribution and derived works. Given these legal, practical, and methodological constraints, we found synthetic data generation to be the most viable alternative.
We generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro, a large language model fine-tuned for spatial reasoning (Team et al., 2025). While employing an LLM for both data generation and evaluation may raise concerns of circularity, our evaluation pipeline is entirely independent: the LLM is used solely for data generation, and evaluation is performed via deterministic spatial reasoning tasks.
The generation process comprises two stages. First, we specify room geometries using explicit constraints on shape, adjacency, and design principles related to circulation, symmetry, and zoning. These constraints are encoded directly in the LLM prompt. Second, each room is furnished according to style-specific rules, also defined in structured prompts, to encourage both visual realism and functional plausibility. Approximately one-third of candidate layouts are filtered out by automated spatial validity checks, which remove configurations with inaccessible furniture or implausible adjacencies. This curation pipeline yields a diverse and geometrically coherent dataset. Full prompts, generation templates, and validation scripts are provided in the supplemental materials. Representative examples of layout types are shown in Fig. 1. A breakdown of room types and layout geometries is provided in Table 1.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Representative layouts from the PlanQA dataset, spanning three room types (kitchen, living room, bedroom) and three geometric configurations (rectangular, L-shaped, open). Layouts are procedurally generated using explicit spatial constraints and validated for functional feasibility.
Table 1: Layout distribution in PlanQA by room type, style, and geometry, with room-wise totals.
Room Type Style Rectangular L-Shaped Open Total
Kitchen L-Shaped (incl. counter) 208 36 105 600
U-Shaped / G-Shaped 42 128 2
Island-Based 2 0 33
Wall / Galley (compact layouts) 17 10 17
Living Room Fireplace-Centric 55 1 34 600
Conversational 66 3 37
Multi-Zone 73 176 46
TV-Focused 78 3 28
Bedroom Bedroom with Workstation 98 35 28 600
Traditional Bedroom 114 59 57
Efficient Small Bedroom 65 3 8
Welcoming Guest Bedroom 75 30 28
Total 893 484 423 1800
3.2 Question Taxonomy, Prompting, and Cognitive Scope
PlanQA assesses spatial reasoning by presenting models with a single natural language question per symbolic layout. Questions span a range of geometric, topological, and functional types, including numeric computations (distances, areas), spatial feasibility (object fit, placement), occlusion and visibility, and environment-level assessments such as missing elements or requirement violations. Examples include queries like whether a 2 m $\times$ 3 m table can be placed without blocking passage, whether the television is visible from a particular seat, or which essential appliances are absent from a proposed kitchen. Some questions require local metric reasoning; others test whether a model can simulate constraints that would arise in physical occupancy or usability. A categorized list of question types is shown in Table 2.
Each question is generated by instantiating a parameterized template against the specific geometry of a layout. These templates are natural language prompts that fill in object names, measurements, and task-specific context. Prompts are issued in zero-shot settings with no few-shot scaffolding or role-based instruction, but include consistent structural cues to support stepwise inference. Each prompt ends with the string ”Reason step by step to solve the problem.” followed by “Final Answer:”, enabling unambiguous output extraction. For instance, a placement query might read: “Given the room layout below, compute the area of the largest rectangle that can be placed in the kitchen without overlapping any object. Reason step by step to solve the problem. Final Answer: [value in square meters].”
Table 2: PlanQA question taxonomy. The example question shown is an instantiation of the template used to generate all questions of that type. Each task is labeled with a format code, cognitive skill category (see Tables 4 and 4), and the room types in which it is applied: K = kitchen, L = living room, B = bedroom.
Category Example Question Format Target Rooms
Distance What is the distance between the refrigerator and the stove? N M K, L, B
Area What is the total sitting area? N M K, L, B
Free Space What is the unoccupied floor area in the room? N T K, L, B
View Angle What is the angle between the TV’s facing direction and the viewer’s line of sight from the armchair? N V K, L, B
Repositioning How far can the ottoman be moved to the left direction until it touches another object or the wall? N G K, L, B
Max Box What is the area of the largest axis-aligned rectangle that fits in the free space? N G K, L, B
Fit/Placement Can a 2m × 3m desk table fit in the room without overlap? B G K, L, B
Shortest Path What is the valid shortest path from the stove to the exit with clearance? S G, R K, L, B
Missing Object What is a possible placement for the missing sink? S R, T K, L, B
Obstruction What are all the objects that intersect the straight line from the fireplace to the door? L V K, L, B
Table 3: Answer format and scoring codes.
Code Format Scoring
N Scalar Rel. error $\leq$ 2%
B Boolean Exact match
S Sequence Fréchet $\leq$ 0.6 m
L List Set equality
C Coord. Spatial predicate
Table 4: Cognitive skill clusters.
Code Cognitive Target
M Metric computation
T Topology
G Geometric planning
V Visual logic / occlusion
R Rule-based validation
Structured layouts are flattened into lists of labeled objects, each defined by an axis-aligned bounding box. Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents are unique and stable across prompt construction and answer evaluation. All coordinates are expressed in meters, using a bottom-left origin and $(x,y)$ ordering. There is no hierarchical or typed schema beyond the flat object list, and no implicit conventions are assumed. A single system prompt is prepended to every interaction, specifying general and room-specific constraints: “You are a spatial assistant. Assume that all windows have zero thickness.” For layouts of type living room or bedroom, the prompt further clarifies that rugs may be overlapped by other objects, allowing for more realistic spatial arrangements involving soft surfaces.
Some question types are instantiated deterministically given the layout geometry (e.g., exact distances between centroids), while others are drawn from constrained samplers over feasible object pairs, free-space regions, or task-specific heuristics. The prompt templates themselves remain fixed across layouts and models. This uniformity allows us to isolate spatial reasoning under consistent linguistic and structural conditions.
These cognitive categories correspond to distinct reasoning faculties. Metric tasks (M) require explicit numerical computations over distances, areas, or positions, often grounded in real-valued geometric attributes. Topological tasks (T) depend on spatial relationships like adjacency or containment that are invariant under metric transformations. Geometric planning (G) involves multistep spatial inference over object configurations, including feasibility, collision avoidance, or spatial optimization (e.g., placing a large box without overlap). Visual logic (V) tasks such as visibility and angle estimation rely on implicit occlusion reasoning or directional projection. Rule-based tasks (R) involve symbolic constraints, such as verifying whether a layout satisfies clearance or adjacency requirements or whether essential objects are missing. These clusters are intended to reflect the diversity of tasks while enabling diagnostic grouping of error types. They also reflect a spectrum of abstraction, fromm low-level metric properties to high-level design affordances. While these clusters are not mutually exclusive, they provide a useful abstraction for analyzing model behavior and diagnosing failures. For example, we hypothesize that questions involving geometric planning and occlusion (G, V) will be disproportionately challenging for models without grounded spatial priors. This hypothesis is revisited in our results analysis (§4). The formatrmat and target codes are defined in Tables 4 and 4, and used throughout Table 2.
3.3 Evaluation Protocol and Scoring
Each question in PlanQA is paired with a reference answer derived entirely from the symbolic layout, allowing deterministic and automated evaluation of model outputs. Depending on the response type, correctness is assessed using tolerance-based numeric comparison, symbolic string matching, or geometric validation routines. The scoring logic referenced by each question format code is detailed in Table 4, and is applied uniformly across models.
For numerical questions (e.g., distances, areas, angles), predictions are accepted if they fall within a fixed relative error of the reference: $\left|\frac{\hat{y}-y^{*}}{y^{*}}\right|\leq\varepsilon$ where $\varepsilon$ is typically 0.02. This threshold is applied uniformly across models, and reflects spatial resolution limits in architectural layouts.
Binary and categorical responses, such as visibility, object fit, or violation detection, are validated against geometry-aware checks. For example, visibility is determined by ray intersection tests over the layout’s free space, while object fit questions are scored by verifying non-overlap and clearance constraints for a candidate placement. Boolean answers must match the reference exactly.
Structured outputs, including bounding boxes and paths, are checked for both format and semantic correctness. Rather than using exact coordinates, we validate these using geometric plausibility conditions—e.g., sufficient overlap with ground truth, or collision-freeness for planned paths. Thresholds for spatial alignment (e.g., bounding box proximity or path deviation) are set conservatively to account for minor lexical variation without over-penalizing near-correct reasoning.
To differentiate reasoning failures from generation or parsing errors, we annotate each model response with a status flag: Correct, Incorrect, Malformed, or Missing. Malformed responses include answers lacking the required structure “Final Answer:” or acceptable variations thereof, or containing ill-typed data. These annotations allow us to separate breakdowns in spatial inference from extraction failures and language model formatting idiosyncrasies.
3.4 Model Inference Setup
All models are queried using standard chat-based completion APIs, with prompts constructed as described above. We evaluate a mixture of open-weight and proprietary language models across three resource tiers, distinguished by their token limits and relative inference cost. Models in the “reasoning” tier are allocated up to 8192 tokens per completion, enabling them to process longer layout descriptions and multi-step CoT outputs. Other models operate within a 4096-token context window and are evaluated only on the non-reasoning subset of questions.
No model receives fine-tuning or prompt adaptation specific to PlanQA. All prompts are zero-shot, with the system message and output formatting constraints held fixed. For each model, we use the default inference configuration provided by its vendor, including temperature settings. In practice, these are near-deterministic across runs and sufficiently stable for head-to-head comparison. In cases where a model produces variable-length completions or truncates early, we retain the first parseable instance of the “Final Answer:” line, if present.
Each model is evaluated over a batch of 600 layouts per room type, with one question per layout. Requests are grouped into new-line-delimited JSON (JSONL) payloads for efficient submission via batched APIs. Layout serialization, prompt insertion, and output parsing are fully automated, with no manual curation or intervention.
We include both high-capacity generalist models and compact instruction-tuned variants to measure the degree to which spatial reasoning emerges with scale and supervision alone. All models are evaluated on identical input distributions and scoring criteria, enabling cross-system comparisons that are both cost-aware and architecture-agnostic.
4 Results
We evaluated the performance of the model on a dataset of 1,800 layouts, consisting of 600 kitchens, 600 living rooms and 600 bedrooms. Each layout is paired with one question sampled from a pool of 10 parameterized templates, filtered by room applicability. This yields 18,000 layout–question pairs. The full taxonomy is shown in Table 2.
Eleven language models are evaluated, covering a wide range of parameter scales, architectures, and training regimes. The reasoning-oriented models include Qwen3-32B Yang et al. (2025) and DeepSeek-R1 DeepSeek-AI (2025). The large general-purpose models include DeepSeek-V3 DeepSeek-AI (2024), LLaMA 3.3-70B-Instruct, Gemma 2-27B-it Riviere et al. (2024), phi-4 Abdin et al. (2024b), and GPT-4.1 OpenAI et al. (2024). The smaller models include LLaMA 3.1-8B-Instruct Grattafiori et al. (2024), Gemma 2-9B-it Riviere et al. (2024), Phi-3.5-mini Abdin et al. (2024a), and GPT-4o-mini. All models are evaluated in identical zero-shot conditions using standardized prompts and serialized layout inputs, as described in Section 3. The complete results disaggregated by question, room type, and model are included in the appendix.
4.1 Quantitative Results
We begin by aggregating accuracy across cognitive targets as defined in Table 4. Figure 2 shows a heatmap of mean accuracy by model and target category. Metric reasoning (M) and visual logic (V) show the highest cross-model scores, while Topology (T), rule-based validation (R), and geometric planning (G) are more difficult across the board. DeepSeek-R1 (reasoning), GPT-4.1 (large), and GPT-4o-mini (compact, tuned) consistently lead the performance in the harder categories, although even large models struggle in certain occlusion and clearance tasks. Surprisingly, GPT-4.1’s performance on vector-direction (V) questions nearly matches DeepSeek-R1 and GPT-4o-mini—despite those models’ explicit chain-of-thought optimization—highlighting an unexpected parity in occlusion reasoning.
Figure 2: Accuracy by model and cognitive target. Each cell shows average accuracy across questions associated with a given skill cluster. Cognitive targets: M = Metric computation, T = Topology, G = Geometric planning, V = Visual logic / occlusion, R = Rule-based validation.
Performance improves with model scale, but not uniformly across tasks. As shown in Figure 3, accuracy improves steeply up to  30B parameters, suggesting that scale alone cannot fully close the gap on complex spatial tasks. Notably, gains largely plateau beyond 30 B parameters, indicating diminishing returns for raw scale on planning and constraint problems.
Figure 3: Accuracy vs. model size. Models are arranged from the smallest to largest in terms of trainable parameters. While larger models perform better overall, instruction tuning and task specialization also contribute significantly.
We also compare accuracy across room types (Figure 4). Kitchens exhibit the highest overall accuracy, likely because their regular counter-and-appliance grids reduce occlusion and simplify clearance reasoning. In contrast, living rooms and bedrooms generally yield lower performance across models, due to denser layouts, complex furniture adjacencies, and higher occlusion likelihood, with frequent object groupings such as televisions on stands, sofas and coffee tables on rugs, or lamps on nightstands. These room-type gaps suggest that layouts with strong grid structure provide easier inductive priors for LLMs than organic, cluttered arrangements.
Figure 4: Accuracy by question and room type. Kitchens show the highest average accuracy, likely due to their regular, grid-like counter layouts and lower occlusion. The inability of most models to estimate free space in living rooms and bedrooms may be caused by more clutter, or the presence of objects like area rugs which the model was instructed to ignores for occlusion purposes.
4.2 Ablation
To evaluate the robustness of layout interpretation under alternate encodings, we perform a format ablation that replaces the standard JSON layout with a semantically equivalent XML version. This substitution preserves all geometric and object-level content while modifying only the syntax and structural serialization. To reduce evaluation cost, we focused on the most variance-sensitive question categories in the kitchen domain: Repositioning, Vector Direction, and Free Space. These were selected based on standard deviation in accuracy across the full model set. We apply each ablation to three illustrative models spanning architecture classes and capacity tiers: GPT-4.1 (closed large), GPT-4o-mini (compact tuned), and DeepSeek-R1 (reasoning-optimized). As shown in Table 5, accuracy remains stable across formats for most question types, suggesting that these models encode layout semantics in a manner invariant to low-level representation details.
Table 5: Accuracy using JSON vs XML layout encoding. Each cell shows performance on the original JSON representation and its equivalent XML rendering.
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 76.5 92.2 → 93.7 75.5 → 89.6
GPT-4o-mini 33.8 → 29.6 42.4 → 42.2 12.2 → 15.8
DeepSeek-R1 83.6 → 93.7 74.7 → 74.3 95.7 → 96.1
5 Conclusion
We have presented PlanQA, a benchmark that evaluates spatial reasoning over symbolic 2D layouts by leveraging object-centric encodings (JSON, XML) similar to those used in architecture and robotics. Through a systematic evaluation of 11 language models across five skill clusters: metric computation, topology, geometric planning, visual logic, and rule-based validation, we make critical findings with implications for the use of LLMs in geometric design tasks.
Metric and basic visibility queries (distances, areas, line-of-sight) are handled reliably (>95% accuracy), whereas geometric planning, topological relations, and constraint verification consistently fall below 50% accuracy. Performance gains level off beyond approximately 30 billion parameters, and models optimized for chain-of-thought show only modest improvements on the hardest tasks. Additionally, grid-like layouts (kitchens) yield higher accuracy than more varied configurations (living rooms, bedrooms), indicating that layout regularity influences model performance.
These results reveal that current LLMs do not form sufficiently robust internal representations for complex spatial inference. Addressing this limitation could take two complementary directions. In the near term, integrating external geometric solvers or symbolic planning modules can compensate for the models’ weaknesses in collision avoidance and clearance reasoning. In the longer term, training regimes that incorporate explicit spatial constraints—through specialized objectives, data augmentation with irregular layouts, or constraint-violation examples—may be necessary to improve performance on design-oriented tasks.
Acknowledgments
The work is supported by funding from KAUST, Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program
Appendix A Technical Appendices and Supplementary Material for PlanQA
This technical appendix includes an overview of the dataset generation process, followed by full evaluation results on our generated QA pairs, along with aggregated metrics and ablation studies. It also presents an extended taxonomy of spatial reasoning tasks, explanations of selected failure cases, an expanded analysis of vision-language models (VLMs), and examples of prompts used for both generation and questioning.
The full dataset and all code used for generation, evaluation, and visualization are provided in the supplementary materials.
Appendix B Details on Synthetic Data Generation
Room layouts were generated using Gemini 2.5 Pro, specifically fine-tuned for spatial reasoning and bounding-box tasks Team et al. (2025). Below, we describe the detailed procedure and constraints. All generation scripts, prompts, constraint-checking code, and random seeds are included in the code for full reproducibility.
B.1 Room Shape Generation
We generated 600 layouts each for kitchens, bedrooms, and living rooms. These layouts include a variety of geometries and sizes, with clearly defined room structures containing windows, doors, and corner cutouts where applicable. All layouts follow specifications tailored to each room type.
Each layout falls into one of three shape categories: rectangular, L-shaped, or open. Size categories were defined individually for each room type and were incorporated into the generation prompts based on common assumptions about typical room dimensions.
The distribution of room shapes and sizes is shown in Table 6, and example layout types are illustrated in Fig. 1.
Table 6: Shape and size distribution across generated layouts for each room type.
Room Type Shape Distribution (Rect / L-shaped / Open) Size Categories (in m2) (Small / Medium / Large)
Kitchen 40% / 40% / 20% $\leq$7 / 7–18 / $>$18
Bedroom 50% / 30% / 20% 8–12 / 12–18 / $>$18
Living Room 40% / 40% / 20% 20 / 22 / 24
Geometric and Structural Constraints:
Room geometries were procedurally varied using prompt-based guidance to produce rectangular, L-shaped, and open-plan configurations. The following structural properties were described in the prompts but not explicitly enforced during layout generation:
  • L-shaped rooms were described as rectangular spaces with a square cutout from one corner. To maintain usable proportions, each leg of the L shape was suggested to be at least 1.5,m wide and deep.
  • Open-plan rooms, apart from living rooms, were prompted without doors and with one full wall removed. This was intended to vary room shapes and simplify layout generation.
  • Doors were described with widths between 0.8,m and 1.0,m, selected randomly. Windows were chosen from a fixed set of widths: 0.6,m, 0.75,m, 0.9,m, 1.2,m, and 1.5,m.
  • Prompts included guidance to place all elements such as doors, windows, and cutouts entirely within room boundaries.
Window Placement:
Window placement followed prompt-based guidelines aimed at supporting daylight access and layout clarity:
  • The total window length was set to exceed 15 percent of the room’s floor area. It was used as a simple proxy to ensure visible window openings and visual balance along the walls.
  • Windows on the same wall were the same size to support visual balance. Small, isolated windows were not used.
  • Long windows, over 1.5 m, were split into segments with 0.05 to 0.15 m gaps for a more modular appearance.
  • Windows were not placed opposite each other or on the same wall as a door, as such arrangements are less common and were not emphasized in the prompt design.
B.2 Furniture and Appliance Placement
In the second stage, each room was populated with furniture and appliances based on layout style and object-specific constraints. While the styles differ by room type–such as enforcing a work triangle in kitchens, orienting seating around a focal point in living rooms, or centering beds symmetrically in bedrooms–the overall placement process followed a unified set of rules:
  • All floor-standing objects must be placed without overlaps, except in semantically grouped cases (e.g., lamps on tables, chairs under tables).
  • Clearance zones must be preserved around doors, main pathways, and functional elements such as beds, appliances, and desks.
  • Placement follows a priority order: essential furniture is placed first, followed by optional and decorative elements only if space allows.
  • Major objects like fridges, ovens, and beds must be anchored to structural walls or room boundaries.
Layouts violating any of these hard constraints, due to overlap, clearance issues, or improper attachment, were automatically discarded.
B.3 Layout Selection Criteria
Approximately one-third of the initially generated room layouts were filtered-out using a set of geometric and functional constraints. These filters were designed to ensure realistic object placement, functional usability, and architectural plausibility. While these rules are based on general design principles, they are not based on any specific design standard. Instead, they are the result of iterative development, focusing specifically on addressing cases where our prompts lead to unlikely or implausible layouts. After filtering, we retained 600 valid layouts for each room type.
  • Non-overlapping objects (with exceptions): Each pair of objects must satisfy axis-aligned bounding box (AABB) separation constraints, unless they belong to a known exception category. For two objects $A$ and $B$ with bounding boxes $(x_{1}^{A},y_{1}^{A},x_{2}^{A},y_{2}^{A})$ and $(x_{1}^{B},y_{1}^{B},x_{2}^{B},y_{2}^{B})$, non-overlapping requires:
    $x_{2}^{A}\leq x_{1}^{B}\quad\text{or}\quad x_{2}^{B}\leq x_{1}^{A}\quad\text{or}\quad y_{2}^{A}\leq y_{1}^{B}\quad\text{or}\quad y_{2}^{B}\leq y_{1}^{A}$
    This constraint is not enforced for the following semantically compatible object pairs: (i) rug with any object placed on top of it; (ii) lamp with nightstand, desk, or table; (iii) tv with tv_stand; (iv) chair objects with desk or table.
    These exception pairs are considered contextually collocated or hierarchically related (e.g., support/surface relationships) and are therefore allowed to overlap.
  • Non-blocking door clearance: Doors have physical thickness and are defined by bounding boxes $(x_{1}^{d},y_{1}^{d},x_{2}^{d},y_{2}^{d})$. A clearance zone of $door\_length$ meters is required in front of the door to ensure swing space and accessibility. The position of this zone depends on which wall the door is attached to.
  • No windows on opposite walls: We did not include layouts with windows on directly opposite walls, as such configurations are uncommon in typical residential designs.
  • Appliances against walls or cutout edges: Large fixtures like fridges and ovens must be flush against at least one wall or cutout boundary. This is formalized by enforcing:
    $x_{1}=0\quad\text{or}\quad x_{2}=W\quad\text{or}\quad y_{1}=0\quad\text{or}\quad y_{2}=D\quad\text{or edge of cutout}$
    For rooms with cutouts, an object may align with a cutout boundary, defined as additional wall segments with known coordinates.
These constraints were iteratively selected to address common implausible layouts generated by Gemini 2.5 Pro using our prompts. They are not universal requirements for real layouts, nor a complete set of constraints, but aim to avoid frequent sources of implausibility in generated layouts.
B.4 Room Statistics
Table 7 summarizes key statistics for the generated layouts across room types, including average object count, overlap frequency, and spatial density. Kitchens typically contain fewer objects but have higher density due to their compact size. In contrast, bedrooms and living rooms host more objects with greater overlap, reflecting more open layouts and flexible furnishing.
This structural simplicity, with fewer objects, minimal overlap, and higher density, makes kitchens the simplest environment for spatial reasoning tasks. Due to the limited amount of free space, some complex questions, such as those involving paths, have fewer possible answers. As a result, models tend to perform better on kitchen-related questions.
Table 7: Average layout statistics by room type.
Metric Kitchen Bedroom Living Room
Avg. # of Objects 13.30 14.24 15.29
Avg. # of Overlaps 0.52 6.27 6.26
Avg. Object Density 1.15 0.84 0.71
Appendix C Full Breakdown by Room Type, Question Category, and Model
Table 8 reports per-model accuracy separated by room type (kitchen, living room, bedroom) and question category. This is the full unaggregated results matrix underlying the analyses presented in Section 4. The table includes all eleven evaluated models and eleven question types. Path is assessed using two complementary metrics: validity and Fréchet Distance. Together, these question types span a range of challenges, covering both reasoning-heavy and functionally grounded tasks.
The table presents accuracy for all eleven models across eleven question types, grouped by room type. Each room type–kitchen, bedroom, and living room–includes results for all question types, totaling 33 rows. Each value reflects model accuracy over a fixed set of 600 questions, corresponding to 600 layouts per room type. For every combination of question type and room type, the same set of 600 questions is used across all models to ensure fair comparison.
Table 8: Question-level accuracy by category for each model across room types. Models (left to right): Qwen3-32B, DeepSeek-R1, DeepSeek-V3, LLaMA 3.3-70B-Instruct, Gemma 2-27B-it, Phi-4, GPT-4.1, LLaMA 3.1-8B-Instruct, Gemma 2-9B-it, Phi 3.5-mini, GPT-4o-mini.
Note: High truncation rate due to token limits.
Reasoning Big Models Small Models
N 32.8B 671B 671B 70B 27.2B 14B N/A 8B 9.2B 3.8B N/A
Temp. 0.6 0.6 0.3 0.0 0.5 0.0 1.0 0.6 0.5 0.7 1.0
Type Category [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
K Distance 97.2 99.5 98.5 97.2 94.7 90.7 100 82.3 74.2 35.3 97.3
Area (counters) 85.3 99.5 95.7 81.7 46.5 79.2 98.5 30.8 37.5 9.8 80.3
Free Space 83.7 88.0 52.5 37.0 23.5 42.0 83.5 9.5 13.5 6.3 20.3
View Angle 74.0 58.2 69.2 75.2 12.7 52.8 86.3 11.5 9.2 8.2 42.2
Repositioning 91.5 96.3 69.8 48.7 14.8 22.7 85.8 3.7 6.2 11.5 40.7
Max Box 33.3 30.7 9.2 8.8 3.0 3.0 57.2 1.7 1.2 1.7 3.7
Fit/Placement 92.8 92.5 71.0 66.2 72.0 82.0 89.8 68.7 71.8 68.8 70.7
Path (Valid) 13.2 6.3 39.3 30.3 23.2 26.8 73.0 10.5 17.5 14.8 35.7
Path (Fréchet) 15.0 6.3 32.8 30.2 26.0 26.0 55.8 10.2 19.0 9.3 36.7
Missing Object 87.3 88.7 44.3 56.2 52.3 52.2 79.8 27.5 39.7 14.3 58.0
Obstruction 84.0 95.2 32.7 6.0 3.3 9.7 93.5 1.2 2.7 11.3 14.2
L Distance 98.7 99.8 99.5 98.2 96.3 98.8 99.8 87.0 81.8 58.8 98.5
Area (sitting) 96.7 99.5 84.5 98.0 83.8 88.7 99.5 32.5 41.3 12.8 85.7
Free Space 0.3 4.3 1.0 0.3 3.7 0.3 5.0 0.7 3.2 1.5 1.7
View Angle 81.5 86.0 70.0 76.8 14.3 50.3 96.0 14.8 8.3 10.5 42.2
Repositioning 80.5 93.0 45.0 33.5 12.8 19.2 71.5 6.5 5.7 6.3 29.3
Max Box 1.5 0.8 1.0 3.0 2.0 2.8 6.5 0.8 1.5 1.0 2.3
Fit/Placement 90.7 91.2 71.5 80.0 83.7 87.7 91.8 75.0 75.5 72.8 72.7
Path (Valid) 10.0 14.7 33.0 26.7 30.5 26.7 53.2 7.3 21.7 16.7 33.5
Path (Fréchet) 13.2 14.2 23.7 17.8 19.2 14.7 48.0 2.5 14.2 6.8 25.2
Missing Object 73.0 76.2 51.0 49.3 29.5 36.0 65.5 9.7 28.3 11.7 32.3
Obstruction 80.7 96.5 24.3 7.3 3.8 9.3 84.7 2.5 5.2 4.7 11.7
B Distance 98.7 99.8 99.5 98.2 96.3 98.8 99.8 87.0 81.8 58.8 98.5
Area (storage) 98.7 99.8 94.0 97.0 86.3 88.3 99.3 30.3 66.3 47.7 88.2
Free Space 1.7 5.8 1.2 0.3 2.5 1.2 2.8 1.8 1.0 1.0 1.2
View Angle 76.0 79.8 70.0 78.3 10.8 57.0 94.2 15.3 10.2 7.7 43.0
Repositioning 78.7 94.3 53.8 36.0 11.2 15.7 73.7 4.3 9.0 6.5 31.5
Max Box 0.7 1.0 2.0 1.8 2.0 2.5 7.2 1.0 0.7 1.0 2.0
Fit/Placement 86.3 86.3 65.8 66.7 66.5 73.0 82.0 66.0 63.8 64.7 66.2
Path (Valid) 15.5 20.8 49.7 40.7 38.5 35.8 67.8 11.3 34.5 20.5 48.8
Path (Fréchet) 15.3 21.5 30.3 30.3 27.0 19.0 49.5 3.7 19.7 8.5 36.3
Missing Object 64.3 65.2 39.8 40.3 33.0 25.2 61.0 14.8 19.7 8.3 34.5
Obstruction 87.2 95.7 32.0 4.3 1.3 5.8 89.5 2.2 5.3 9.3 12.2
Appendix D Additional Ablations and Analysis
To further verify the robustness and generalizability of our evaluation setup, we conducted several supplementary analyses beyond the primary results in Table 8.
Mean and Variability Across Room Types
Figure 6 shows radar plots summarizing model performance across question types, averaged by room type. Kitchens, living rooms, and bedrooms demonstrate similar high-level trends: models are generally most accurate on questions involving spatial metrics such as Distance, Area, and Fit/Placement, while struggling more with abstract or multi-object tasks such as Max Box and Path (Valid).
Figure 6 illustrates the standard deviation across models per question category. Notably, Repositioning, View Angle, Area, and Obstruction show the highest variability across models, particularly in bedrooms and living rooms. Free Space, in contrast, exhibits high variance only in kitchens, highlighting that performance inconsistency is not uniform across room types.
Figures 8 and 8 provide complementary radar plots organized by model, highlighting differences in room-specific performance. Kitchens consistently exhibit slightly higher mean accuracy across models, particularly for larger models like GPT-4.1 and DeepSeek-R1. Variance across models is relatively uniform, though larger models tend to show reduced variance overall.
Figure 5: Mean accuracy by question type across room categories. Each axis corresponds to a specific spatial reasoning skill (e.g., distance, placement, path validity).
Figure 6: Standard deviation in model accuracy across question types and room categories. Higher values indicate greater variability across models within a room type.
Figure 7: Mean accuracy across models by room category. Shaded regions reflect performance in bedrooms, kitchens, and living rooms, highlighting which models generalize best across spatial tasks.
Figure 8: Standard deviation of model performance across room types. Captures how consistent each model is across different spatial environments. Narrower shapes indicate more uniform behavior.
Prompt Sensitivity Ablation
To assess model robustness to prompt variation, we conducted an ablation in which each question was regenerated using the same template but with alternate object references. For example, a question originally referring to a "sofa" might instead use a "bookshelf" in the regenerated version. This allowed us to evaluate whether model performance remains stable under changes in object content while preserving linguistic structure.
As shown in Table 9, performance generally remained consistent, with only minor differences across models. Larger models such as GPT-4.1 were more stable, while smaller models like GPT-4o-mini showed slightly higher sensitivity. These results suggest that the evaluation framework is robust to this type of variation.
Table 9: Accuracy under prompt variation. Each cell shows performance on the original prompt followed by a regenerated version with alternate object references.
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 75.3 92.2 → 91.4 75.5 → 74.9
GPT-4o-mini 33.8 → 32.9 42.4 → 46.7 12.2 → 14.1
DeepSeek-R1 83.6 → 95.7 74.7 → 72.1 95.7 → 95.6
Layout Perturbation Ablation
To test model sensitivity to small physical changes, we introduced random spatial jitter of up to ±0.1 m to all object coordinates during evaluation. This simulated minor but plausible noise in the layout. After jittering, we recomputed the correct answers for all QA pairs based on the updated room configuration. The questions remained unchanged, but all evaluations used the perturbed layouts.
The jitter was constrained to ensure that objects did not overlap, remained within room boundaries, and did not block doors. Jittering can affect spatial relationships such as clearances, alignments, and angles. For example, a path that was previously clear might become blocked, or an object may no longer fit within a gap.
As shown in Table 10, model accuracy showed slight fluctuations in response to the spatial changes. Overall, model performance remained robust, indicating resilience to small, physically realistic layout variations.
Table 10: Accuracy before and after layout jitter. Values shown for three spatially sensitive questions.
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 72.5 92.2 → 87.3 75.5 → 80.2
GPT-4o-mini 33.8 → 28.4 42.4 → 42.5 12.2 → 11.8
DeepSeek-R1 83.6 → 89.8 74.7 → 70.3 95.7 → 97.1
Appendix E Spatial Reasoning Question Suite
To extend our taxonomy of question types introduced in Table 2, Table 11 details the specific evaluation procedures used to compute ground truth answers for each category. In addition to categorizing the task, the table shows how answers are derived and what types of spatial reasoning each task demands. This breakdown clarifies not only what is being tested, but how it is measured, supporting both reproducibility and deeper analysis of model performance in later sections.
Table 11: Categories of spatial reasoning questions, answer-checking procedures, and diagnostic insights.
Category Sample Question Answer Checking Diagnostic Insight
Distance What is the Euclidean distance between the center of the door and the fridge? Compute centroids $(x,y)$ of both boxes; calculate Euclidean distance between them; compare with model output allowing a tolerance of $2\%$ of the ground-truth distance. Tests numeric computation, coordinate extraction, and arithmetic precision.
Area What is the area of the kitchen counter? Multiply the width $\times$ height of each cabinet; sum all resulting areas; compare the total against the model output with a 2% tolerance. Probes multiplicative reasoning, aggregation, and unit consistency.
Free Space What is the total unoccupied floor area in the bedroom? Merge overlapping object bounding boxes; sum the resulting areas; subtract from total room area; compare the unoccupied area with model output using a 2% tolerance. Measures the ability to aggregate and subtract multiple regions.
View Angle What is the dot product between the facing vector of the sofa and the vector to the TV? Compute normalized facing vector of reference; get normalized vector to target center; compare dot product with model output using 2% tolerance. Evaluates directional reasoning, vector normalization, and spatial alignment understanding.
Repositioning How far can the sofa move right before hitting a wall or another object? Simulate axis-aligned movement; measure distance to furthest valid non-overlapping position; compare with model output using 2% tolerance. Tests spatial simulation, collision detection, and directional constraint reasoning.
Max Box What is the largest axis-aligned box you can place in the living room? Use brute-force to solve largest empty rectangle in free space; compare resulting dimensions with 2% tolerance. Probes continuous-space optimization and spatial coverage estimation.
Fit/Placement Can a $2\,\mathrm{m}\times 3\,\mathrm{m}$ table fit in the room without overlap? Attempt to place a rectangular object in free space; ensure no intersection with existing bounding boxes. Compare boolean result with model prediction. Tests collision detection, spatial constraint enforcement, and free-space reasoning.
Shortest Path What is the shortest walkable path from the fridge to the exit with $15\,\mathrm{cm}$ clearance? Offset free space by $15\,\mathrm{cm}$ for clearance; use A* algorithm on navigable grid to compute path; compute Fréchet distance (threshold $0.6\,\mathrm{m}$) to model path; check its validity. Tests path planning, clearance-aware modeling, and heuristic graph search.
Missing Object Where should the missing armchair be placed in this layout? Check if predicted bounding box fits in free space without overlaps and satisfies clearance/accessibility constraints. Evaluates spatial reasoning, free-space utilization, and functionally appropriate object placement.
Obstruction Which objects obstruct the line of sight between the bed and the window? Compute vector from target to reference center; check which bounding boxes intersect the segment (excluding endpoints). Tests geometric reasoning, line–rectangle intersection, and occlusion detection.
Appendix F Failure Cases
To better understand the sources of model failure, we conducted a qualitative analysis of representative examples from the benchmark. This section presents visualizations of selected test layouts alongside model responses. By examining both correct and incorrect outputs, we aim to identify common failure patterns and reasoning bottlenecks across different architectures.
F.1 Max Box
As illustrated in Figure 10, this example shows GPT-4.1 producing a suboptimal result in a spatial reasoning task. Instead of identifying the largest free region in the kitchen layout, the model placed a bounding box that extended along a single axis, starting from one edge, without fully exploring the layout. It selected an area from x=0 to 2.1 and y=1.1 to 1.8, yielding 1.47 m² of free space.
A larger available region, shown in Figure 10, spans from (0.7, 0.7) to (2.1, 1.8), and offers 1.54 m². The model did not identify this larger region, which indicates that its search strategy was likely not exhaustive, even though it had access to the full structured input. The reasoning trace in Figure 11 confirms that it evaluated a set of candidate regions but ultimately settled on a suboptimal one. The criteria for candidate selection remain unclear and may vary across queries.
Figure 9: GPT-4.1 prediction
Figure 10: Ground truth
LLM Reasoning Summary Let’s analyze the problem step by step. First, I’ll look for large free areas without obstacles. Option 1: The region to the right of the chairs and table, from x=1.3 to 2.1, and from y=0 to 1.8. This gives an estimated area of 1.44 m². Option 2: The space above the chairs and below the wall units, from x=0 to 2.1 and y=1.1 to 1.8. This gives an area of 1.47 m². Option 3: A small gap near the bottom right corner. It looks too narrow and likely blocked by furniture. Option 4: The space in front of chair_2, but it appears too small to be useful. Based on these, Option 2 seems best. I’ll choose the rectangle [0.0, 1.1, 2.1, 1.8], which gives an area of 1.47 m².
Figure 11: Summarization of chain-of-thought style output from GPT-4.1. The model imitates reasoning by listing and evaluating candidate regions, then selects a suboptimal solution.
F.2 Shortest Path
In this task, the model must find a path from the center of a starting object to the center of a destination object. For simplicity, the start and goal objects are not treated as obstacles.
Our pathfinding system constructs a visibility graph by connecting all mutually visible corner points in the environment. Dijkstra’s algorithm is then used to compute the true shortest path over this graph, ensuring optimality under Euclidean distance. This method avoids the limitations of grid-based approximations and produces precise, flexible paths around obstacles. While multiple valid solutions may exist depending on object layout and clearance constraints, the chosen path always reflects the shortest feasible route.
When interpreting the task, models often fail to make effective use of available free space. Instead of navigating around obstacles or adapting to the environment, they frequently extend a straight line in one direction until it is blocked. This leads to incomplete or inefficient trajectories that do not reflect the geometry of the room or required clearance. In some cases, paths cut too closely around objects, likely due to the model implicitly minimizing Euclidean distance without accounting for spatial constraints.
Figure 13 shows a path that reaches the target but fails the Fréchet distant constraint of 0.6 m, making it non-optimal. Figure 13 depicts a path that violates clearance constraints and does not represent a valid trajectory to the target.
Figure 12: Comparison between the optimal path (green) and the GPT-4.1-generated trajectory (cyan), which, despite reaching the target, exceeds the Fréchet distance threshold of 0.6 m.
Figure 13: Depiction of the optimal path (green) alongside the GPT-4.1-generated trajectory (red), which fails to satisfy clearance requirements and results in an infeasible route.
Appendix G VLM Experiments
To further probe model performance on particularly challenging tasks, we conducted targeted evaluations of two vision-language models (VLMs). We focused on the most difficult question categories–Free Space and Max Box–within the most spatially complex environments: living rooms and bedrooms. The models tested in this focused benchmark include Gemini 2.5 Flash Preview 04-17 Gemini Team, Google (2025) and GPT-4.1. Results are summarized in Table 12, which compares each model’s output under two input conditions: JSON only and JSON combined with a floorplan image. An example of Gemini’s step-by-step reasoning under the JSON+Image condition is shown in Figure 14.
Table 12: Comparison of spatial understanding between Gemini 2.5 Flash and GPT-4.1 for Living Room (L) and Bedroom (B), under different conditions.
Type Category [Uncaptioned image] Gemini 2.5 Flash [Uncaptioned image] GPT-4.1
JSON Only JSON + Image JSON Only JSON + Image
L Free Space 6.4 12.2 5.0 9.8
Max Box 17.0 22.0 6.5 4.3
B Free Space 12.8 15.2 2.8 10.5
Max Box 15.7 21.0 7.2 7.2
[…]
Figure 14: Summarized step-by-step reasoning by Gemini to compute maximum free space from a JSON floorplan and image. The model merges nested objects and analyzes rectangular gaps to estimate usable area.
While our results show that both models benefit from visual input in some settings, the extent of improvement varies by task. For the Free Space category, both Gemini 2.5 Flash and GPT-4.1 show substantial gains when provided with floorplan images. In contrast, for Max Box, only Gemini shows a notable improvement, while GPT-4.1 performs similarly or worse with visual input. This variation suggests that visual grounding supports some spatial tasks more effectively than others.
The stronger visual performance of Gemini 2.5 Flash may stem from architectural or training differences, possibly reflecting design choices aimed at robotics or spatial reasoning tasks.
One challenge lies in the nature of floorplan inputs. Unlike natural images, floorplans vary widely in style, abstraction, and labeling. This inconsistency makes it harder for models to extract spatial meaning reliably.
Smaller models show little benefit from visual input, likely due to limited capacity and lack of pretraining for structured spatial tasks.
In summary, floorplan understanding requires more than object detection. It demands precise spatial reasoning and integration with structured formats like JSON. Without task-specific tuning or standardized visual formats, even advanced models remain limited in this domain.
Appendix H Prompts
This section presents selected prompt templates used in our benchmark. We include a summarized version of the bedroom layout generation prompt, as well as representative examples of question prompts targeting specific spatial reasoning skills. These prompts reflect the structured task design outlined in Table 11. Additional prompts–including full versions, formatting logic, and implementation details–are provided in the supplementary code and materials for reproducibility.
Prompt: Max Box You’ll be given a layout of a {room_type} in JSON format. Your task is to calculate the area (in m²) of the largest axis-aligned rectangle that can fit within the room. The rectangle must not overlap with any other objects, but it may fully occupy free space, including areas above rugs. You can ignore passageways, door clearance, or accessibility constraints. Room layout: {room_json} Reason step by step to calculate the area. Respond only with a float value in square meters using the format: *Final answer*: area
Figure 15: Prompt for computing the largest empty rectangle area within a room layout using Chain-of-Thought reasoning.
Prompt: Missing Object Placement You’ll be given a layout of a {room_type} in JSON format. Your task is to suggest a suitable placement for the missing object {object_name} by providing its bounding box coordinates. The placement must avoid overlaps, respect obstacles, and maintain accessibility. Room layout: {room_json} Reason step by step to identify the placement. Respond only with bounding box coordinates (float values) using the format: *Final answer*: [x_min, y_min, x_max, y_max]
Figure 16: Prompt for placing a missing object in a room layout using Chain-of-Thought reasoning. The model must return valid coordinates that satisfy geometric constraints and maintain functional space.
[…]
Figure 17: Summarized data generation prompt for producing structured and constrained bedroom layouts.
Prompt: Fill Bedroom Layout with Objects Given a predefined bedroom layout style and room geometry in JSON format, generate a filled 2D bird-view layout. Include a list of placed objects with their bounding boxes and explanatory comments. Essential fields: Each object must have a "label", "bbox" ([y0, x0, y1, x1]), and a descriptive "comment". Furniture labels include: "bed", "nightstand", "dresser", "wardrobe", "desk", "chair", "armchair", "rug", "lamp", etc. Architectural elements ("door", "window", "cutout_area", "fireplace", "closet_alcove") must match the input layout and remain unmodified. Placement priorities: 1. Place the "bed" according to the bed_size_suggestion and layout style. 2. Add essential storage: "dresser", "wardrobe", or use "closet_alcove" if defined. 3. Add secondary items (e.g., "nightstand", "desk", "chair") only if space and clearance allow. 4. Add decorative or optional items ("rug", "mirror", "floor_lamp", "plant") last. Constraints: Maintain at least 0.75 m clearance for walkways and door swing. Beds require 0.6–0.75 m of access space on sides and foot (unless against wall). Wardrobes/dressers need 0.6–0.8 m clearance for drawer/door use. No object overlap (except table lamps on nightstands or rugs under furniture). Use walls efficiently; avoid blocking windows unless unavoidable. Ensure mirror has 0.75 m clearance in front; treat "rug" as an anchor but optional. Final output: a JSON list of objects, including placement and comments. No layout geometry should be altered.
Figure 18: Prompt for populating a bedroom layout with functionally and spatially valid object placements, following layout-specific design rules.

### v2

@@
-PlanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations
+FloorplanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations
 Abstract
-We introduce PlanQA, a diagnostic benchmark for evaluating geometric and spatial reasoning in large-language models (LLMs).
-PlanQA is grounded in structured representations of indoor scenes, such as kitchens, living rooms, and bedrooms, encoded in a symbolic format (e.g., JSON, XML layouts).
-The benchmark includes diverse question types that test not only metric and topological reasoning (e.g., distance, visibility, shortest paths) but also interior design constraints such as affordance, clearance, balance, and usability.
-Our results across a variety of frontier open-source and commercial LLMs show that while models may succeed in shallow queries, they often fail to simulate physical constraints, preserve spatial coherence, or generalize under layout perturbation.
-PlanQA uncovers a clear blind spot in today’s LLMs: they don’t consistently reason about real-world layouts.
-We hope that this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
+We introduce FloorplanQA, a diagnostic benchmark for evaluating spatial reasoning in large-language models (LLMs).
+FloorplanQA is grounded in structured representations of indoor scenes, such as (e.g., kitchens, living rooms, bedrooms, bathrooms, and others), encoded symbolically in JSON or XML layouts.
+The benchmark covers core spatial tasks, including distance measurement, visibility, path finding, and object placement within constrained spaces.
+Our results across a variety of frontier open-source and commercial LLMs reveal that while models may succeed in shallow queries, they often fail to respect physical constraints, preserve spatial coherence, though they remain mostly robust to small spatial perturbations.
+FloorplanQA uncovers a blind spot in today’s LLMs: inconsistent reasoning about indoor layouts.
+We hope this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
 1 Introduction
-Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over physically grounded environments remains poorly understood.
-In numerous real-world applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images or natural language.
+Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over plausible, physically feasible environments such as indoor layouts remains poorly understood.
+In numerous practical applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images or natural language.
 Reasoning in these contexts requires geometric inference over symbolic layouts, not pixel-level perception.
-We introduce PlanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floor plans represented in structured text-based formats.
+We introduce FloorplanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floorplans represented in structured text-based formats.
 Each instance consists of a JSON-encoded layout paired with natural language questions that require the model to compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints.
-PlanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments.
+FloorplanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments.
 Although LLMs can increasingly be used in tool-assisted pipelines, for example to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities.
-PlanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding.
+FloorplanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding, in order to measure their unassisted capabilities.
+This baseline is important because even in tool-rich systems, models benefit from some unaided spatial ability to anticipate outputs and avoid trivial errors.
 Specifically, our contributions are as follows:
-  • We introduce a dataset of 1,800 structured 2D layouts, 600 each from kitchens, living rooms, and bedrooms, represented in JSON and paired with spatial reasoning questions.
-  • We provide a diverse suite of 18,000 questions spanning geometric, functional, and design-aware reasoning, with one question per layout across multiple categories.
+  • We introduce a dataset of 2,000 structured 2D layouts, including 600 each from synthetically generated kitchens, living rooms, and bedrooms, plus 200 layouts sourced from the Habitat Synthetic Scenes Dataset (HSSD) (Khanna et al., 2023), providing a realism check.
+All are represented in JSON and paired with spatial reasoning questions.
+  • We provide a diverse suite of 16,000 spatial reasoning questions, eight questions per layout, covering geometric relations, placement feasibility, spatial occupancy, and navigation.
   • We establish structured evaluation protocols and scoring metrics that enable a fine-grained diagnosis of reasoning performance by task type and error mode.
-  • We conduct a comparative analysis of 11 LLMs, including GPT-4.1 OpenAI et al.
-(2024) and open models such as DeepSeek-R1 DeepSeek-AI (2025) and Llama-3.3-70B-Instruct Grattafiori et al.
-(2024), revealing consistent failure patterns in spatial inference from symbolic input.
-PlanQA provides a diagnostic substrate for measuring spatial competence in language models, with applications in layout validation, agent planning, and symbolic grounding.
+  • We conduct a comparative analysis of 15 LLMs, including 7 reasoning-focused models, as well as 8 standard models, revealing consistent failure patterns in spatial inference from symbolic input.
+FloorplanQA provides a benchmark of layouts, questions, and evaluation metrics for assessing spatial reasoning in language models, focusing on symbolic floorplans that integrate geometry and semantics in ways that mirror real architectural abstractions.
 2 Related Work
@@
 CLEVR Johnson et al.
-(2017) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relationships.
+(2017) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relations.
 In real-world settings, SpatialSense Yang et al.
@@
 (2024) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts.
-Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but suffers from poor object placement, often yielding physically implausible or random layouts due to non-convergent simulated annealing.
+Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but often produces implausible object placement due to non-convergent simulated annealing.
 LayoutVLM Sun et al.
@@
 (2025) improve layout generation via optimization and constraint solving, respectively.
-But they assess spatial reasoning through output realism, not direct inference.
+But they assess output realism, not the model’s ability to infer constraints directly.
 In contrast, our benchmark tests symbolic reasoning without tool-assisted refinement.
@@
 While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning in structured layouts, providing fine-grained and task-specific insights into models’ spatial competence.
-Recent 3D-LLM surveys such as Ma et al.
-(2024b) cover tasks like navigation and interaction, but not symbolic spatial reasoning.
-PlanQA fills this gap by testing raw spatial competence from structured layouts without multimodal input.
-PlanQA addresses the gap in existing benchmarks by directly evaluating structured spatial inference from symbolic room layouts.
-Unlike prior benchmarks relying on raw images or focusing on commonsense spatial language, PlanQA provides explicit spatial representations (object coordinates and dimensions) and tests models’ abilities to perform precise spatial reasoning tasks, such as calculating distances, assessing visibility, and verifying object fit within a controlled setting.
+Recent 3D-LLM surveys such as (Ma et al., 2024b) cover tasks like navigation and interaction, but not symbolic spatial reasoning.
+FloorplanQA fills this gap by testing raw spatial competence from structured layouts without multimodal input.
+FloorplanQA addresses the gap in existing benchmarks by directly evaluating structured spatial inference from symbolic room layouts.
+Unlike prior benchmarks relying on raw images or focusing on commonsense spatial language, FloorplanQA provides explicit spatial representations (object coordinates and dimensions) and tests models’ abilities to perform precise spatial reasoning tasks, such as calculating distances, assessing visibility, and verifying object fit within a controlled setting.
 3 Method
 3.1 Synthetic Layout Generation
-Our initial aim was to utilize publicly available real-world floorplan datasets.
-However, a comprehensive review revealed critical limitations.
-Prominent datasets such as SUNCG (Song et al., 2017) and HouseExpo (Li et al., 2019) were withdrawn due to unresolved copyright claims, rendering them unusable.
+Our initial aim was to use publicly available real-world floorplan datasets.
+However, a comprehensive review revealed significant limitations, as several prominent datasets such as SUNCG (Song et al., 2017) and HouseExpo (Li et al., 2019) are not accessible due to unresolved copyright claims.
 Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2020), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse.
@@
 RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al.
-(2020), while large, appears procedurally generated albeit with realtor supervision and imposes restrictions on redistribution and derived works.
+(2020), while large, is procedurally generated with realtor supervision and imposes restrictions on reuse.
 Given these legal, practical, and methodological constraints, we found synthetic data generation to be the most viable alternative.
-We generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro, a large language model fine-tuned for spatial reasoning (Team et al., 2025).
-While employing an LLM for both data generation and evaluation may raise concerns of circularity, our evaluation pipeline is entirely independent: the LLM is used solely for data generation, and evaluation is performed via deterministic spatial reasoning tasks.
+We generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro, a large language model fine-tuned for spatial reasoning (Google Gemini Robotics Team, 2025).
+Although our evaluation includes multiple LLMs (including Gemini variants), there is no circularity: the data-generation step is one-off and separate from evaluation.
+For evaluation, models solely produce answers to the benchmark questions, and correctness is computed by deterministic geometric routines against ground-truth solutions derived from the layouts, so no model output is used to generate data or to grade itself.
 The generation process comprises two stages.
@@
 These constraints are encoded directly in the LLM prompt.
-Second, each room is furnished according to style-specific rules, also defined in structured prompts, to encourage both visual realism and functional plausibility.
-Approximately one-third of candidate layouts are filtered out by automated spatial validity checks, which remove configurations with inaccessible furniture or implausible adjacencies.
-This curation pipeline yields a diverse and geometrically coherent dataset.
-Full prompts, generation templates, and validation scripts are provided in the supplemental materials.
-Representative examples of layout types are shown in Fig.
-1.
-A breakdown of room types and layout geometries is provided in Table 1.
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Figure 1: Representative layouts from the PlanQA dataset, spanning three room types (kitchen, living room, bedroom) and three geometric configurations (rectangular, L-shaped, open).
-Layouts are procedurally generated using explicit spatial constraints and validated for functional feasibility.
-Table 1: Layout distribution in PlanQA by room type, style, and geometry, with room-wise totals.
+Second, each room is furnished according to style-specific guidelines (e.g., a bedroom must contain a bed and storage), also defined in structured prompts, to encourage both visual realism and functional plausibility.
+Approximately one-third of candidate layouts are filtered out by a rule-based spatial validity filter that enforces basic clearance and accessibility constraints.
+The checks remove scenes with inaccessible furniture and implausible adjacencies, such as sofas blocking doors or a refrigerator overlapping a table; see Appendix B.3 for the full set of cases.
+Full prompts, generation templates, and validation scripts are provided in the Supplementary Material.
+3.2 Layout Extraction from HSSD Dataset
+To complement the synthetically generated layouts, we further incorporated 200 layouts extracted from the Habitat Synthetic Scenes Dataset (HSSD-200).
+HSSD provides 211 high-quality, human-authored 3D scenes designed with the Floorplanner interface and populated with 18,656 objects across 466 semantic categories.
+Unlike purely procedural datasets, HSSD offers fine-grained semantics, 3D assets, and close correspondence to real interiors, making it an effective proxy for real-world interior layouts.
+For our purposes, we project each 3D scene to a 2D floorplan and retain only select structural and furniture elements.
+Decorative or auxiliary objects (e.g., vases, plants, cushions, artworks, posters, bottles, shoes, candles) are removed to reduce clutter; see Appendix C for the full details.
+We then use an $\alpha$-convex hull (Asaeedi et al., 2014; Edelsbrunner & Mücke, 1994) to smooth object boundaries, yielding polygonal layouts that are not restricted to axis-aligned rectangles.
+This step is necessary because raw HSSD projections often produce overly dense polygons, with redundant vertices along straight or nearly straight segments; applying an $\alpha$-hull reduces spurious complexity while preserving concavity, which avoids unnecessary token overhead in downstream LLM processing.
+This ensures compatibility with our synthetic layouts, while maintaining the richer geometric variety of HSSD.
+3.3 Unified Dataset
+Together, these two sources yield a dataset of 2,000 layouts: 1,800 synthetically generated via Gemini 2.5 Pro and 200 extracted from HSSD.
+The two subsets share a unified polygonal representation, enabling consistent downstream processing.
+Figure 1 shows examples from both sources while Table 1 summarizes room, style, and geometry distributions across the synthetic subset with the figure next to it providing a breakdown by room type.
+Table 1: Layout distribution in FloorplanQA by room type, style, and geometry, with room-wise totals.
 Room Type Style Rectangular L-Shaped Open Total
-Kitchen L-Shaped (incl. counter) 208 36 105 600
-U-Shaped / G-Shaped 42 128 2
+Kitchen L-Shaped 208 36 105 600
+U/G-Shaped 42 128 2
 Island-Based 2 0 33
-Wall / Galley (compact layouts) 17 10 17
+Wall / Galley 17 10 17
 Living Room Fireplace-Centric 55 1 34 600
@@
 TV-Focused 78 3 28
-Bedroom Bedroom with Workstation 98 35 28 600
-Traditional Bedroom 114 59 57
-Efficient Small Bedroom 65 3 8
-Welcoming Guest Bedroom 75 30 28
+Bedroom With Workstation 98 35 28 600
+Traditional 114 59 57
+Efficient Small 65 3 8
+Welcoming Guest 75 30 28
 Total 893 484 423 1800
-3.2 Question Taxonomy, Prompting, and Cognitive Scope
-PlanQA assesses spatial reasoning by presenting models with a single natural language question per symbolic layout.
-Questions span a range of geometric, topological, and functional types, including numeric computations (distances, areas), spatial feasibility (object fit, placement), occlusion and visibility, and environment-level assessments such as missing elements or requirement violations.
-Examples include queries like whether a 2 m $\times$ 3 m table can be placed without blocking passage, whether the television is visible from a particular seat, or which essential appliances are absent from a proposed kitchen.
-Some questions require local metric reasoning; others test whether a model can simulate constraints that would arise in physical occupancy or usability.
+Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 1: Representative layouts from FloorplanQA. Generated: kitchen, living room, bedroom. HSSD: last image. Generated objects are axis-aligned boxes; HSSD uses arbitrary polygons.
+3.4 Question Taxonomy and Prompting
+FloorplanQA assesses spatial reasoning by presenting models with a single natural language question per symbolic layout.
+Questions span a range of topological and functional types, including numeric computations (distances, areas), spatial feasibility (object placement), visibility, and requirement violations.
+Some questions require fine-grained metric reasoning, others test whether a model can respect physical constraints.
 A categorized list of question types is shown in Table 2.
-Each question is generated by instantiating a parameterized template against the specific geometry of a layout.
-These templates are natural language prompts that fill in object names, measurements, and task-specific context.
-Prompts are issued in zero-shot settings with no few-shot scaffolding or role-based instruction, but include consistent structural cues to support stepwise inference.
-Each prompt ends with the string ”Reason step by step to solve the problem.” followed by “Final Answer:”, enabling unambiguous output extraction.
-For instance, a placement query might read: “Given the room layout below, compute the area of the largest rectangle that can be placed in the kitchen without overlapping any object.
-Reason step by step to solve the problem.
-Final Answer: [value in square meters].”
-Table 2: PlanQA question taxonomy.
+Each question is generated by filling a parameterized template with layout-specific variables such as object names, measurements, and task-specific contextual information (e.g.
+units for distance, clearance for paths, or which object-types should not occlude visibility).
+Prompts are issued in zero-shot settings, without few-shot examples or role-based instructions.
+Instead, we enforce simple structural markers—such as a required checklist and a final-answer line—to encourage stepwise reasoning.
+To ensure verifiable outputs, each prompt specifies a response schema consisting of a brief structured justification and a final answer line. For example, a distance query is phrased as:
+Prompt: Distance Query If the format, object names, or required inputs are missing, invalid, or inconsistent, the model must return:  *Final answer*: ERROR.
+Otherwise, responses must follow the scheme, for example: Response Schema This structure invokes step-by-step reasoning and each question ends with *Final Answer*: <answer>, enabling robust extraction even when APIs lack native structured-output support.
+Table 2: FloorplanQA question taxonomy.
 The example question shown is an instantiation of the template used to generate all questions of that type.
-Each task is labeled with a format code, cognitive skill category (see Tables 4 and 4), and the room types in which it is applied: K = kitchen, L = living room, B = bedroom.
-Category Example Question Format Target Rooms
-Distance What is the distance between the refrigerator and the stove? N M K, L, B
-Area What is the total sitting area? N M K, L, B
-Free Space What is the unoccupied floor area in the room? N T K, L, B
-View Angle What is the angle between the TV’s facing direction and the viewer’s line of sight from the armchair? N V K, L, B
-Repositioning How far can the ottoman be moved to the left direction until it touches another object or the wall? N G K, L, B
-Max Box What is the area of the largest axis-aligned rectangle that fits in the free space? N G K, L, B
-Fit/Placement Can a 2m × 3m desk table fit in the room without overlap? B G K, L, B
-Shortest Path What is the valid shortest path from the stove to the exit with clearance? S G, R K, L, B
-Missing Object What is a possible placement for the missing sink? S R, T K, L, B
-Obstruction What are all the objects that intersect the straight line from the fireplace to the door? L V K, L, B
-Table 3: Answer format and scoring codes.
-Code Format Scoring
-N Scalar Rel. error $\leq$ 2%
-B Boolean Exact match
-S Sequence Fréchet $\leq$ 0.6 m
-L List Set equality
-C Coord. Spatial predicate
-Table 4: Cognitive skill clusters.
-Code Cognitive Target
-M Metric computation
-T Topology
-G Geometric planning
-V Visual logic / occlusion
-R Rule-based validation
-Structured layouts are flattened into lists of labeled objects, each defined by an axis-aligned bounding box.
-Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents are unique and stable across prompt construction and answer evaluation.
-All coordinates are expressed in meters, using a bottom-left origin and $(x,y)$ ordering.
-There is no hierarchical or typed schema beyond the flat object list, and no implicit conventions are assumed.
-A single system prompt is prepended to every interaction, specifying general and room-specific constraints: “You are a spatial assistant.
-Assume that all windows have zero thickness.” For layouts of type living room or bedroom, the prompt further clarifies that rugs may be overlapped by other objects, allowing for more realistic spatial arrangements involving soft surfaces.
-Some question types are instantiated deterministically given the layout geometry (e.g., exact distances between centroids), while others are drawn from constrained samplers over feasible object pairs, free-space regions, or task-specific heuristics.
-The prompt templates themselves remain fixed across layouts and models.
-This uniformity allows us to isolate spatial reasoning under consistent linguistic and structural conditions.
-These cognitive categories correspond to distinct reasoning faculties.
-Metric tasks (M) require explicit numerical computations over distances, areas, or positions, often grounded in real-valued geometric attributes.
-Topological tasks (T) depend on spatial relationships like adjacency or containment that are invariant under metric transformations.
-Geometric planning (G) involves multistep spatial inference over object configurations, including feasibility, collision avoidance, or spatial optimization (e.g., placing a large box without overlap).
-Visual logic (V) tasks such as visibility and angle estimation rely on implicit occlusion reasoning or directional projection.
-Rule-based tasks (R) involve symbolic constraints, such as verifying whether a layout satisfies clearance or adjacency requirements or whether essential objects are missing.
-These clusters are intended to reflect the diversity of tasks while enabling diagnostic grouping of error types.
-They also reflect a spectrum of abstraction, fromm low-level metric properties to high-level design affordances.
-While these clusters are not mutually exclusive, they provide a useful abstraction for analyzing model behavior and diagnosing failures.
-For example, we hypothesize that questions involving geometric planning and occlusion (G, V) will be disproportionately challenging for models without grounded spatial priors.
-This hypothesis is revisited in our results analysis (§4).
-The formatrmat and target codes are defined in Tables 4 and 4, and used throughout Table 2.
+Each task is labeled with a format code: N (scalar), B (boolean), S (sequence), and L (list), and a question reasoning category.
+Type Example Question Format Category
+Distance Calculate the Euclidean distance in meters between the centroids of the fridge and the stove N Metric
+Free Space Calculate the total non-occupied floor area in square meters N Topology
+View Angle Compute the smallest absolute angle in degrees between the vector from the centroid of the sofa to the centroid of the TV and the global north vector (0, 1) N Metric
+Repositioning Calculate how far the ottoman be moved in the left direction until it touches another object or the wall N Dynamic
+Max Box Calculate the area in square meters (m²) of the largest rectangle that can fit inside the room N Topology
+Placement Check if a 2m × 3m desk table can fit fully inside the room without overlaps B Topology
+Shortest Path Determine the shortest valid path that maintains a clearance of 15 cm from all other objects, starting from centroid of the stove and ending at the centroid of door S Dynamic
+Visibility Find all objects that intersect the vector from the centroid of the window to the centroid of the fireplace L Topology
+Layouts are represented in a structured JSON format.
+Each entry contains a layout_id, the room_type, and explicit geometric descriptions.
+The room_boundary is stored as a closed polygon, while walls are represented separately as a list of line segments.
+Openings such as windows and doors are included in a dedicated openings field rather than flattened into the object list.
+All furnishings and functional elements (e.g., bed, sofa, table) are stored in the objects list, with each object defined by a labeled polygon.
+In the synthetic data, these polygons are axis-aligned bounding boxes (four points), whereas in HSSD they can exhibit arbitrary shapes and orientations.
+Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents remain unique and stable across prompt construction and answer evaluation.
+Coordinates are expressed in meters in a right-handed 2D Cartesian frame (+x+x to the right, +y+y upward).
+The global origin $(0,0)$ is not fixed to the layout’s lower-left corner and may vary across layouts.
+The prompt used to generate examples according to this schema is in Appendix H.
+We group these questions into three reasoning categories.
+Metric tasks require explicit numerical computation, such as calculating centroids, measuring distances between objects, or evaluating the angle between an inter-object vector and a reference axis.
+Topology category involves geometric and relational reasoning, including checking placement feasibility, computing free space, or identifying whether an object blocks the direct line of sight between two others.
+Dynamic category addresses layout-changing procedures, such as repositioning an object until contact with a boundary or another object, or computing a valid collision-free path between two objects.
+These categories are intended to capture the core modes of spatial reasoning in FloorplanQA, ranging from low-level geometric calculation to higher-level relational and procedural inference.
+While not strictly disjoint, they provide a diagnostic framework for analyzing model behavior and diagnosing failure modes.
+In addition to categorizing by reasoning type, each task is also associated with an answer format code that specifies the expected output structure and the corresponding scoring rule.
+Scalar outputs (N) are scored by relative error with a default tolerance of 2%; for complex area-computation tasks (e.g., Free Space), the tolerance is relaxed to 5%.
+Sequence outputs (S) are evaluated with a Fréchet threshold of 0.6 m, approximating minimal human clearance, and must be valid (collision-free, no overlaps).
+List outputs (L) are evaluated by set equality.
+Together, the categories and format codes define the taxonomy summarized in Table 2, which reports task coverage and examples for each case.
 3.5 Evaluation Protocol and Scoring
-Each question in PlanQA is paired with a reference answer derived entirely from the symbolic layout, allowing deterministic and automated evaluation of model outputs.
-Depending on the response type, correctness is assessed using tolerance-based numeric comparison, symbolic string matching, or geometric validation routines.
-The scoring logic referenced by each question format code is detailed in Table 4, and is applied uniformly across models.
-For numerical questions (e.g., distances, areas, angles), predictions are accepted if they fall within a fixed relative error of the reference: $\left|\frac{\hat{y}-y^{*}}{y^{*}}\right|\leq\varepsilon$ where $\varepsilon$ is typically 0.02.
-This threshold is applied uniformly across models, and reflects spatial resolution limits in architectural layouts.
-Binary and categorical responses, such as visibility, object fit, or violation detection, are validated against geometry-aware checks.
-For example, visibility is determined by ray intersection tests over the layout’s free space, while object fit questions are scored by verifying non-overlap and clearance constraints for a candidate placement.
-Boolean answers must match the reference exactly.
-Structured outputs, including bounding boxes and paths, are checked for both format and semantic correctness.
-Rather than using exact coordinates, we validate these using geometric plausibility conditions—e.g., sufficient overlap with ground truth, or collision-freeness for planned paths.
-Thresholds for spatial alignment (e.g., bounding box proximity or path deviation) are set conservatively to account for minor lexical variation without over-penalizing near-correct reasoning.
-To differentiate reasoning failures from generation or parsing errors, we annotate each model response with a status flag: Correct, Incorrect, Malformed, or Missing.
-Malformed responses include answers lacking the required structure “Final Answer:” or acceptable variations thereof, or containing ill-typed data.
-These annotations allow us to separate breakdowns in spatial inference from extraction failures and language model formatting idiosyncrasies.
+Each question in FloorplanQA is paired with a reference answer computed directly from the symbolic layout, enabling fully automated and deterministic evaluation of model outputs.
+Depending on the response type, correctness is assessed using numeric comparison with fixed tolerances, string matching, or geometric validation checks.
+For numerical questions (e.g., distances, areas, angles), predictions are accepted if they fall within a relative error threshold.
+Sequence outputs are evaluated for both format and semantic correctness.
+Rather than requiring exact coordinate matches, we validate paths using geometric plausibility conditions of collision avoidance and sufficient proximity to a ground-truth trajectory.
+Deviation thresholds are set conservatively to tolerate minor geometric variations without crediting qualitatively wrong routes.
+To differentiate reasoning failures from extraction or formatting issues, we apply a regex-based parsing pipeline, covering a fixed set of expected answer patterns (e.g., ‘*Final answer*’ tokens in lower case or with surrounding symbols).
+If no answer is produced, or if the extracted content does not match a valid format, we count the response as an error (which is also considered incorrect).
+We provide a detailed breakdown on accuracy in Appendix E.
+In our evaluations, the proportion of invalidly formatted answers is below 1%.
+We also explicitly track cases where no answer is returned due to truncation (API: stop_reason = Token Limit), a failure mode that disproportionately affects reasoning-heavy models.
+To account for this, we report in Appendix I both the percentage of responses truncated by token limits and an adjusted accuracy computed only over valid (non-truncated) answers.
+These adjusted accuracies can be interpreted as an approximate upper bound on overall dataset performance, since with unlimited token budgets models could in principle reach that quality.
 3.6 Model Inference Setup
 All models are queried using standard chat-based completion APIs, with prompts constructed as described above.
-We evaluate a mixture of open-weight and proprietary language models across three resource tiers, distinguished by their token limits and relative inference cost.
-Models in the “reasoning” tier are allocated up to 8192 tokens per completion, enabling them to process longer layout descriptions and multi-step CoT outputs.
-Other models operate within a 4096-token context window and are evaluated only on the non-reasoning subset of questions.
-No model receives fine-tuning or prompt adaptation specific to PlanQA.
+We evaluate both large and mid-size models, including reasoning and standard variants.
+Large reasoning-oriented and standard models are allocated up to 12,288 tokens per completion, enabling them to process long layout descriptions and produce multi-step outputs.
+Mid-size models, including the GPT family variants optimized for speed and Qwen3-30B, are limited to 8,192 tokens.
+These budgets reflect our observations that larger models generate longer intermediate justifications and thus consume more tokens.
+GPT-5 is evaluated under a distinct configuration: its reasoning intensity can’t be disabled, so we run it with reasoning and verbosity set to “low,” with a maximum output length of 4,096 tokens, while GPT-5-mini is run with reasoning and verbosity set to “medium,” with a maximum output length of 8,192 tokens.
+No model receives fine-tuning or prompt adaptation specific to FloorplanQA.
 All prompts are zero-shot, with the system message and output formatting constraints held fixed.
-For each model, we use the default inference configuration provided by its vendor, including temperature settings.
-In practice, these are near-deterministic across runs and sufficiently stable for head-to-head comparison.
-In cases where a model produces variable-length completions or truncates early, we retain the first parseable instance of the “Final Answer:” line, if present.
-Each model is evaluated over a batch of 600 layouts per room type, with one question per layout.
-Requests are grouped into new-line-delimited JSON (JSONL) payloads for efficient submission via batched APIs.
-Layout serialization, prompt insertion, and output parsing are fully automated, with no manual curation or intervention.
-We include both high-capacity generalist models and compact instruction-tuned variants to measure the degree to which spatial reasoning emerges with scale and supervision alone.
-All models are evaluated on identical input distributions and scoring criteria, enabling cross-system comparisons that are both cost-aware and architecture-agnostic.
+For each model, we use the default inference configuration provided by its vendor, with temperature set to 0.
+The only exception is GPT-5, for which the temperature cannot be modified and defaults to 1.
+Each model is evaluated over 1800 generated layouts and 200 layouts from HSSD, with one question from each type posed per layout.
+Evaluation is fully automated, from layout serialization and prompt insertion to parsing, with no manual curation.
+All models are evaluated on identical input distributions and scoring criteria, enabling cross-system comparisons that are architecture-agnostic and directly comparable.
 4 Results
-We evaluated the performance of the model on a dataset of 1,800 layouts, consisting of 600 kitchens, 600 living rooms and 600 bedrooms.
-Each layout is paired with one question sampled from a pool of 10 parameterized templates, filtered by room applicability.
-This yields 18,000 layout–question pairs.
+We evaluated the performance of the model on a dataset of 2,000 layouts, consisting of 600 kitchens, 600 living rooms, 600 bedrooms, and 200 additional layouts from HSSD.
+Each layout is paired with one question sampled from a pool of 8 parameterized templates, filtered by room applicability.
+This yields 16,000 layout–question pairs.
 The full taxonomy is shown in Table 2.
-Eleven language models are evaluated, covering a wide range of parameter scales, architectures, and training regimes.
-The reasoning-oriented models include Qwen3-32B Yang et al.
-(2025) and DeepSeek-R1 DeepSeek-AI (2025).
-The large general-purpose models include DeepSeek-V3 DeepSeek-AI (2024), LLaMA 3.3-70B-Instruct, Gemma 2-27B-it Riviere et al.
-(2024), phi-4 Abdin et al.
-(2024b), and GPT-4.1 OpenAI et al.
-(2024).
-The smaller models include LLaMA 3.1-8B-Instruct Grattafiori et al.
-(2024), Gemma 2-9B-it Riviere et al.
-(2024), Phi-3.5-mini Abdin et al.
-(2024a), and GPT-4o-mini.
+We evaluate fifteen models, spanning a wide range of parameter scales, architectures, and training regimes.
+The reasoning-oriented models include GPT-5 (OpenAI, 2025a), GPT-OSS-120B (OpenAI, 2025b), DeepSeek-R1-0528 (DeepSeek-AI, 2025), GPT-5-mini, Gemini Flash 2.5 (Google DeepMind, 2025), GPT-OSS-20B, and Qwen3-30B-A3B-Thinking-2507 (Yang et al., 2025).
+The general-purpose (standard) models include Claude Sonnet 4 (Anthropic, 2025), GPT-4.1 (OpenAI et al., 2024), Moonshot Kimi-K2-Instruct (Kimi Team, 2025), Qwen3-Coder-480B-A35B-Instruct, Qwen3-235B-A22B-Instruct-2507, GPT-4.1-mini, Qwen3-30B-A3B-Instruct-2507, and Devstral-Small-2505 (MistralAI, 2025).
 All models are evaluated in identical zero-shot conditions using standardized prompts and serialized layout inputs, as described in Section 3.
-The complete results disaggregated by question, room type, and model are included in the appendix.
+Complete results disaggregated by question type, room type, and model are provided in Appendix E.
 4.1 Quantitative Results
-We begin by aggregating accuracy across cognitive targets as defined in Table 4.
-Figure 2 shows a heatmap of mean accuracy by model and target category.
-Metric reasoning (M) and visual logic (V) show the highest cross-model scores, while Topology (T), rule-based validation (R), and geometric planning (G) are more difficult across the board.
-DeepSeek-R1 (reasoning), GPT-4.1 (large), and GPT-4o-mini (compact, tuned) consistently lead the performance in the harder categories, although even large models struggle in certain occlusion and clearance tasks.
-Surprisingly, GPT-4.1’s performance on vector-direction (V) questions nearly matches DeepSeek-R1 and GPT-4o-mini—despite those models’ explicit chain-of-thought optimization—highlighting an unexpected parity in occlusion reasoning.
-Figure 2: Accuracy by model and cognitive target.
-Each cell shows average accuracy across questions associated with a given skill cluster.
-Cognitive targets: M = Metric computation, T = Topology, G = Geometric planning, V = Visual logic / occlusion, R = Rule-based validation.
-Performance improves with model scale, but not uniformly across tasks.
-As shown in Figure 3, accuracy improves steeply up to  30B parameters, suggesting that scale alone cannot fully close the gap on complex spatial tasks.
-Notably, gains largely plateau beyond 30 B parameters, indicating diminishing returns for raw scale on planning and constraint problems.
-Figure 3: Accuracy vs.
-model size.
-Models are arranged from the smallest to largest in terms of trainable parameters.
-While larger models perform better overall, instruction tuning and task specialization also contribute significantly.
-We also compare accuracy across room types (Figure 4).
-Kitchens exhibit the highest overall accuracy, likely because their regular counter-and-appliance grids reduce occlusion and simplify clearance reasoning.
-In contrast, living rooms and bedrooms generally yield lower performance across models, due to denser layouts, complex furniture adjacencies, and higher occlusion likelihood, with frequent object groupings such as televisions on stands, sofas and coffee tables on rugs, or lamps on nightstands.
-These room-type gaps suggest that layouts with strong grid structure provide easier inductive priors for LLMs than organic, cluttered arrangements.
-Figure 4: Accuracy by question and room type.
-Kitchens show the highest average accuracy, likely due to their regular, grid-like counter layouts and lower occlusion.
-The inability of most models to estimate free space in living rooms and bedrooms may be caused by more clutter, or the presence of objects like area rugs which the model was instructed to ignores for occlusion purposes.
+We begin by aggregating accuracy across models and question types for both reasoning and general model families.
+Figure 2 summarizes accuracy: the top row shows general models, and the bottom row shows reasoning models.
+In each figure, the left panel summarizes accuracy by model, and the right panel summarizes accuracy by question across room types.
+Kitchens lead across models because overlaps are rare, so most queries are straightforward.
+Scores on HSSD tend to lag behind the other room types; irregular, non-axis-aligned geometry and denser overlap make these layouts more demanding, see Appendix D for detailed statistics.
+Bedrooms and living rooms are mid-tier and nearly equal, lying between kitchens and HSSD.
+Comparing model families, general models struggle when many overlaps must be merged; they often treat object areas independently (no union), which hurts Free Space and Max Box and carries over into path planning.
+Reasoning models handle those cases with unions and rotations better, so they show gains on Free Space and Max Box.
+A practical limitation remains: Gemini Flash 2.5 and DeepSeek-R1 hit token limits on larger layouts, which drops scores, especially in HSSD.
+Task difficulty follows a consistent pattern in terms of accuracy.
+Metric-category questions, such as Pair Distance and View Angle, achieve the highest accuracies.
+Repositioning, Visibility, and Placement yield mid-range accuracies.
+Max Box and Free Space benefit most from reasoning-oriented models, yet their accuracies remain low and are comparable in difficulty to Shortest Path.
+Accuracy decreases as object count and overlap density increase, most notably on HSSD layouts.
+Detailed analysis on each question category is provided in Appendix G
+Figure 2: Accuracy of general (top) and reasoning (bottom) models.
+Left: by model.
+Right: by question.
+Each column corresponds to a specific room type represented in our dataset: Kitchens, Living Rooms, Bedrooms (synthetic subsets), plus HSSD layouts.
 4.2 Ablation
@@
 This substitution preserves all geometric and object-level content while modifying only the syntax and structural serialization.
-To reduce evaluation cost, we focused on the most variance-sensitive question categories in the kitchen domain: Repositioning, Vector Direction, and Free Space.
-These were selected based on standard deviation in accuracy across the full model set.
-We apply each ablation to three illustrative models spanning architecture classes and capacity tiers: GPT-4.1 (closed large), GPT-4o-mini (compact tuned), and DeepSeek-R1 (reasoning-optimized).
-As shown in Table 5, accuracy remains stable across formats for most question types, suggesting that these models encode layout semantics in a manner invariant to low-level representation details.
+The prompt level and semantic ablations are described in Appendix, Section F.
+To avoid recomputing the full suite, we focus on the most variance-sensitive question types for HSSD layouts, as shown in Appendix F (Figs.
+5 and 6): View Angle, Visibility, and one task from a different category—Repositioning (Dynamic).
+We apply each input-format ablation to three representative models: two reasoning models (GPT-OSS-120B, GPT-OSS-20B) and one large general model (Qwen3-235B-A22B) for the same subset of HSSD layouts.
+As shown in Table 3, accuracy is largely stable in both formats for most categories; the GPT-OSS models change minimally, while Qwen shows modest fluctuations.
+This suggests that these models encode layout semantics in a manner that is relatively invariant to low-level representation details.
 Table 3: Accuracy using JSON vs XML layout encoding. Each cell shows performance on the original JSON representation and its equivalent XML rendering.
-Model Repositioning View Angle Obstruction
-GPT-4.1 77.0 → 76.5 92.2 → 93.7 75.5 → 89.6
-GPT-4o-mini 33.8 → 29.6 42.4 → 42.2 12.2 → 15.8
-DeepSeek-R1 83.6 → 93.7 74.7 → 74.3 95.7 → 96.1
+Model Repositioning View Angle Visibility
+GPT-OSS-120B 60.5 → 59.0 74.0 → 74.0 70.0 → 72.0
+GPT-OSS-20B 40.0 → 39.0 37.5 → 38.5 45.5 → 43.5
+Qwen3-235B-A22B 39.0 → 42.0 50.5 → 54.0 70.5 → 65.5
 5 Conclusion
-We have presented PlanQA, a benchmark that evaluates spatial reasoning over symbolic 2D layouts by leveraging object-centric encodings (JSON, XML) similar to those used in architecture and robotics.
-Through a systematic evaluation of 11 language models across five skill clusters: metric computation, topology, geometric planning, visual logic, and rule-based validation, we make critical findings with implications for the use of LLMs in geometric design tasks.
-Metric and basic visibility queries (distances, areas, line-of-sight) are handled reliably (>95% accuracy), whereas geometric planning, topological relations, and constraint verification consistently fall below 50% accuracy.
-Performance gains level off beyond approximately 30 billion parameters, and models optimized for chain-of-thought show only modest improvements on the hardest tasks.
-Additionally, grid-like layouts (kitchens) yield higher accuracy than more varied configurations (living rooms, bedrooms), indicating that layout regularity influences model performance.
-These results reveal that current LLMs do not form sufficiently robust internal representations for complex spatial inference.
-Addressing this limitation could take two complementary directions.
-In the near term, integrating external geometric solvers or symbolic planning modules can compensate for the models’ weaknesses in collision avoidance and clearance reasoning.
-In the longer term, training regimes that incorporate explicit spatial constraints—through specialized objectives, data augmentation with irregular layouts, or constraint-violation examples—may be necessary to improve performance on design-oriented tasks.
+We introduced FloorplanQA, a benchmark for spatial reasoning over symbolic 2D layouts aligned with architecture and robotics practice.
+We evaluated 15 language models (8 general, 7 reasoning) on 2,000 layouts (1,800 synthetic; 200 semi–real HSSD) across tasks ranging from metric queries to visibility, placement, and shortest path with clearance.
+Empirically, metric and simple visibility queries are reliable; kitchens score highest because overlaps are rare.
+HSSD layouts are more demanding—irregular, non–axis-aligned shapes and dense overlaps expose weaknesses such as centroid miscalculation and missed unions.
+Reasoning models improve notably on Free Space and Max Box by handling overlaps and rotations more consistently, while general models often subtract object areas independently and fail under heavy overlap.
+Shortest Path is sufficiently challenging because it requires multiple correct steps (clearance buffering, collision checks, and path search), where errors compound.
+These results indicate that current LLMs lack sufficiently robust internal geometric representations for complex spatial inference.
+Two complementary directions follow.
+Near term: hybridize with external geometric solvers or symbolic planning modules—set operations (unions/differences), centroid via shoelace, clearance buffering, oriented rectangle search, and A* path planning—to compensate for the models’ weaknesses in collision avoidance and clearance reasoning.
+Longer term: train with explicit spatial constraints and harder distributions (irregular, overlap–heavy layouts), and include constraint–violation exemplars and geometry-aware objectives so models learn to maintain coherence under rotation, clearance, and union operations in design-oriented tasks.
 Acknowledgments
-The work is supported by funding from KAUST, Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program
-Appendix A Technical Appendices and Supplementary Material for PlanQA
-This technical appendix includes an overview of the dataset generation process, followed by full evaluation results on our generated QA pairs, along with aggregated metrics and ablation studies.
-It also presents an extended taxonomy of spatial reasoning tasks, explanations of selected failure cases, an expanded analysis of vision-language models (VLMs), and examples of prompts used for both generation and questioning.
-The full dataset and all code used for generation, evaluation, and visualization are provided in the supplementary materials.
+The work is supported by funding from King Abdullah University of Science and Technology (KAUST)—Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program
+Appendix A Technical Appendix and Supplementary Material for FloorplanQA
+This appendix details the dataset generation pipeline, dataset statistics, and the full evaluation of FloorplanQA across reasoning and general models.
+It also includes ablation studies, analysis of failure cases, and the exact prompts used for data generation and question asking.
+All artifacts (code, generation prompts, evaluation scripts, and visualization utilities) are included with the submission as supplementary material.
 Appendix B Details on Synthetic Data Generation
-Room layouts were generated using Gemini 2.5 Pro, specifically fine-tuned for spatial reasoning and bounding-box tasks Team et al.
-(2025).
-Below, we describe the detailed procedure and constraints.
+Room layouts were generated using Gemini 2.5 Pro, specifically fine-tuned for spatial reasoning and bounding-box tasks.
+In the following sections, we describe the detailed procedure and its constraints.
 All generation scripts, prompts, constraint-checking code, and random seeds are included in the code for full reproducibility.
 B.1 Room Shape Generation
-We generated 600 layouts each for kitchens, bedrooms, and living rooms.
-These layouts include a variety of geometries and sizes, with clearly defined room structures containing windows, doors, and corner cutouts where applicable.
+We generated 600 layouts for each of the kitchens, bedrooms, and living rooms.
+These layouts feature a range of geometries and sizes, with clearly defined room structures that incorporate windows, doors, and corner cutouts where applicable.
 All layouts follow specifications tailored to each room type.
 Each layout falls into one of three shape categories: rectangular, L-shaped, or open.
-Size categories were defined individually for each room type and were incorporated into the generation prompts based on common assumptions about typical room dimensions.
+Size categories were defined individually for each room type and were incorporated into the generation prompts based on standard assumptions about typical room dimensions.
 The distribution of room shapes and sizes is shown in Table 4, and example layout types are illustrated in Fig. 1.
 Table 4: Shape and size distribution across generated layouts for each room type.
-Room Type Shape Distribution (Rect / L-shaped / Open) Size Categories (in m2) (Small / Medium / Large)
-Kitchen 40% / 40% / 20% $\leq$7 / 7–18 / $>$18
-Bedroom 50% / 30% / 20% 8–12 / 12–18 / $>$18
+Room Type
+Shape Distribution
+(Rect / L-shaped / Open)
+Size Categories (in m2)
+(Small / Medium / Large)
+Kitchen 40% / 40% / 20% $\leq$7 / 7–18 / >>18
+Bedroom 50% / 30% / 20% 8–12 / 12–18 / >>18
 Living Room 40% / 40% / 20% 20 / 22 / 24
 Geometric and Structural Constraints:
-Room geometries were procedurally varied using prompt-based guidance to produce rectangular, L-shaped, and open-plan configurations.
+The geometry of the room was procedurally varied using prompt-based guidance to produce rectangular, L-shaped, and open-plan configurations.
 The following structural properties were described in the prompts but not explicitly enforced during layout generation:
-  • L-shaped rooms were described as rectangular spaces with a square cutout from one corner. To maintain usable proportions, each leg of the L shape was suggested to be at least 1.5,m wide and deep.
-  • Open-plan rooms, apart from living rooms, were prompted without doors and with one full wall removed. This was intended to vary room shapes and simplify layout generation.
-  • Doors were described with widths between 0.8,m and 1.0,m, selected randomly. Windows were chosen from a fixed set of widths: 0.6,m, 0.75,m, 0.9,m, 1.2,m, and 1.5,m.
-  • Prompts included guidance to place all elements such as doors, windows, and cutouts entirely within room boundaries.
+  • L-shaped rooms were described as rectangular spaces with a square cutout in one corner. To maintain usable proportions, each leg of the L shape was suggested to be at least 1.5 m wide and deep.
+  • Open-plan rooms, apart from living rooms, were prompted without doors and with one whole wall removed. This was intended to vary the room’s shape and simplify the layout generation process.
+  • Doors were described with widths between 0.8 m and 1.0 m, randomly selected. The windows were chosen from a fixed set of widths: 0.6 m, 0.75 m, 0.9 m, 1.2 m, and 1.5 m.
+  • The prompts included instructions on placing all elements, such as doors, windows, and cutouts, entirely within the room boundaries.
 Window Placement:
@@
   • The total window length was set to exceed 15 percent of the room’s floor area. It was used as a simple proxy to ensure visible window openings and visual balance along the walls.
-  • Windows on the same wall were the same size to support visual balance. Small, isolated windows were not used.
+  • The windows on the same wall were the same size to support visual balance. Small, isolated windows were not used.
   • Long windows, over 1.5 m, were split into segments with 0.05 to 0.15 m gaps for a more modular appearance.
-  • Windows were not placed opposite each other or on the same wall as a door, as such arrangements are less common and were not emphasized in the prompt design.
+  • The windows were not placed opposite each other or on the same wall as a door, as such arrangements are less common and were not emphasized in the prompt design.
 B.2 Furniture and Appliance Placement
@@
 B.3 Layout Selection Criteria
-Approximately one-third of the initially generated room layouts were filtered-out using a set of geometric and functional constraints.
+Approximately one-third of the initially generated room layouts were filtered out using a set of geometric and functional constraints.
 These filters were designed to ensure realistic object placement, functional usability, and architectural plausibility.
@@
   • Non-blocking door clearance: Doors have physical thickness and are defined by bounding boxes $(x_{1}^{d},y_{1}^{d},x_{2}^{d},y_{2}^{d})$.
-A clearance zone of $door\_length$ meters is required in front of the door to ensure swing space and accessibility.
+A clearance zone of door_lengthdoor\_length meters is required in front of the door to ensure swing space and accessibility.
 The position of this zone depends on which wall the door is attached to.
@@
 They are not universal requirements for real layouts, nor a complete set of constraints, but aim to avoid frequent sources of implausibility in generated layouts.
-B.4 Room Statistics
-Table 7 summarizes key statistics for the generated layouts across room types, including average object count, overlap frequency, and spatial density.
-Kitchens typically contain fewer objects but have higher density due to their compact size.
-In contrast, bedrooms and living rooms host more objects with greater overlap, reflecting more open layouts and flexible furnishing.
-This structural simplicity, with fewer objects, minimal overlap, and higher density, makes kitchens the simplest environment for spatial reasoning tasks.
-Due to the limited amount of free space, some complex questions, such as those involving paths, have fewer possible answers.
-As a result, models tend to perform better on kitchen-related questions.
+Appendix C Details on HSSD layouts selection
+We curate a subset of HSSD layouts to ensure clean geometry and unambiguous supervision for spatial reasoning tasks.
+Starting from the raw scenes, we generate a 2D floor-plan projection and apply filtering and normalization steps (e.g., geometry cleanup and category unification).
+Object filtering.
+To reduce visual clutter and retain only objects essential for spatial reasoning, we exclude purely decorative or small accessory categories.
+The banned labels are: accessory, air conditioner, artwork, blanket, boots, bottle, bowl, box, book, brush, candle, cushion, decor, dog, fan, flower, frame, guitar, herb, hook, jar, light, lightbulb, orchids, pendant, pendulum, pet, some plants, poster, pot, shoes, socket, switch, vase, wreath.
+We also drop subcomponents that fragment footprints without changing free space (e.g., chair/table legs; bed/armchair frames), remove partial cabinet-door leaves, and consolidate multiple near-duplicate small items by keeping a single representative instance.
+The resulting scenes preserve the functional layout while simplifying geometry.
+In Figure 3, the left panel shows the layout immediately after 2D projection; the right panel shows the simplified layout after filtering and normalization.
+Projection and cleanup.
+During 2D projection, we correct mislabeled windows and doors, snap nearly collinear edges, and resolve self-intersections.
+To smooth rectilinear artifacts and irregular edges in box-like footprints, we apply alpha-shape–based convex hulls, then recompute centroids and areas on the cleaned geometry.
+Figure 3: HSSD layout selection: comparison of the original (left) and simplified (right) 2D layouts.
+The simplified version removes redundant or overlapping objects and cleans geometry to produce well-structured input for spatial reasoning tasks.
+Appendix D Room Statistics
+Table 5 reports summary statistics for the layouts.
+Kitchens are typically the smallest spaces, with relatively few objects and overlaps, but a high density due to their compact geometry.
+Living rooms are the largest, with slightly more objects overall but lower density, reflecting their open layout.
+Bedrooms fall in between, with similar object counts to kitchens but more frequent overlaps.
+The HSSD layouts are comparable in scale to bedrooms and living rooms in terms of area and object count, but differ in structure: objects are represented with detailed, non-axis-aligned polygons, resulting in a significantly higher vertex count.
+They also exhibit more overlaps than the generated layouts, reflecting their closer alignment with human-authored floorplans.
+In other respects, however, the distributions remain broadly consistent.
 Table 5: Average layout statistics by room type.
-Metric Kitchen Bedroom Living Room
-Avg. # of Objects 13.30 14.24 15.29
-Avg. # of Overlaps 0.52 6.27 6.26
-Avg. Object Density 1.15 0.84 0.71
-Appendix C Full Breakdown by Room Type, Question Category, and Model
-Table 8 reports per-model accuracy separated by room type (kitchen, living room, bedroom) and question category.
-This is the full unaggregated results matrix underlying the analyses presented in Section 4.
-The table includes all eleven evaluated models and eleven question types.
-Path is assessed using two complementary metrics: validity and Fréchet Distance.
+Metric Kitchen Bedroom Living Room HSSD
+Avg. Area (m2) 12.00 17.76 20.75 17.95
+Avg. # of Objects 10.35 10.76 11.69 12.20
+Avg. # of Overlaps 0.52 1.82 1.52 4.39
+Avg. Object Density 0.95 0.66 0.57 0.83
+Avg. Vertices per Layout 41.39 43.03 46.77 152.29
+To further illustrate these statistics, Figure 4 shows the distribution of object counts across all layouts.
+The histograms confirm the averages reported in Table 5: kitchens are concentrated at lower counts, typically around 10 objects; bedrooms and living rooms exhibit broader distributions with slightly higher counts; and HSSD layouts overlap in range but extend to higher counts in the tail.
+Overall, the object distributions remain comparable across sources, with no extreme outliers.
+Figure 4: Distribution of object counts across kitchens, bedrooms, living rooms, and HSSD layouts.
+Appendix E Full Breakdown by Room Type, Question Type, and Model
+E.1 Detailed Accuracy by full dataset
+Tables 6 and 7 report the full per-model accuracy matrix for the base and reasoning model groups, respectively.
+Each table covers nine question categories across four room types, resulting in a total of 36 rows.
+The path task is assessed using two complementary metrics: validity and Fréchet Distance.
 Together, these question types span a range of challenges, covering both reasoning-heavy and functionally grounded tasks.
-The table presents accuracy for all eleven models across eleven question types, grouped by room type.
-Each room type–kitchen, bedroom, and living room–includes results for all question types, totaling 33 rows.
-Each value reflects model accuracy over a fixed set of 600 questions, corresponding to 600 layouts per room type.
-For every combination of question type and room type, the same set of 600 questions is used across all models to ensure fair comparison.
-Table 8: Question-level accuracy by category for each model across room types.
-Models (left to right): Qwen3-32B, DeepSeek-R1, DeepSeek-V3, LLaMA 3.3-70B-Instruct, Gemma 2-27B-it, Phi-4, GPT-4.1, LLaMA 3.1-8B-Instruct, Gemma 2-9B-it, Phi 3.5-mini, GPT-4o-mini.
-Note: High truncation rate due to token limits.
-Reasoning Big Models Small Models
-N 32.8B 671B 671B 70B 27.2B 14B N/A 8B 9.2B 3.8B N/A
-Temp. 0.6 0.6 0.3 0.0 0.5 0.0 1.0 0.6 0.5 0.7 1.0
-Type Category [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-K Distance 97.2 99.5 98.5 97.2 94.7 90.7 100 82.3 74.2 35.3 97.3
-Area (counters) 85.3 99.5 95.7 81.7 46.5 79.2 98.5 30.8 37.5 9.8 80.3
-Free Space 83.7 88.0 52.5 37.0 23.5 42.0 83.5 9.5 13.5 6.3 20.3
-View Angle 74.0 58.2 69.2 75.2 12.7 52.8 86.3 11.5 9.2 8.2 42.2
-Repositioning 91.5 96.3 69.8 48.7 14.8 22.7 85.8 3.7 6.2 11.5 40.7
-Max Box 33.3 30.7 9.2 8.8 3.0 3.0 57.2 1.7 1.2 1.7 3.7
-Fit/Placement 92.8 92.5 71.0 66.2 72.0 82.0 89.8 68.7 71.8 68.8 70.7
-Path (Valid) 13.2 6.3 39.3 30.3 23.2 26.8 73.0 10.5 17.5 14.8 35.7
-Path (Fréchet) 15.0 6.3 32.8 30.2 26.0 26.0 55.8 10.2 19.0 9.3 36.7
-Missing Object 87.3 88.7 44.3 56.2 52.3 52.2 79.8 27.5 39.7 14.3 58.0
-Obstruction 84.0 95.2 32.7 6.0 3.3 9.7 93.5 1.2 2.7 11.3 14.2
-L Distance 98.7 99.8 99.5 98.2 96.3 98.8 99.8 87.0 81.8 58.8 98.5
-Area (sitting) 96.7 99.5 84.5 98.0 83.8 88.7 99.5 32.5 41.3 12.8 85.7
-Free Space 0.3 4.3 1.0 0.3 3.7 0.3 5.0 0.7 3.2 1.5 1.7
-View Angle 81.5 86.0 70.0 76.8 14.3 50.3 96.0 14.8 8.3 10.5 42.2
-Repositioning 80.5 93.0 45.0 33.5 12.8 19.2 71.5 6.5 5.7 6.3 29.3
-Max Box 1.5 0.8 1.0 3.0 2.0 2.8 6.5 0.8 1.5 1.0 2.3
-Fit/Placement 90.7 91.2 71.5 80.0 83.7 87.7 91.8 75.0 75.5 72.8 72.7
-Path (Valid) 10.0 14.7 33.0 26.7 30.5 26.7 53.2 7.3 21.7 16.7 33.5
-Path (Fréchet) 13.2 14.2 23.7 17.8 19.2 14.7 48.0 2.5 14.2 6.8 25.2
-Missing Object 73.0 76.2 51.0 49.3 29.5 36.0 65.5 9.7 28.3 11.7 32.3
-Obstruction 80.7 96.5 24.3 7.3 3.8 9.3 84.7 2.5 5.2 4.7 11.7
-B Distance 98.7 99.8 99.5 98.2 96.3 98.8 99.8 87.0 81.8 58.8 98.5
-Area (storage) 98.7 99.8 94.0 97.0 86.3 88.3 99.3 30.3 66.3 47.7 88.2
-Free Space 1.7 5.8 1.2 0.3 2.5 1.2 2.8 1.8 1.0 1.0 1.2
-View Angle 76.0 79.8 70.0 78.3 10.8 57.0 94.2 15.3 10.2 7.7 43.0
-Repositioning 78.7 94.3 53.8 36.0 11.2 15.7 73.7 4.3 9.0 6.5 31.5
-Max Box 0.7 1.0 2.0 1.8 2.0 2.5 7.2 1.0 0.7 1.0 2.0
-Fit/Placement 86.3 86.3 65.8 66.7 66.5 73.0 82.0 66.0 63.8 64.7 66.2
-Path (Valid) 15.5 20.8 49.7 40.7 38.5 35.8 67.8 11.3 34.5 20.5 48.8
-Path (Fréchet) 15.3 21.5 30.3 30.3 27.0 19.0 49.5 3.7 19.7 8.5 36.3
-Missing Object 64.3 65.2 39.8 40.3 33.0 25.2 61.0 14.8 19.7 8.3 34.5
-Obstruction 87.2 95.7 32.0 4.3 1.3 5.8 89.5 2.2 5.3 9.3 12.2
-Appendix D Additional Ablations and Analysis
-To further verify the robustness and generalizability of our evaluation setup, we conducted several supplementary analyses beyond the primary results in Table 8.
-Mean and Variability Across Room Types
-Figure 6 shows radar plots summarizing model performance across question types, averaged by room type.
-Kitchens, living rooms, and bedrooms demonstrate similar high-level trends: models are generally most accurate on questions involving spatial metrics such as Distance, Area, and Fit/Placement, while struggling more with abstract or multi-object tasks such as Max Box and Path (Valid).
-Figure 6 illustrates the standard deviation across models per question category.
-Notably, Repositioning, View Angle, Area, and Obstruction show the highest variability across models, particularly in bedrooms and living rooms.
-Free Space, in contrast, exhibits high variance only in kitchens, highlighting that performance inconsistency is not uniform across room types.
-Figures 8 and 8 provide complementary radar plots organized by model, highlighting differences in room-specific performance.
-Kitchens consistently exhibit slightly higher mean accuracy across models, particularly for larger models like GPT-4.1 and DeepSeek-R1.
-Variance across models is relatively uniform, though larger models tend to show reduced variance overall.
-Figure 5: Mean accuracy by question type across room categories. Each axis corresponds to a specific spatial reasoning skill (e.g., distance, placement, path validity).
-Figure 6: Standard deviation in model accuracy across question types and room categories. Higher values indicate greater variability across models within a room type.
-Figure 7: Mean accuracy across models by room category. Shaded regions reflect performance in bedrooms, kitchens, and living rooms, highlighting which models generalize best across spatial tasks.
-Figure 8: Standard deviation of model performance across room types. Captures how consistent each model is across different spatial environments. Narrower shapes indicate more uniform behavior.
+The base group includes eight general-purpose models, while the reasoning group includes seven models with explicit reasoning capabilities.
+For every model, question type accuracy is computed over a fixed set of 600 synthetic layouts (Kitchens, Living Rooms, Bedrooms) and 200 HSSD layouts, ensuring that results are directly comparable across models.
+E.2 Token-Limit Analysis and Valid-Only Accuracy
+In addition to overall accuracy, we analyze two complementary aspects of model performance.
+First, Tables 10 and 12 quantify the fraction of responses that were terminated due to the TOKEN LIMIT stop reason.
+This failure mode is particularly relevant for reasoning-oriented models, which often generate longer chain-of-thought outputs.
+Second, Tables 11 and 13 report accuracy over completed answers only, excluding truncated outputs (e.g., token-limit terminations).
+This metric isolates models’ reasoning performance on successfully produced, well-formed responses.
+Together, these analyses complement the full accuracy tables by disentangling reasoning failures from generation truncation and formatting issues.
+Table 6: Question-level accuracy on full dataset by room type for standard models.
+Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
+Pair distance K 99.8 96.5 95.7 96.8 99.2 90.7 89.5 58.8
+LR 99.5 95.0 94.2 96.8 99.7 88.5 88.8 62.3
+B 99.7 96.3 93.3 96.5 99.5 87.8 85.2 60.0
+HSSD 88.0 56.0 75.5 66.0 67.0 37.0 44.5 56.0
+Placement K 87.8 78.0 82.2 80.3 90.2 86.5 85.5 74.2
+LR 80.5 69.0 73.8 83.2 89.2 75.2 85.7 71.8
+B 68.8 59.8 67.5 70.8 76.7 68.7 77.0 56.3
+HSSD 72.0 64.5 73.0 70.0 82.0 76.5 71.5 54.5
+Reposi- tioning K 73.8 63.8 48.7 45.3 83.5 66.3 73.7 14.8
+LR 79.3 60.3 56.7 64.5 91.3 76.8 72.2 25.0
+B 71.0 55.5 48.3 59.5 79.0 72.5 70.0 21.2
+HSSD 42.0 47.0 28.0 34.0 39.0 40.5 33.0 10.0
+Free space K 97.8 93.2 83.0 84.2 95.0 95.2 83.8 66.2
+LR 0.2 14.2 2.8 1.8 3.5 1.2 0.5 2.7
+B 2.7 31.2 1.0 1.3 8.8 0.8 1.0 0.7
+HSSD 35.0 16.0 24.0 22.0 17.0 15.5 7.5 6.5
+Visibility K 87.7 63.2 54.5 67.0 98.3 90.2 91.5 18.5
+LR 81.7 52.7 43.3 52.2 98.3 86.8 88.7 10.0
+B 74.8 57.5 41.0 54.0 96.7 86.3 86.3 10.8
+HSSD 46.5 20.0 22.5 26.5 70.5 52.0 45.5 9.0
+View angle K 92.0 95.3 69.8 78.8 97.0 95.8 74.7 49.7
+LR 87.7 93.2 59.7 75.3 94.0 93.0 72.2 40.5
+B 88.0 90.2 60.3 72.8 95.0 91.2 76.8 35.8
+HSSD 67.5 55.0 28.5 46.5 50.5 46.0 34.5 30.0
+Max box K 47.2 31.8 32.0 26.8 65.5 27.5 26.5 4.5
+LR 7.8 7.0 5.0 5.7 22.3  4.5 8.8 0.5
+B 5.8 6.8 7.3 5.0 29.2 4.8 7.0 1.7
+HSSD 5.0 7.5 2.5 2.0 11.5 4.5 3.0 1.0
+Shortest path (valid) K 59.2 61.7 52.7 39.7 45.2 55.3 21.8 34.2
+LR 53.0 51.3 42.8 37.3 44.8 47.2 20.3 30.3
+B 48.5 52.2 40.7 34.7 44.2 45.8 18.5 33.5
+HSSD 28.5 25.0 18.5 18.0 26.0 15.0 5.5 10.5
+Shortest path (Fréchet) K 45.3 56.8 51.3 28.2 44.2 39.5 17.8 36.2
+LR 24.7 42.5 27.3 12.3 34.7 26.5 9.2 15.5
+B 23.0 42.2 27.7 14.2 38.0 30.5 8.2 18.3
+HSSD 15.5 22.5 18.0 8.0 20.0 12.5 2.5 11.5
+Table 7: Question-level accuracy on full dataset by room type for reasoning models.
+Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
+Pair distance K 99.8 99.3 98.0 96.3 100.0 94.2 97.7
+LR 98.8 99.3 99.0 96.0 99.7 93.5 97.5
+B 98.3 99.5 96.8 95.5 99.7 93.8 95.3
+HSSD 69.0 78.5 25.5 12.5 32.5 40.5 18.0
+Placement K 84.7 92.0 89.0 59.7 90.8 85.7 68.2
+LR 75.5 89.0 82.2 53.3 86.3 78.5 46.5
+B 61.2 83.5 72.0 35.8 81.2 62.0 34.5
+HSSD 70.0 85.0 79.0 16.0 75.5 74.5 28.5
+Reposi- tioning K 83.0 85.5 79.2 90.5 84.5 70.8 61.5
+LR 85.5 89.8 86.2 91.2 92.8 87.8 77.0
+B 77.8 83.3 83.3 85.2 84.3 78.0 69.2
+HSSD 49.5 60.5 47.5 18.5 53.5 40.0 27.0
+Free Space K 82.5 99.0 93.0 93.3 99.5 94.8 97.0
+LR 47.0 83.3 18.3 17.7 78.5 53.2 0.0
+B 50.5 87.5 34.8 33.3 82.2 74.0 1.2
+HSSD 19.5 31.0 6.5 1.0 5.0 9.0 1.0
+Visibility K 94.8 94.2 71.3 26.8 98.0 91.5 78.8
+LR 95.2 94.0 52.0 11.3 98.0 89.3 70.8
+B 94.2 92.5 53.5 11.2 95.5 89.2 64.2
+HSSD 57.0 70.0 10.0 0.5 39.0 45.5 3.0
+View Angle K 96.2 98.5 73.7 92.5 88.3 93.5 98.5
+LR 93.3 97.3 68.2 91.5 84.5 92.0 98.3
+B 95.2 98.2 75.2 93.8 86.8 91.3 98.3
+HSSD 59.5 74.0 13.5 20.0 25.5 37.5 26.0
+Max Box K 48.5 62.8 50.5 3.7 85.2 31.3 16.7
+LR 17.3 28.2 8.0 0.0 60.3 3.8 0.8
+B 13.0 30.3 11.0 0.0 61.2 5.8 0.5
+HSSD 5.0 9.5 2.5 0.0 17.0 0.5 0.0
+Shortest path (valid) K 64.7 64.2 12.3 3.2 52.3 43.0 14.2
+LR 21.2 66.0 12.5 1.5 53.7 37.7 16.7
+B 58.2 57.8 7.8 1.0 52.0 30.0 14.3
+HSSD 28.5 33.5 8.5 0.0 16.5 13.5 1.0
+Shortest path (Fréchet) K 56.3 55.5 11.5 3.2 50.5 42.2 14.0
+LR 12.3 40.5 9.3 1.3 47.5 28.5 16.3
+B 39.3 44.8 6.0 0.8 47.7 24.2 14.3
+HSSD 22.5 26.5 2.5 0.0 12.0 14.0 1.0
+Appendix F Additional Ablations and Analysis
+To further verify the robustness and generalizability of our evaluation setup, we conducted supplementary analyses beyond the main accuracy tables.
+General models.
+Figure 5 shows mean accuracy (top left) and variability (top right) across general models, as well as accuracy (bottom left) and variability (bottom right) by question type.
+Kitchens are consistently easier, while HSSD is the most challenging.
+Among models, Devstral Small performs noticeably worse, whereas Qwen3-30B achieves a level comparable to that of larger models.
+Across question types, Pair Distance and View Angle from the Metric category yield the highest accuracy, while more complex tasks such as Max Box and Shortest Path show lower scores and higher variance across rooms.
+Reasoning models.
+Figure 6 presents analogous plots for reasoning models.
+GPT-family models show stronger results overall.
+In contrast, DeepSeek-R1 and Gemini Flash 2.5 struggle with token limits, as these models tend to produce very long outputs according to Table 12.
+By question type, Repositioning and Placement are handled reliably, whereas Max Box and Shortest Path remain the most difficult as well, with high variance across rooms.
+Figure 5: Radar plots for general models, showing mean and standard deviation of accuracy across (top) models and (bottom) question type.
+Figure 6: Radar plots for reasoning models, showing mean and standard deviation of accuracy across (top) models and (bottom) question type.
 Prompt Sensitivity Ablation
 To assess model robustness to prompt variation, we conducted an ablation in which each question was regenerated using the same template but with alternate object references.
-For example, a question originally referring to a "sofa" might instead use a "bookshelf" in the regenerated version.
+For example, a question originally referring to a ”sofa” might instead use a ”bookshelf” in the regenerated version.
 This allowed us to evaluate whether model performance remains stable under changes in object content while preserving linguistic structure.
-As shown in Table 9, performance generally remained consistent, with only minor differences across models.
-Larger models such as GPT-4.1 were more stable, while smaller models like GPT-4o-mini showed slightly higher sensitivity.
-These results suggest that the evaluation framework is robust to this type of variation.
-Table 9: Accuracy under prompt variation. Each cell shows performance on the original prompt followed by a regenerated version with alternate object references.
-Model Repositioning View Angle Obstruction
-GPT-4.1 77.0 → 75.3 92.2 → 91.4 75.5 → 74.9
-GPT-4o-mini 33.8 → 32.9 42.4 → 46.7 12.2 → 14.1
-DeepSeek-R1 83.6 → 95.7 74.7 → 72.1 95.7 → 95.6
-Layout Perturbation Ablation
-To test model sensitivity to small physical changes, we introduced random spatial jitter of up to ±0.1 m to all object coordinates during evaluation.
-This simulated minor but plausible noise in the layout.
-After jittering, we recomputed the correct answers for all QA pairs based on the updated room configuration.
-The questions remained unchanged, but all evaluations used the perturbed layouts.
-The jitter was constrained to ensure that objects did not overlap, remained within room boundaries, and did not block doors.
-Jittering can affect spatial relationships such as clearances, alignments, and angles.
-For example, a path that was previously clear might become blocked, or an object may no longer fit within a gap.
-As shown in Table 10, model accuracy showed slight fluctuations in response to the spatial changes.
-Overall, model performance remained robust, indicating resilience to small, physically realistic layout variations.
-Table 10: Accuracy before and after layout jitter. Values shown for three spatially sensitive questions.
-Model Repositioning View Angle Obstruction
-GPT-4.1 77.0 → 72.5 92.2 → 87.3 75.5 → 80.2
-GPT-4o-mini 33.8 → 28.4 42.4 → 42.5 12.2 → 11.8
-DeepSeek-R1 83.6 → 89.8 74.7 → 70.3 95.7 → 97.1
-Appendix E Spatial Reasoning Question Suite
-To extend our taxonomy of question types introduced in Table 2, Table 11 details the specific evaluation procedures used to compute ground truth answers for each category.
-In addition to categorizing the task, the table shows how answers are derived and what types of spatial reasoning each task demands.
-This breakdown clarifies not only what is being tested, but how it is measured, supporting both reproducibility and deeper analysis of model performance in later sections.
-Table 11: Categories of spatial reasoning questions, answer-checking procedures, and diagnostic insights.
-Category Sample Question Answer Checking Diagnostic Insight
-Distance What is the Euclidean distance between the center of the door and the fridge?
-Compute centroids $(x,y)$ of both boxes; calculate Euclidean distance between them; compare with model output allowing a tolerance of $2\%$ of the ground-truth distance.
-Tests numeric computation, coordinate extraction, and arithmetic precision.
-Area What is the area of the kitchen counter?
-Multiply the width $\times$ height of each cabinet; sum all resulting areas; compare the total against the model output with a 2% tolerance.
-Probes multiplicative reasoning, aggregation, and unit consistency.
-Free Space What is the total unoccupied floor area in the bedroom?
-Merge overlapping object bounding boxes; sum the resulting areas; subtract from total room area; compare the unoccupied area with model output using a 2% tolerance.
-Measures the ability to aggregate and subtract multiple regions.
-View Angle What is the dot product between the facing vector of the sofa and the vector to the TV?
-Compute normalized facing vector of reference; get normalized vector to target center; compare dot product with model output using 2% tolerance.
-Evaluates directional reasoning, vector normalization, and spatial alignment understanding.
-Repositioning How far can the sofa move right before hitting a wall or another object?
-Simulate axis-aligned movement; measure distance to furthest valid non-overlapping position; compare with model output using 2% tolerance.
-Tests spatial simulation, collision detection, and directional constraint reasoning.
-Max Box What is the largest axis-aligned box you can place in the living room?
-Use brute-force to solve largest empty rectangle in free space; compare resulting dimensions with 2% tolerance.
-Probes continuous-space optimization and spatial coverage estimation.
-Fit/Placement Can a $2\,\mathrm{m}\times 3\,\mathrm{m}$ table fit in the room without overlap?
-Attempt to place a rectangular object in free space; ensure no intersection with existing bounding boxes.
-Compare boolean result with model prediction.
-Tests collision detection, spatial constraint enforcement, and free-space reasoning.
-Shortest Path What is the shortest walkable path from the fridge to the exit with $15\,\mathrm{cm}$ clearance?
-Offset free space by $15\,\mathrm{cm}$ for clearance; use A* algorithm on navigable grid to compute path; compute Fréchet distance (threshold $0.6\,\mathrm{m}$) to model path; check its validity.
-Tests path planning, clearance-aware modeling, and heuristic graph search.
-Missing Object Where should the missing armchair be placed in this layout?
-Check if predicted bounding box fits in free space without overlaps and satisfies clearance/accessibility constraints.
-Evaluates spatial reasoning, free-space utilization, and functionally appropriate object placement.
-Obstruction Which objects obstruct the line of sight between the bed and the window?
-Compute vector from target to reference center; check which bounding boxes intersect the segment (excluding endpoints).
-Tests geometric reasoning, line–rectangle intersection, and occlusion detection.
-Appendix F Failure Cases
+As shown in Table 9, accuracy under prompt regeneration is broadly stable; larger models change little, and smaller ones fluctuate modestly.
+We observe somewhat higher sensitivity for Repositioning: paraphrases can implicitly select different target objects or motion directions, occasionally introducing additional complexity or non-movable cases.
+Overall, the evaluation appears robust to prompt-level variations.
+Table 8: Accuracy under prompt variation. Each cell displays performance on the original prompt, followed by a regenerated version with alternative object references.
+Model Repositioning View Angle Visibility
+GPT-OSS-120B 60.5 → 59.0 74.0 → 80.0 70.0 → 77.0
+GPT-OSS-20B 40.0 → 50.0 37.5 → 35.5 45.5 → 50.0
+Qwen3-235B-A22B 39.0 → 49.0 50.5 → 48.0 70.5 → 74.0
+Semantic Ablation
+To evaluate the model’s reliance on semantic object labels rather than purely geometric reasoning, we conduct a semantic ablation experiment.
+In this setting, object identifiers in the scene are permuted (e.g., swapping bed and chair labels) while keeping all geometry unchanged.
+The regenerated prompts thus refer to the same physical configuration but with altered object semantics, for instance, a question originally phrased as “move the chair left” becomes “move the bed left,” even though the underlying geometry is identical.
+As summarized in Table 9, tasks that rely primarily on pure metric computation or topological cues, such as View Angle and Visibility, show minimal changes in accuracy, with only small fluctuations due to prompt phrasing.
+However, performance on the action-based Repositioning task drops sharply under semantic perturbation, indicating that the model partially grounds its reasoning in object semantics rather than spatial configuration alone.
+This suggests that large language models may entangle linguistic priors with geometric inference when tasks involve physical movement or interaction.
+Table 9: Accuracy under prompt variation. Each cell displays performance on the original prompt, followed by a regenerated version with alternative object references.
+Model Repositioning View Angle Visibility
+GPT-OSS-120B 60.5 → 40.0 74.0 → 73.0 70.0 → 77.0
+GPT-OSS-20B 40.0 → 28.0 37.5 → 35.5 45.5 → 44.5
+Qwen3-235B-A22B 39.0 → 37.0 50.5 → 43.0 70.5 → 74.0
+Appendix G Case Studies by Question Type
 To better understand the sources of model failure, we conducted a qualitative analysis of representative examples from the benchmark.
@@
 By examining both correct and incorrect outputs, we aim to identify common failure patterns and reasoning bottlenecks across different architectures.
-F.1 Max Box
-As illustrated in Figure 10, this example shows GPT-4.1 producing a suboptimal result in a spatial reasoning task.
-Instead of identifying the largest free region in the kitchen layout, the model placed a bounding box that extended along a single axis, starting from one edge, without fully exploring the layout.
-It selected an area from x=0 to 2.1 and y=1.1 to 1.8, yielding 1.47 m² of free space.
-A larger available region, shown in Figure 10, spans from (0.7, 0.7) to (2.1, 1.8), and offers 1.54 m².
-The model did not identify this larger region, which indicates that its search strategy was likely not exhaustive, even though it had access to the full structured input.
-The reasoning trace in Figure 11 confirms that it evaluated a set of candidate regions but ultimately settled on a suboptimal one.
-The criteria for candidate selection remain unclear and may vary across queries.
-Figure 9: GPT-4.1 prediction
-Figure 10: Ground truth
-LLM Reasoning Summary Let’s analyze the problem step by step.
-First, I’ll look for large free areas without obstacles.
-Option 1: The region to the right of the chairs and table, from x=1.3 to 2.1, and from y=0 to 1.8.
-This gives an estimated area of 1.44 m².
-Option 2: The space above the chairs and below the wall units, from x=0 to 2.1 and y=1.1 to 1.8.
-This gives an area of 1.47 m².
-Option 3: A small gap near the bottom right corner.
-It looks too narrow and likely blocked by furniture.
-Option 4: The space in front of chair_2, but it appears too small to be useful.
-Based on these, Option 2 seems best.
-I’ll choose the rectangle [0.0, 1.1, 2.1, 1.8], which gives an area of 1.47 m².
-Figure 11: Summarization of chain-of-thought style output from GPT-4.1. The model imitates reasoning by listing and evaluating candidate regions, then selects a suboptimal solution.
-F.2 Shortest Path
-In this task, the model must find a path from the center of a starting object to the center of a destination object. For simplicity, the start and goal objects are not treated as obstacles.
-Our pathfinding system constructs a visibility graph by connecting all mutually visible corner points in the environment.
-Dijkstra’s algorithm is then used to compute the true shortest path over this graph, ensuring optimality under Euclidean distance.
-This method avoids the limitations of grid-based approximations and produces precise, flexible paths around obstacles.
-While multiple valid solutions may exist depending on object layout and clearance constraints, the chosen path always reflects the shortest feasible route.
-When interpreting the task, models often fail to make effective use of available free space.
-Instead of navigating around obstacles or adapting to the environment, they frequently extend a straight line in one direction until it is blocked.
-This leads to incomplete or inefficient trajectories that do not reflect the geometry of the room or required clearance.
-In some cases, paths cut too closely around objects, likely due to the model implicitly minimizing Euclidean distance without accounting for spatial constraints.
-Figure 13 shows a path that reaches the target but fails the Fréchet distant constraint of 0.6 m, making it non-optimal.
-Figure 13 depicts a path that violates clearance constraints and does not represent a valid trajectory to the target.
-Figure 12: Comparison between the optimal path (green) and the GPT-4.1-generated trajectory (cyan), which, despite reaching the target, exceeds the Fréchet distance threshold of 0.6 m.
-Figure 13: Depiction of the optimal path (green) alongside the GPT-4.1-generated trajectory (red), which fails to satisfy clearance requirements and results in an infeasible route.
-Appendix G VLM Experiments
-To further probe model performance on particularly challenging tasks, we conducted targeted evaluations of two vision-language models (VLMs).
-We focused on the most difficult question categories–Free Space and Max Box–within the most spatially complex environments: living rooms and bedrooms.
-The models tested in this focused benchmark include Gemini 2.5 Flash Preview 04-17 Gemini Team, Google (2025) and GPT-4.1.
-Results are summarized in Table 12, which compares each model’s output under two input conditions: JSON only and JSON combined with a floorplan image.
-An example of Gemini’s step-by-step reasoning under the JSON+Image condition is shown in Figure 14.
-Table 12: Comparison of spatial understanding between Gemini 2.5 Flash and GPT-4.1 for Living Room (L) and Bedroom (B), under different conditions.
-Type Category [Uncaptioned image] Gemini 2.5 Flash [Uncaptioned image] GPT-4.1
-JSON Only JSON + Image JSON Only JSON + Image
-L Free Space 6.4 12.2 5.0 9.8
-Max Box 17.0 22.0 6.5 4.3
-B Free Space 12.8 15.2 2.8 10.5
-Max Box 15.7 21.0 7.2 7.2
-[…]
-Figure 14: Summarized step-by-step reasoning by Gemini to compute maximum free space from a JSON floorplan and image.
-The model merges nested objects and analyzes rectangular gaps to estimate usable area.
-While our results show that both models benefit from visual input in some settings, the extent of improvement varies by task.
-For the Free Space category, both Gemini 2.5 Flash and GPT-4.1 show substantial gains when provided with floorplan images.
-In contrast, for Max Box, only Gemini shows a notable improvement, while GPT-4.1 performs similarly or worse with visual input.
-This variation suggests that visual grounding supports some spatial tasks more effectively than others.
-The stronger visual performance of Gemini 2.5 Flash may stem from architectural or training differences, possibly reflecting design choices aimed at robotics or spatial reasoning tasks.
-One challenge lies in the nature of floorplan inputs.
-Unlike natural images, floorplans vary widely in style, abstraction, and labeling.
-This inconsistency makes it harder for models to extract spatial meaning reliably.
-Smaller models show little benefit from visual input, likely due to limited capacity and lack of pretraining for structured spatial tasks.
-In summary, floorplan understanding requires more than object detection.
-It demands precise spatial reasoning and integration with structured formats like JSON.
-Without task-specific tuning or standardized visual formats, even advanced models remain limited in this domain.
+G.1 Pair Distance
+Task Definition
+In this task, the model is asked: “Calculate the euclidean distance between the centroids of the obj_1 and the obj_2.” In the visualization in Figure 8, these two polygons correspond to the sink and the shower; the goal is to compute their centroid-to-centroid distance.
+Figure 7: Pair Distance (bathroom): the red line indicates the centroid-to-centroid segment.
+Figure 8: Repositioning (bedroom): red fill indicates the initial pose of bin_2 and green fill indicates the final pose.
+Ground-Truth Computation
+To establish the correct answer, we first compute the centroid $(x,y)$ of each polygon.
+This is done using the shoelace formula, which calculates the centroid based on the polygon’s vertices.
+Once both centroids are obtained, the Euclidean distance between them is computed.
+A predicted answer is considered correct if it falls within a tolerance of $2\%$ of the ground-truth distance.
+Main Issue
+Models often fail on the HSSD dataset because they compute the centroid incorrectly.
+In earlier experiments, some models used the center of mass instead of the centroid; therefore, the word centroid is now explicitly stated in the prompt.
+Almost all wrong answers come from calculation mistakes in the centroid formula (areas, sums, divisions), not from the distance step.
+This error does not depend on polygon complexity (number of vertices).
+G.2 Repositioning
+Task Definition
+We pose the question: “Calculate how far the object can be moved in the direction before it touches another object or the wall.” In the visualization in Figure 8, the object of interest is bin_2 and the direction is leftward; the goal is to compute its maximum leftward translation before contacting a bathtub.
+Ground-Truth Computation
+Simulate a leftward, axis-aligned slide of bin_2.
+Advance until the next step would overlap a wall or another object; take the last non-overlapping pose.
+Measure the travel distance from the initial position to that pose.
+Accept model answers within $2\%$ tolerance.
+Main Issue
+Narrow gaps and obstacles with arbitrary orientations make clearance difficult to estimate.
+Models often fall back to axis-aligned bounding boxes (discarding shape orientation) or omit the obstacle-union step, which leads to systematic over- or underestimation of feasible travel distances in any direction (left, right, up, or down).
+G.3 Free Space
+Task Definition
+We consider the following question: “Calculate the total non-occupied floor area in the room?” In the visualization in Figure 10, the mint-highlighted area represents the space that remains free of objects within the game room; the goal is to compute its total area.
+Figure 9: Unoccupied Floor Area (game room): the mint fill indicates the unoccupied (free-space) region within the game room.
+Figure 10: View Angle (kitchen): smallest absolute angle between the vector from the centroid of the chair_2 to the centroid of the window_1 and global north $(0,1)$.
+Ground-Truth Computation
+We compute the free floor area using geometric operations provided by Shapely (Gillies et al., 2007).
+All object polygons within the room are merged using unary_union to correctly handle overlaps.
+The occupied area is then obtained from this union, and the free area is computed as
+$A_{\text{free}}=A_{\text{room}}-A_{\text{union(objects)}}.$
+A model prediction is considered correct if it falls within $5\%$ of the ground-truth free area.
+Main Issue
+Accurate free-space estimation hinges on correctly handling overlapping obstacles.
+A common failure mode is to subtract each object’s area independently, rather than forming their geometric union, which double-counts overlaps and systematically underestimates available area.
+For instance, when we evaluate on HSSD layouts using GPT-OSS-120B, cumulative accuracy declines as object count and overlap increase (see Figure 11), consistent with this union-omission error.
+Figure 11: Cumulative accuracy versus layout complexity for HSSD using GPT-OSS-120B on the Free-Space task.
+Accuracy declines sharply as object count and overlap increase, reflecting the model’s difficulty in handling overlapping geometries.
+G.4 View Angle
+Task Definition
+The prompt is: “Compute the smallest absolute angle between the vector from the centroid of obj_1 to the centroid of obj_2 and the global north vector $(0,1)$; report $\theta$ in degrees.” In the visualization in Figure 10 (kitchen), obj_1 is the chair_2 and obj_2 is the window_1; the goal is to return $\theta$, judged correct within a $2\%$ tolerance.
+Ground-Truth Computation
+(1) Compute centroids $\mathbf{c}_{s}$ (for sofa) and $\mathbf{c}_{v}$ (for TV) using the polygon shoelace formula.
+(2) Form the displacement vector $\mathbf{d}=\mathbf{c}_{v}-\mathbf{c}_{s}$ and its unit vector $\hat{\mathbf{d}}=\dfrac{\mathbf{d}}{\lVert\mathbf{d}\rVert}$.
+(3) Let the global north vector be $\mathbf{n}=(0,1)$ (already unit length). Compute the cosine via $\cos\theta=\mathrm{clip}\!\big(\hat{\mathbf{d}}\cdot\mathbf{n},\,-1,\,1\big)$.
+(4) Convert to degrees and take the smallest absolute angle: $\theta=\arccos(\cos\theta)\cdot\dfrac{180}{\pi}\ \in[0^{\circ},180^{\circ}]$.
+A prediction is correct if it is within $2\%$ of the ground-truth angle $\theta$.
+Main Issue
+On HSSD layouts, most errors come from centroid calculation mistakes (areas/sums/divisions in the shoelace step), not from the dot product.
+To avoid ambiguity, the prompt explicitly says centroid.
+On synthetic layouts (4-point, axis-aligned boxes), centroids are trivial, and this issue does not appear.
+G.5 Placement
+Task Definition
+We consider the following question: “Check if a given object can be placed in the room without overlapping walls or other objects?” In the visualization in Figure 13, the object is a $2.5\,\mathrm{m}\times 1.0\,\mathrm{m}$ antique storage chest and the room is the living room; the goal is to determine whether a collision-free placement is possible.
+This task evaluates collision detection, spatial constraints, and free-space reasoning.
+Figure 12: Placement (living room): determine whether an antique storage chest can be placed without overlapping walls or existing objects (collision-free feasibility).
+Figure 13: Shortest Path (living room): ground-truth shortest walkable path with $15\,\mathrm{cm}$ clearance is shown in purple; the model’s predicted path in red intersects armchair_1, illustrating a failure case for GPT-OSS-120B due to incorrect obstacle/clearance handling.
+Ground-Truth Computation
+Form the room polygon and the union of all existing object polygons.
+Allow arbitrary rotation (non–axis-aligned) for the $2\times 1.5$ m rectangle.
+Search over poses: for each orientation $\theta$, test placements where the rotated rectangle is strictly inside the room polygon and has no intersection with the object union (i.e., contains check for the room and disjoint check for obstacles).
+If any collision-free pose exists, return True; otherwise False.
+Compare the model’s Boolean prediction to this result.
+Main Issue
+The task is harder when non–axis-aligned placements are allowed.
+Models often mis-handle overlap checks under rotation and falsely report feasibility/infeasibility due to incorrect intersection computations.
+G.6 Shortest Path
+Task Definition
+The question asks: “Determine the shortest valid path that maintains a clearance of $d$ cm from all other objects, starting from centroid of the obj_1 and ending at the centroid of the obj_2.” In Figure 13 (living room), obj_1 is the TV, obj_2 is armchair_2, and $d=15\,\mathrm{cm}$; the goal is to compute the minimum-length collision-free path (and its length).
+The figure shows a failure case for GPT-OSS-120B, where the predicted path violates the clearance by intersecting the armchair_1.
+Ground-Truth Computation
+Offset obstacles (equivalently, erode free space) by $0.15\,\mathrm{m}$ to enforce clearance.
+Run A* on the navigable grid to obtain the shortest collision-free path polyline between TV and armchair_2.
+A model path is valid if it is collision-free under the same clearance; it is judged correct if its Fréchet distance to the ground-truth path is $\leq 0.6\,\mathrm{m}$.
+Main Issue
+More objects and overlaps make clearance buffering, merge obstacles, and narrow corridors, increasing failure modes.
+Models often mishandle overlaps, producing paths that cut through obstacles or declaring no path when one exists.
+G.7 Visibility
+Task Definition
+The prompt is: “Find all objects that intersect the vector from the centroid of the obj_1 to the centroid of the obj_2.” In the visualization in Figure 15 (office), obj_1 is the window and obj_2 is the bin; the goal is to return the set of objects that intersect this segment (excluding the endpoints).
+Figure 14: Visibility (office): the table and armchair (highlighted in orange) intersect the line segment from the centroid of the window to the centroid of the bin and constitute the correct answer.
+Figure 15: Max Box (bedroom): the red rectangle shows the largest rectangular region that can be placed without overlaps, excluding soft coverings such as rugs.
+Ground-Truth Computation
+Compute the centroids of window and bin; form the line segment between these centroids.
+Return the set of objects whose bounding boxes intersect this segment, excluding endpoint touches (i.e., ignore cases where the segment only touches at its endpoints).
+Main Issue
+On HSSD layouts, accuracy is slightly worse due to the larger number of objects and overlaps; the increased number of polygons along the line increases intersection ambiguity and error rates.
+G.8 Max Box
+Task Definition
+We consider the following question: “Calculate the area in square meters of the largest rectangle that can fit inside the room”.
+In the visualization in Figure 15 (bedroom), the goal is to compute the maximum-area non–axis-aligned rectangle that fits without overlapping any obstacles, and to report its area.
+Ground-Truth Computation
+Let $R$ be the room polygon and $O$ the union of all object polygons except rugs (soft coverings).
+Compute free space $F=R\setminus O$.
+Search over orientations $\theta\in[0,\pi)$: rotate $F$ by $\theta$, find the largest axis-aligned empty rectangle inside the rotated $F$, record $(w_{\theta},h_{\theta})$ and area $A_{\theta}=w_{\theta}h_{\theta}$, then map back to get $(w^{*},h^{*},\theta^{*})$ with $A^{*}=\max_{\theta}A_{\theta}$.
+Main Issue
+Harder than simple placement: the model must optimize size and orientation, not just answer yes/no.
+Allowing rotation makes the search non-convex; more objects and overlaps increase combinatorial complexity.
+Models often (i) ignore rotation and return an axis-aligned box, or (ii) mis-handle overlaps in free space, leading to under- or over-estimated maxima.
 Appendix H Prompts
-This section presents selected prompt templates used in our benchmark.
-We include a summarized version of the bedroom layout generation prompt, as well as representative examples of question prompts targeting specific spatial reasoning skills.
-These prompts reflect the structured task design outlined in Table 11.
-Additional prompts–including full versions, formatting logic, and implementation details–are provided in the supplementary code and materials for reproducibility.
-Prompt: Max Box You’ll be given a layout of a {room_type} in JSON format.
-Your task is to calculate the area (in m²) of the largest axis-aligned rectangle that can fit within the room.
-The rectangle must not overlap with any other objects, but it may fully occupy free space, including areas above rugs.
-You can ignore passageways, door clearance, or accessibility constraints.
-Room layout: {room_json} Reason step by step to calculate the area.
-Respond only with a float value in square meters using the format: *Final answer*: area
+This section illustrates the design of prompt templates used in our benchmark.
+We first show a representative example of a question prompt, demonstrating how natural language templates are instantiated to elicit spatial reasoning skills (Figure 16).
+Next, we present two examples of layout-generation prompts for bedrooms.
+The first specifies the creation of base room boundaries and openings (walls and windows) (Figure 17).
+The second demonstrates how furniture and objects are placed within the generated layout to yield a complete scene (Figure 18).
+For completeness, full prompt templates, formatting rules, and implementation details are provided in the supplementary code to support reproducibility.
+Prompt: Free Space Given the {room_type} layout in {format}, calculate the total non-occupied (free) floor area in square meters (m2).
+Room layout: {room} Begin with printing a concise checklist (3–7 bullets) of the conceptual steps necessary for calculating the free space.
+Then, carefully walk through each reasoning step required to calculate the area.
+If the format, object names, or required input data are missing, invalid, or inconsistent, reply with: *Final answer*: ERROR Limit your output to the step-by-step reasoning only, and do not include any internal reasoning unless explicitly requested.
+Clearly state the final answer on the last line using the exact format specified below.
+### Output Format
+<step-by-step calculations>
+*Final answer*: <area>
+Where <area> is a float rounded to three decimal places, representing the free area in m2. For example: *Final answer*: 12.347
 Figure 16: Prompt for computing the largest empty rectangle area within a room layout using Chain-of-Thought reasoning.
-Prompt: Missing Object Placement You’ll be given a layout of a {room_type} in JSON format.
-Your task is to suggest a suitable placement for the missing object {object_name} by providing its bounding box coordinates.
-The placement must avoid overlaps, respect obstacles, and maintain accessibility.
-Room layout: {room_json} Reason step by step to identify the placement.
-Respond only with bounding box coordinates (float values) using the format: *Final answer*: [x_min, y_min, x_max, y_max]
-Figure 16: Prompt for placing a missing object in a room layout using Chain-of-Thought reasoning.
-The model must return valid coordinates that satisfy geometric constraints and maintain functional space.
-[…]
+Prompt: Generate Bedroom Layouts Generate a dataset of {N} bedroom layouts in JSON format.
+Each layout must include: A unique layout_id A room dictionary with: width, depth, units (meters) shape (”rectangular”, ”L-shaped”, or ”open”) shape_description, intended_use, and bed_size_suggestion An objects list with dictionaries containing: label, bbox [y0, x0, y1, x1], and a descriptive comment Layouts must obey structural and spatial constraints: 50% rectangular, 30% L-shaped, 20% open.
+L-shape cutouts in corners; each remaining segment $\geq$ 1.5 m.
+All layouts must include a door (except open types); avoid placing doors and windows on the same short wall.
+Windows must span >>15% of usable floor area, with equal sizing on shared walls and valid grouping logic.
+Optional elements: fireplace (for master bedrooms), closet alcove.
+No overlap or out-of-bound placement.
+Fireplace must not overlap with doors/windows.
+Follow a consistent coordinate system: top-left origin, x=width (left to right), y=depth (top to bottom).
+Return a JSON list of {N} valid layouts.
+No comments or trailing metadata.
 Figure 17: Summarized data generation prompt for producing structured and constrained bedroom layouts.
@@
 Figure 18: Prompt for populating a bedroom layout with functionally and spatially valid object placements, following layout-specific design rules.
+Appendix I Supplementary Accuracy Analyses
+Table 10: % token-limit stop reason for general models.
+Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
+Pair distance K 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.2
+LR 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.2
+B 0.0 0.0 0.2 0.0 0.0 0.2 0.0 0.8
+HSSD 0.0 0.5 0.0 0.0 17.5 8.5 3.5 1.5
+Placement K 0.0 0.0 0.0 0.0 6.5 0.0 1.2 1.2
+LR 0.0 0.0 0.0 0.0 6.0 0.0 1.0 1.5
+B 0.0 0.0 0.0 0.0 17.2 0.0 2.2 0.5
+HSSD 0.0 0.0 0.5 0.0 6.5 0.0 1.5 4.5
+Reposi- tioning K 0.0 0.2 0.0 0.0 5.7 0.0 2.3 0.2
+LR 0.0 0.2 0.0 0.0 1.7 0.0 1.0 1.0
+B 0.0 0.0 0.0 0.0 7.5 0.2 0.8 1.3
+HSSD 0.0 0.0 0.0 0.0 47.0 2.0 29.0 4.5
+Free space K 0.0 0.0 0.2 0.0 0.2 0.0 0.3 2.5
+LR 0.0 0.3 0.0 0.0 4.0 0.0 1.0 10.2
+B 0.0 0.0 0.0 0.0 4.5 0.0 0.3 6.3
+HSSD 0.0 2.0 0.0 0.0 52.5 1.0 39.0 28.5
+Visibility K 0.2 0.0 0.0 0.0 0.0 0.0 0.2 0.3
+LR 0.2 0.0 0.0 0.0 0.2 0.2 0.5 0.3
+B 1.0 0.0 0.0 0.0 0.2 0.3 1.3 0.7
+HSSD 0.0 0.0 0.0 0.0 11.5 2.5 11.0 0.0
+View angle K 0.0 0.0 0.0 0.0 0.5 0.0 0.0 0.3
+LR 0.0 0.0 0.2 0.2 1.5 0.0 0.2 0.7
+B 0.0 0.0 0.0 0.3 0.5 0.0 0.5 1.2
+HSSD 0.0 0.5 0.0 0.0 26.5 3.0 17.5 2.5
+Max box K 0.0 0.0 0.0 0.0 22.5 0.0 6.2 1.5
+LR 0.0 0.0 0.7 0.0 49.8 0.0 29.3 0.8
+B 0.0 0.0 1.5 0.0 46.5 0.3 21.8 0.8
+HSSD 0.0 0.0 0.5 0.0 45.5 0.5 20.0 3.0
+Shortest path K 0.0 0.2 0.5 0.0 44.3 0.0 41.5 6.7
+LR 0.0 0.0 0.0 0.0 37.8 0.2 46.3 7.8
+B 0.2 0.0 0.0 0.0 40.2 0.0 49.0 9.2
+HSSD 0.0 0.0 0.0 0.0 35.5 0.0 43.0 27.0
+Table 11: Question-level accuracy on completed answers for general models.
+Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
+Pair distance K 99.8 96.5 95.7 96.8 99.2 90.7 89.5 58.9
+LR 99.5 95.2 94.2 96.8 99.7 88.5 88.8 62.4
+B 99.7 96.3 93.5 96.5 99.5 88.0 85.2 60.5
+HSSD 88.0 56.3 75.5 66.0 81.2 40.4 46.1 56.9
+Placement K 87.8 78.0 82.2 80.3 96.4 86.5 86.5 75.0
+LR 80.5 69.0 73.8 83.2 94.9 75.2 86.5 72.9
+B 68.8 59.8 67.5 70.8 92.6 68.7 78.7 56.6
+HSSD 72.0 64.5 73.4 70.0 87.7 76.5 72.6 57.1
+Reposi- tioning K 73.8 63.9 48.7 45.3 88.5 66.3 75.4 14.9
+LR 79.3 60.4 56.7 64.5 92.9 76.8 72.9 25.3
+B 71.0 55.5 48.3 59.5 85.4 72.6 70.6 21.5
+HSSD 42.0 47.0 28.0 34.0 73.6 41.3 46.5 10.5
+Free space K 97.8 93.2 83.1 84.2 95.2 95.2 84.1 67.9
+LR 0.2 14.2 2.8 1.8 3.7 1.2 0.5 3.0
+B 2.7 31.2 1.0 1.3 9.3 0.8 1.0 0.7
+HSSD 35.0 16.3 24.0 22.0 35.8 15.7 12.3 9.1
+Visibility K 63.3 87.7 54.5 67.0 98.3 90.2 91.6 18.6
+LR 52.8 81.7 43.3 52.2 98.5 87.0 89.1 10.0
+B 58.1 74.8 41.0 54.0 96.8 86.6 87.5 10.9
+HSSD 20.0 46.5 22.5 26.5 79.7 53.3 51.1 9.0
+View angle K 92.0 95.3 69.8 78.8 97.5 95.8 74.7 49.8
+LR 87.7 93.2 59.8 75.3 95.4 93.0 72.3 40.8
+B 88.0 90.2 60.3 72.8 95.5 91.2 77.2 36.3
+HSSD 67.5 55.3 28.5 46.5 68.7 47.4 41.8 30.8
+Max box K 47.2 31.8 32.2 26.9 84.5 27.5 28.2 4.6
+LR 7.8 7.0 5.1 5.7 44.5 4.5 12.5 0.5
+B 5.8 6.8 7.4 5.0 54.5 4.9 9.0 1.7
+HSSD 5.0 7.5 2.5 2.0 21.1 4.5 3.8 1.0
+Shortest path (valid) K 59.2 61.7 52.7 39.7 81.1 55.3 37.3 36.6
+LR 53.0 51.3 42.8 37.3 72.1 47.3 37.9 32.9
+B 48.6 52.2 40.7 34.7 73.8 45.8 36.3 36.9
+HSSD 28.5 25.0 18.5 18.0 40.3 15.0 9.7 14.4
+Shortest path (Fréchet) K 45.3 56.8 51.3 28.2 79.3 39.5 30.5 38.8
+LR 24.7 42.5 27.3 12.3 55.8 26.5 17.1 16.8
+B 23.0 42.2 27.7 14.2 63.5 30.5 16.0 20.2
+HSSD 15.5 22.5 18.0 8.0 31.0 12.5 4.4 15.8
+Table 12: % token-limit stop reason for reasoning models.
+Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
+Pair distance K 0.0 0.0 1.5 3.3 0.0 0.7 2.0
+LR 0.0 0.0 0.8 4.0 0.2 0.7 2.3
+B 0.0 0.0 2.5 4.3 0.2 0.7 4.7
+HSSD 0.0 11.0 70.0 86.5 67.0 39.5 72.0
+Placement K 13.2 0.0 5.2 39.2 6.5 4.5 29.5
+LR 22.7 0.2 7.2 45.8 10.3 12.3 52.5
+B 36.2 0.2 8.2 63.3 12.8 25.7 64.5
+HSSD 27.0 0.0 6.0 84.0 16.5 15.0 70.5
+Reposi- tioning K 0.3 0.0 8.3 1.0 0.0 0.8 7.2
+LR 0.0 0.0 7.5 2.2 0.0 0.5 6.7
+B 0.2 0.0 3.2 1.7 0.0 0.8 8.3
+HSSD 2.0 0.5 33.5 76.0 28.5 22.5 63.0
+Free space K 1.2 0.0 5.5 2.0 0.0 0.8 2.3
+LR 0.2 0.2 41.7 41.7 1.7 13.7 2.3
+B 0.2 0.0 12.2 30.7 1.0 7.8 3.0
+HSSD 5.0 28.5 79.5 96.5 85.0 77.5 98.5
+Visibility K 0.0 1.5 26.7 72.8 0.0 0.3 19.2
+LR 0.0 1.7 46.8 88.2 0.3 1.0 26.3
+B 0.0 1.2 44.5 88.3 1.5 0.2 33.3
+HSSD 0.0 5.0 87.0 99.5 56.5 29.0 96.0
+View angle K 0.0 0.0 25.5 5.5 11.2 1.0 1.0
+LR 0.0 0.0 30.8 5.0 14.3 1.3 0.7
+B 0.0 0.0 23.8 3.7 12.7 1.2 0.8
+HSSD 0.0 7.5 84.5 75.5 73.0 38.0 66.0
+Max box K 0.0 0.0 30.2 95.7 2.0 28.8 62.7
+LR 0.0 0.0 63.5 100.0 5.7 61.0 98.0
+B 0.0 0.0 59.3 100.0 4.7 56.0 98.2
+HSSD 0.0 0.0 58.5 100.0 22.5 33.0 99.0
+Shortest path K 0.2 0.5 78.5 96.2 23.5 36.2 85.2
+LR 61.7 1.2 77.8 97.8 19.8 40.3 81.3
+B 0.5 0.5 79.5 98.0 17.5 50.3 83.7
+HSSD 0.0 3.0 82.0 100.0 65.5 32.5 98.0
+Table 13: Question-level accuracy on completed answers for reasoning models.
+Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
+Pair distance K 99.8 99.3 98.0 96.3 100.0 94.2 97.7
+LR 98.8 99.3 99.0 96.0 99.7 93.5 97.5
+B 98.3 99.5 96.8 95.5 99.7 93.8 95.3
+HSSD 69.0 78.5 25.5 12.5 32.5 40.5 18.0
+Placement K 84.7 92.0 89.0 59.7 90.8 85.7 68.2
+LR 75.5 89.0 82.2 53.3 86.3 78.5 46.5
+B 61.2 83.5 72.0 35.8 81.2 62.0 34.5
+HSSD 70.0 85.0 79.0 16.0 75.5 74.5 28.5
+Reposi- tioning K 83.0 85.5 79.2 90.5 84.5 70.8 61.5
+LR 85.5 89.8 86.2 91.2 92.8 87.8 77.0
+B 77.8 83.3 83.3 85.2 84.3 78.0 69.2
+HSSD 49.5 60.5 47.5 18.5 53.5 40.0 27.0
+Free space K 82.5 99.0 93.0 93.3 99.5 94.8 97.0
+LR 47.0 83.3 18.3 17.7 78.5 53.2 0.0
+B 50.5 87.5 34.8 33.3 82.2 74.0 1.0
+HSSD 19.5 31.0 6.5 1.0 5.0 9.0 1.0
+Visibility K 94.8 94.2 71.3 26.8 98.0 91.5 78.8
+LR 95.2 94.0 52.0 11.3 98.0 89.3 70.8
+B 94.2 92.5 53.5 11.2 95.5 89.2 64.2
+HSSD 57.0 70.0 10.0 0.5 39.0 45.5 3.0
+View Angle K 96.2 98.5 73.7 92.5 88.3 93.5 98.5
+LR 93.3 97.3 68.2 91.5 84.5 92.0 98.3
+B 95.2 98.2 75.2 93.8 86.8 91.3 98.3
+HSSD 59.5 74.0 13.5 20.0 25.5 37.5 26.0
+Max Box K 48.5 62.8 50.5 3.7 85.2 31.3 16.7
+LR 17.3 28.2 8.0 0.0 60.3 3.8 0.8
+B 13.0 30.3 11.0 0.0 61.2 5.8 0.5
+HSSD 5.0 9.5 2.5 0.0 17.0 0.5 0.0
+Shortest path (valid) K 64.7 64.2 12.3 3.2 52.3 43.0 14.2
+LR 21.2 66.0 12.5 1.5 53.7 37.7 16.7
+B 58.2 57.8 7.8 1.0 52.0 30.0 14.3
+HSSD 28.5 33.5 8.5 0.0 16.5 13.5 1.0
+Shortest path (Fréchet) K 56.3 55.5 11.5 3.2 50.5 42.2 14.0
+LR 12.3 40.5 9.3 1.3 47.5 28.5 16.3
+B 39.3 44.8 6.0 0.8 47.7 24.2 14.3
+HSSD 22.5 26.5 2.5 0.0 12.0 14.0 1.0

### v3

@@
 We introduce FloorplanQA, a diagnostic benchmark for evaluating spatial reasoning in large-language models (LLMs).
-FloorplanQA is grounded in structured representations of indoor scenes, such as (e.g., kitchens, living rooms, bedrooms, bathrooms, and others), encoded symbolically in JSON or XML layouts.
+FloorplanQA is grounded in structured representations of indoor scenes (e.g., kitchens, living rooms, bedrooms, bathrooms, and others), encoded symbolically in JSON or XML layouts.
 The benchmark covers core spatial tasks, including distance measurement, visibility, path finding, and object placement within constrained spaces.
-Our results across a variety of frontier open-source and commercial LLMs reveal that while models may succeed in shallow queries, they often fail to respect physical constraints, preserve spatial coherence, though they remain mostly robust to small spatial perturbations.
+Our results across a variety of frontier open-source and commercial LLMs reveal that while models may succeed on shallow queries, they often fail to respect physical constraints and preserve spatial coherence, though they remain mostly robust to small spatial perturbations.
 FloorplanQA uncovers a blind spot in today’s LLMs: inconsistent reasoning about indoor layouts.
 We hope this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
+Spatial Reasoning, Layout Reasoning, Scene Understanding, Structured Scene Representations, Benchmark, Large Language Models (LLMs)
 1 Introduction
 Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over plausible, physically feasible environments such as indoor layouts remains poorly understood.
-In numerous practical applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images or natural language.
+In numerous practical applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images.
 Reasoning in these contexts requires geometric inference over symbolic layouts, not pixel-level perception.
 We introduce FloorplanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floorplans represented in structured text-based formats.
-Each instance consists of a JSON-encoded layout paired with natural language questions that require the model to compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints.
+Each instance consists of a symbolic (JSON-encoded) layout paired with natural language questions that require the model to reason about spatial and geometric aspects of the scene, compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints.
 FloorplanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments.
-Although LLMs can increasingly be used in tool-assisted pipelines, for example to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities.
+In contrast to many layout benchmarks that focus on qualitative relations or output realism, FloorplanQA emphasizes geometric consistency under explicit metric, clearance, and collision constraints.
+Although LLMs can increasingly be used in tool-assisted pipelines, for example, to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities.
 FloorplanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding, in order to measure their unassisted capabilities.
@@
 Specifically, our contributions are as follows:
-  • We introduce a dataset of 2,000 structured 2D layouts, including 600 each from synthetically generated kitchens, living rooms, and bedrooms, plus 200 layouts sourced from the Habitat Synthetic Scenes Dataset (HSSD) (Khanna et al., 2023), providing a realism check.
+  • We introduce a dataset of 2,000 structured 2D layouts, including 200 layouts sourced from the Habitat Synthetic Scenes Dataset (HSSD) (Khanna et al., 2023) based on real floorplans, and 600 each from synthetically generated kitchens, living rooms, and bedrooms.
 All are represented in JSON and paired with spatial reasoning questions.
@@
 Prior benchmarks have explored spatial reasoning across vision and language domains.
-CLEVR Johnson et al.
-(2017) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relations.
-In real-world settings, SpatialSense Yang et al.
-(2019) focuses on recognizing spatial relations in images through adversarially mined examples.
-Benchmarks like BabyAI Chevalier-Boisvert et al.
-(2019), ALFRED Shridhar et al.
-(2020), and Room-to-Room (R2R) Anderson et al.
-(2018) integrate spatial understanding into embodied tasks, requiring agents to follow instructions involving navigation and object manipulation in simulated environments.
-Recent datasets such as ScanQA Azuma et al.
-(2022) and 3DSRBench Ma et al.
-(2024a) extend spatial reasoning evaluation into 3D environments, emphasizing the need for models to comprehend and reason about spatial relationships in three dimensions.
+CLEVR (Johnson et al., 2016) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relations.
+In real-world settings, SpatialSense (Yang et al., 2019) focuses on recognizing spatial relations in images through adversarially mined examples.
+Benchmarks like BabyAI (Chevalier-Boisvert et al., 2018), ALFRED (Shridhar et al., 2019), and Room-to-Room (R2R) (Anderson et al., 2017) integrate spatial understanding into embodied tasks, requiring agents to follow instructions involving navigation and object manipulation in simulated environments.
+Recent datasets such as ScanQA (Azuma et al., 2021) and 3DSRBench (Ma et al., 2024a) extend spatial reasoning evaluation into 3D environments, emphasizing the need for models to comprehend and reason about spatial relationships in three dimensions.
 Vision-language models have advanced spatial reasoning but often handle it qualitatively.
-The VQA dataset Antol et al.
-(2015) challenges models to answer questions about images, while VL-T5 Cho et al.
-(2021) unifies vision-and-language tasks via text generation.
-Recent work on 3D scene graphs Armeni et al.
-(2019) introduces structured representations of environments, facilitating spatial reasoning.
+The VQA dataset (Agrawal et al., 2015) challenges models to answer questions about images, while VL-T5 (Cho et al., 2021) unifies vision-and-language tasks via text generation.
+Recent work on 3D scene graphs (Armeni et al., 2019) introduces structured representations of environments, facilitating spatial reasoning.
 However, these approaches may miss fine-grained geometric details necessary for precise spatial inference.
-Efforts like SpatialVLM Chen et al.
-(2024) aim to endow vision-language models with enhanced spatial reasoning capabilities, addressing some of these limitations.
+Efforts like SpatialVLM (Chen et al., 2024) aim to endow vision-language models with enhanced spatial reasoning capabilities, addressing some of these limitations.
 Advancements in generative models have also contributed to spatial reasoning tasks.
-LayoutGPT Feng et al.
-(2023) leverages large language models for compositional visual planning and layout generation, while Holodeck Yang et al.
-(2024) enables language-guided generation of 3D embodied AI environments.
-Similarly, AnyHome Fu et al.
-(2024) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts.
+LayoutGPT (Feng et al., 2023) leverages large language models for compositional visual planning and layout generation, while Holodeck (Yang et al., 2023) enables language-guided generation of 3D embodied AI environments.
+Similarly, AnyHome (Wen et al., 2023) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts.
 Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but often produces implausible object placement due to non-convergent simulated annealing.
-LayoutVLM Sun et al.
-(2024) and FirePlace Huang et al.
-(2025) improve layout generation via optimization and constraint solving, respectively.
+LayoutVLM (Sun et al., 2024) and FirePlace (Huang et al., 2025) improve layout generation via optimization and constraint solving, respectively.
 But they assess output realism, not the model’s ability to infer constraints directly.
 In contrast, our benchmark tests symbolic reasoning without tool-assisted refinement.
-Evaluations of large language models’ spatial understanding have been conducted in studies like Evaluating Spatial Understanding of Large Language Models Yamada et al.
-(2024), which assesses the spatial reasoning capabilities of LLMs through structured tasks.
-Additionally, benchmarks such as BALROG Paglieri et al.
-(2025) test agentic reasoning in game environments, further exploring the spatial and decision-making abilities of language and vision-language models.
-While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning in structured layouts, providing fine-grained and task-specific insights into models’ spatial competence.
-Recent 3D-LLM surveys such as (Ma et al., 2024b) cover tasks like navigation and interaction, but not symbolic spatial reasoning.
-FloorplanQA fills this gap by testing raw spatial competence from structured layouts without multimodal input.
-FloorplanQA addresses the gap in existing benchmarks by directly evaluating structured spatial inference from symbolic room layouts.
-Unlike prior benchmarks relying on raw images or focusing on commonsense spatial language, FloorplanQA provides explicit spatial representations (object coordinates and dimensions) and tests models’ abilities to perform precise spatial reasoning tasks, such as calculating distances, assessing visibility, and verifying object fit within a controlled setting.
+Evaluations of large language models’ spatial understanding have been conducted in studies like Evaluating Spatial Understanding of Large Language Models (Yamada et al., 2023), which assesses spatial reasoning through natural language descriptions of qualitative relations (e.g., ”left of,” ”above”), without numeric coordinates or metric computation.
+Additionally, benchmarks such as BALROG (Paglieri et al., 2025) test agentic reasoning in game environments (NetHack, BabyAI, etc.), where spatial reasoning emerges implicitly through sequential navigation commands rather than explicit geometric inference.
+While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning over structured coordinate-based layouts, requiring models to compute distances, angles, collision-free paths, and spatial unions directly from symbolic representations.
+Recent 3D-LLM surveys, such as Ma et al.
+(2024b), cover tasks like embodied navigation and interaction, but not symbolic geometric reasoning from structured layouts.
+FloorplanQA addresses this gap by providing explicit spatial representations (object coordinates and dimensions).
+Unlike prior benchmarks relying on raw images or commonsense spatial language, it tests precise geometric tasks: computing distances, angles, collision-free paths, and spatial unions directly from symbolic input.
 3 Method
-3.1 Synthetic Layout Generation
-Our initial aim was to use publicly available real-world floorplan datasets.
-However, a comprehensive review revealed significant limitations, as several prominent datasets such as SUNCG (Song et al., 2017) and HouseExpo (Li et al., 2019) are not accessible due to unresolved copyright claims.
-Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2020), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse.
-Datasets like CubiCasa5K (Kalervo et al., 2019) and Rent3D (Liu et al., 2015) offer fixed architectural plans from real environments but lack furnishing annotations.
-RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al.
-(2020), while large, is procedurally generated with realtor supervision and imposes restrictions on reuse.
-Given these legal, practical, and methodological constraints, we found synthetic data generation to be the most viable alternative.
-We generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro, a large language model fine-tuned for spatial reasoning (Google Gemini Robotics Team, 2025).
-Although our evaluation includes multiple LLMs (including Gemini variants), there is no circularity: the data-generation step is one-off and separate from evaluation.
-For evaluation, models solely produce answers to the benchmark questions, and correctness is computed by deterministic geometric routines against ground-truth solutions derived from the layouts, so no model output is used to generate data or to grade itself.
-The generation process comprises two stages.
-First, we specify room geometries using explicit constraints on shape, adjacency, and design principles related to circulation, symmetry, and zoning.
-These constraints are encoded directly in the LLM prompt.
-Second, each room is furnished according to style-specific guidelines (e.g., a bedroom must contain a bed and storage), also defined in structured prompts, to encourage both visual realism and functional plausibility.
-Approximately one-third of candidate layouts are filtered out by a rule-based spatial validity filter that enforces basic clearance and accessibility constraints.
-The checks remove scenes with inaccessible furniture and implausible adjacencies, such as sofas blocking doors or a refrigerator overlapping a table; see Appendix B.3 for the full set of cases.
-Full prompts, generation templates, and validation scripts are provided in the Supplementary Material.
-3.2 Layout Extraction from HSSD Dataset
-To complement the synthetically generated layouts, we further incorporated 200 layouts extracted from the Habitat Synthetic Scenes Dataset (HSSD-200).
-HSSD provides 211 high-quality, human-authored 3D scenes designed with the Floorplanner interface and populated with 18,656 objects across 466 semantic categories.
-Unlike purely procedural datasets, HSSD offers fine-grained semantics, 3D assets, and close correspondence to real interiors, making it an effective proxy for real-world interior layouts.
-For our purposes, we project each 3D scene to a 2D floorplan and retain only select structural and furniture elements.
-Decorative or auxiliary objects (e.g., vases, plants, cushions, artworks, posters, bottles, shoes, candles) are removed to reduce clutter; see Appendix C for the full details.
-We then use an $\alpha$-convex hull (Asaeedi et al., 2014; Edelsbrunner & Mücke, 1994) to smooth object boundaries, yielding polygonal layouts that are not restricted to axis-aligned rectangles.
-This step is necessary because raw HSSD projections often produce overly dense polygons, with redundant vertices along straight or nearly straight segments; applying an $\alpha$-hull reduces spurious complexity while preserving concavity, which avoids unnecessary token overhead in downstream LLM processing.
-This ensures compatibility with our synthetic layouts, while maintaining the richer geometric variety of HSSD.
-3.3 Unified Dataset
-Together, these two sources yield a dataset of 2,000 layouts: 1,800 synthetically generated via Gemini 2.5 Pro and 200 extracted from HSSD.
-The two subsets share a unified polygonal representation, enabling consistent downstream processing.
-Figure 1 shows examples from both sources while Table 1 summarizes room, style, and geometry distributions across the synthetic subset with the figure next to it providing a breakdown by room type.
-Table 1: Layout distribution in FloorplanQA by room type, style, and geometry, with room-wise totals.
 Room Type Style Rectangular L-Shaped Open Total
@@
 Total 893 484 423 1800
+Figure 1: (Left) The illustration of the overall breakdown by room type for the entire 2,000-layout benchmark, encompassing both the synthetic component and the layouts extracted from the HSSD.
+(Right) Detailed distribution of the synthetic subset of the FloorplanQA benchmark (1,800 layouts) across room type, internal style, and geometric configuration.
+3.1 Layouts Used in This Study
+We aim to create a benchmark based on floorplan datasets, with some critical requirements; first, they need to be publicly accessible and second, they need to include furnishings (couches, tables, chairs, etc) at a minimally plausible level.
+These requirements matter because FloorplanQA’s questions require object-level geometry (sizes, orientations, clearances), and we need a dataset that others can actually obtain, so that results are reproducible.
+In practice, we were unable to find a dataset that meets these requirements, as several prominent datasets such as SUNCG (Song et al., 2016) and HouseExpo (Li et al., 2019) are not accessible due to unresolved copyright claims.
+Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2019), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse.
+Datasets like CubiCasa5K (Kalervo et al., 2019) and Rent3D (Liu et al., 2015) offer fixed architectural plans from real environments but lack furnishing annotations.
+RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al.
+(2020), while large, is procedurally generated with realtor supervision and imposes restrictions on reuse.
+Given these legal, practical, and methodological constraints, we found a combination of data derived from HSSD Khanna et al.
+(2023) based on real data, extended with LLM-based synthetic data generation, to be the most viable alternative.
+We incorporated 200 layouts extracted from HSSD (Khanna et al., 2023).
+HSSD provides 211 high-quality, human-authored 3D scenes populated with 18,656 objects across 466 semantic categories.
+Unlike purely procedural datasets, HSSD offers fine-grained semantics, 3D assets, and close correspondence to real interiors, making it an effective proxy for real-world interior layouts.
+Decorative or auxiliary objects (e.g., vases, plants, cushions, artworks, posters, bottles, shoes, candles) are removed to reduce clutter; see Appendix C for the full details.
+The original data is 3D, so we project each 3D scene to a 2D floor plan.
+This can result in dense polygons for each asset (repeated points, points on straight lines, excessive tokens), which would confound our ability to test 2D geometric reasoning, so we use the Douglas-Peucker algorithm (Douglas and Peucker, 1973) with $\epsilon=0.01$m to simplify the shapes (see Figure 4 for an example, Appendix C).
+To increase the diversity of our benchmark, we generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro (at the time the most capable publicly available variant), with all layouts generated up-front and fixed prior to evaluation.
+The Gemini 2.5 Pro model has been trained on spatial reasoning and robotic manipulation tasks (Google Gemini Robotics Team, 2025), making it well-suited for generating geometrically plausible indoor layouts.
+Although our evaluation includes multiple LLMs (including Gemini variants), there is no circularity: (1) layout generation is a one-time process separate from benchmark evaluation, (2) all ground-truth answers are computed deterministically using geometric algorithms (not model outputs), and (3) correctness is verified via rule-based spatial calculations.
+The generation process comprises two stages.
+First, we specify room geometries using explicit constraints on shape, adjacency, and design principles related to clearance, symmetry, and zoning (see Appendix B for more details).
+These constraints are encoded directly in the LLM prompt.
+Second, each room is furnished according to style-specific guidelines (e.g., a bedroom must contain a bed and storage), also defined in structured prompts, to encourage both visual realism and functional plausibility.
+Approximately one-third of candidate layouts are filtered out by a rule-based spatial validity filter that enforces basic clearance and accessibility constraints.
+The checks remove scenes with inaccessible furniture and implausible adjacencies, such as sofas blocking doors or a refrigerator overlapping a table.
+Full prompts, generation templates, and validation scripts are provided in the Supplementary Material.
+3.2 Unified Dataset
 Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 1: Representative layouts from FloorplanQA. Generated: kitchen, living room, bedroom. HSSD: last image. Generated objects are axis-aligned boxes; HSSD uses arbitrary polygons.
+Figure 2: Representative layouts from FloorplanQA. HSSD: living room (first image) Generated: bedroom, kitchen and living room. Generated objects are axis-aligned boxes; HSSD uses arbitrary polygons.
+Together, these two sources yield a publicly available dataset of 2,000 layouts: 1,800 synthetically generated via Gemini 2.5 Pro and 200 extracted from HSSD.
+The two subsets share a unified polygonal representation, enabling consistent downstream processing.
+Figure 2 shows examples from both sources, while Figure 1 (left) plots the room-type distribution; Figure 1 (right) reports the corresponding room/style/geometry counts for the synthetic subset.
+Layouts are represented in a structured JSON format.
+Each entry contains a layout_id, the room_type, and explicit geometric descriptions.
+The room_boundary is stored as a closed polygon, while walls are represented separately as a list of line segments.
+Openings such as windows and doors are included in a dedicated openings field rather than flattened into the object list.
+All furnishings and functional elements (e.g., bed, sofa, table) are stored in the objects list, with each object defined by a labeled polygon.
+In the synthetic data, these polygons are axis-aligned bounding boxes (four points), whereas in HSSD they can exhibit arbitrary shapes and orientations.
+Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents remain unique and stable across prompt construction and answer evaluation.
+Coordinates are expressed in meters in a right-handed 2D Cartesian frame (+x+x to the right, +y+y upward), with origin at $(0,0)$.
+During synthetic data generation, layouts are initially produced in a top-left origin frame (+y downward) following Gemini 2.5 Pro’s native output format.
+All generated layouts are automatically transformed to the right-handed coordinate system before benchmark construction, ensuring consistency across synthetic and HSSD data.
+HSSD layouts are similarly transformed from their original 3D projection coordinates to this unified 2D frame.
+The prompt used to generate examples according to this schema is in Appendix J.
 3.3 Question Taxonomy and Prompting
-FloorplanQA assesses spatial reasoning by presenting models with a single natural language question per symbolic layout.
-Questions span a range of topological and functional types, including numeric computations (distances, areas), spatial feasibility (object placement), visibility, and requirement violations.
-Some questions require fine-grained metric reasoning, others test whether a model can respect physical constraints.
-A categorized list of question types is shown in Table 2.
-Each question is generated by filling a parameterized template with layout-specific variables such as object names, measurements, and task-specific contextual information (e.g.
-units for distance, clearance for paths, or which object-types should not occlude visibility).
-Prompts are issued in zero-shot settings, without few-shot examples or role-based instructions.
-Instead, we enforce simple structural markers—such as a required checklist and a final-answer line—to encourage stepwise reasoning.
-To ensure verifiable outputs, each prompt specifies a response schema consisting of a brief structured justification and a final answer line. For example, a distance query is phrased as:
-Prompt: Distance Query If the format, object names, or required inputs are missing, invalid, or inconsistent, the model must return:  *Final answer*: ERROR.
-Otherwise, responses must follow the scheme, for example: Response Schema This structure invokes step-by-step reasoning and each question ends with *Final Answer*: <answer>, enabling robust extraction even when APIs lack native structured-output support.
 Table 1: FloorplanQA question taxonomy.
@@
 View Angle Compute the smallest absolute angle in degrees between the vector from the centroid of the sofa to the centroid of the TV and the global north vector (0, 1) N Metric
-Repositioning Calculate how far the ottoman be moved in the left direction until it touches another object or the wall N Dynamic
+Repositioning Calculate how far the ottoman can be moved in the left direction until it touches another object or the wall N Dynamic
 Max Box Calculate the area in square meters (m²) of the largest rectangle that can fit inside the room N Topology
 Placement Check if a 2m × 3m desk table can fit fully inside the room without overlaps B Topology
-Shortest Path Determine the shortest valid path that maintains a clearance of 15 cm from all other objects, starting from centroid of the stove and ending at the centroid of door S Dynamic
+Shortest Path Determine the shortest valid path that maintains a clearance of 15 cm from all other objects, starting from the centroid of the stove and ending at the centroid of the door S Dynamic
 Visibility Find all objects that intersect the vector from the centroid of the window to the centroid of the fireplace L Topology
-Layouts are represented in a structured JSON format.
-Each entry contains a layout_id, the room_type, and explicit geometric descriptions.
-The room_boundary is stored as a closed polygon, while walls are represented separately as a list of line segments.
-Openings such as windows and doors are included in a dedicated openings field rather than flattened into the object list.
-All furnishings and functional elements (e.g., bed, sofa, table) are stored in the objects list, with each object defined by a labeled polygon.
-In the synthetic data, these polygons are axis-aligned bounding boxes (four points), whereas in HSSD they can exhibit arbitrary shapes and orientations.
-Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents remain unique and stable across prompt construction and answer evaluation.
-Coordinates are expressed in meters in a right-handed 2D Cartesian frame (+x+x to the right, +y+y upward).
-The global origin $(0,0)$ is not fixed to the layout’s lower-left corner and may vary across layouts.
-The prompt used to generate examples according to this schema is in Appendix H.
+FloorplanQA assesses spatial reasoning by presenting models with a single natural language question per symbolic layout.
+Questions span a range of topological and functional types, including numeric computations (distances, areas), spatial feasibility (object placement), visibility, and requirement violations.
+Some questions require fine-grained metric reasoning, others test whether a model can respect physical constraints.
+A categorized list of question types is shown in Table 1.
 We group these questions into three reasoning categories.
@@
 Scalar outputs (N) are scored by relative error with a default tolerance of 2%; for complex area-computation tasks (e.g., Free Space), the tolerance is relaxed to 5%.
-Sequence outputs (S) are evaluated with a Fréchet threshold of 0.6 m, approximating minimal human clearance, and must be valid (collision-free, no overlaps).
+These tolerances are chosen to accommodate minor numerical instability in LLM outputs while remaining strict enough to distinguish correct spatial reasoning from approximation.
+Sensitivity studies supporting these thresholds, including tolerance sweeps showing that they affect only absolute accuracies and not model rankings, are provided in Appendix G.1.
 List outputs (L) are evaluated by set equality.
-Together, the categories and format codes define the taxonomy summarized in Table 2, which reports task coverage and examples for each case.
+Sequence outputs (S) must be valid (collision-free with required clearance) and are additionally scored by Fréchet distance to the reference path.
+We use a threshold of 0.6 m, which tolerates minor deviations in waypoint placement while rejecting qualitatively different routes (e.g., opposite sides of an obstacle).
+Threshold sweeps are reported in Appendix G.2.
+Each question is generated by filling a parameterized template with layout-specific variables such as object names, measurements, and task-specific contextual information (e.g.
+units for distance, clearance for paths, or which object-types should not occlude visibility).
+Prompts are issued in zero-shot settings, without few-shot examples or role-based instructions.
+Instead, we enforce simple structural markers—such as a required checklist and a final-answer line—to encourage stepwise reasoning.
+To ensure verifiable outputs, each prompt specifies a response schema consisting of a brief structured justification and a final answer line.
+Some models expose API-level mechanisms for structured output enforcement.
+We instead encode the same schema as formatting requirements in the prompt for all models to keep the prompting interface fixed across systems.
+For example, a distance query is phrased as:
+Prompt: Distance Query If the format, object names, or required inputs are missing, invalid, or inconsistent, the model must return:  *Final answer*: ERROR.
+Otherwise, responses must follow the scheme, for example: Prompt: Response Schema This structure invokes step-by-step reasoning and each question ends with *Final Answer*: <answer>, enabling robust extraction even when APIs lack native structured-output support.
 3.4 Evaluation Protocol and Scoring
 Each question in FloorplanQA is paired with a reference answer computed directly from the symbolic layout, enabling fully automated and deterministic evaluation of model outputs.
-Depending on the response type, correctness is assessed using numeric comparison with fixed tolerances, string matching, or geometric validation checks.
-For numerical questions (e.g., distances, areas, angles), predictions are accepted if they fall within a relative error threshold.
-Sequence outputs are evaluated for both format and semantic correctness.
-Rather than requiring exact coordinate matches, we validate paths using geometric plausibility conditions of collision avoidance and sufficient proximity to a ground-truth trajectory.
-Deviation thresholds are set conservatively to tolerate minor geometric variations without crediting qualitatively wrong routes.
-To differentiate reasoning failures from extraction or formatting issues, we apply a regex-based parsing pipeline, covering a fixed set of expected answer patterns (e.g., ‘*Final answer*’ tokens in lower case or with surrounding symbols).
-If no answer is produced, or if the extracted content does not match a valid format, we count the response as an error (which is also considered incorrect).
-We provide a detailed breakdown on accuracy in Appendix E.
-In our evaluations, the proportion of invalidly formatted answers is below 1%.
-We also explicitly track cases where no answer is returned due to truncation (API: stop_reason = Token Limit), a failure mode that disproportionately affects reasoning-heavy models.
-To account for this, we report in Appendix I both the percentage of responses truncated by token limits and an adjusted accuracy computed only over valid (non-truncated) answers.
-These adjusted accuracies can be interpreted as an approximate upper bound on overall dataset performance, since with unlimited token budgets models could in principle reach that quality.
+Answer extraction.
+To differentiate reasoning failures from formatting issues, we apply a regex-based parsing pipeline covering expected answer patterns (e.g., ‘*Final answer*’ tokens).
+If no answer is produced or the extracted content does not match a valid format, we count the response as an error.
+The proportion of invalidly formatted answers is below 1%.
+Per-model truncation rates and invalid-format proportions are summarized in Table 7 (Appendix E.3).
+Truncation tracking.
+We track cases where no answer is returned due to token limits (API: stop_reason = Token Limit), a failure mode that disproportionately affects reasoning-heavy models.
+We report both the raw, detailed accuracy breakdown by model, room type, and question type (Appendix E, Tables 5, 6) and the adjusted accuracy computed only over non-truncated responses (Appendix K).
+The adjusted scores serve as an upper bound on performance, since larger token budgets could enable more complete solutions.
+However, truncation occurs primarily on the most challenging questions, so valid-only accuracy may inflate performance and should not be interpreted as a standalone metric.
 3.5 Model Inference Setup
-All models are queried using standard chat-based completion APIs, with prompts constructed as described above.
 We evaluate both large and mid-size models, including reasoning and standard variants.
@@
 These budgets reflect our observations that larger models generate longer intermediate justifications and thus consume more tokens.
-GPT-5 is evaluated under a distinct configuration: its reasoning intensity can’t be disabled, so we run it with reasoning and verbosity set to “low,” with a maximum output length of 4,096 tokens, while GPT-5-mini is run with reasoning and verbosity set to “medium,” with a maximum output length of 8,192 tokens.
-No model receives fine-tuning or prompt adaptation specific to FloorplanQA.
-All prompts are zero-shot, with the system message and output formatting constraints held fixed.
+GPT-5 is evaluated under a distinct configuration, since its reasoning intensity cannot be disabled.
+We run GPT-5 with reasoning enabled and “low” verbosity and limit its output budget to 4,096 tokens, counting both reasoning and generated tokens, while accounting for its higher computational cost.
+GPT-5-mini is evaluated with “medium” reasoning and verbosity and an 8,192-token output budget.
 For each model, we use the default inference configuration provided by its vendor, with temperature set to 0.
-The only exception is GPT-5, for which the temperature cannot be modified and defaults to 1.
+The only exception is GPT-5 family, for which the temperature parameter is not user-configurable and defaults to 1 when reasoning is enabled.
 Each model is evaluated over 1800 generated layouts and 200 layouts from HSSD, with one question from each type posed per layout.
@@
 4 Results
-We evaluated the performance of the model on a dataset of 2,000 layouts, consisting of 600 kitchens, 600 living rooms, 600 bedrooms, and 200 additional layouts from HSSD.
-Each layout is paired with one question sampled from a pool of 8 parameterized templates, filtered by room applicability.
-This yields 16,000 layout–question pairs.
-The full taxonomy is shown in Table 2.
-We evaluate fifteen models, spanning a wide range of parameter scales, architectures, and training regimes.
-The reasoning-oriented models include GPT-5 (OpenAI, 2025a), GPT-OSS-120B (OpenAI, 2025b), DeepSeek-R1-0528 (DeepSeek-AI, 2025), GPT-5-mini, Gemini Flash 2.5 (Google DeepMind, 2025), GPT-OSS-20B, and Qwen3-30B-A3B-Thinking-2507 (Yang et al., 2025).
-The general-purpose (standard) models include Claude Sonnet 4 (Anthropic, 2025), GPT-4.1 (OpenAI et al., 2024), Moonshot Kimi-K2-Instruct (Kimi Team, 2025), Qwen3-Coder-480B-A35B-Instruct, Qwen3-235B-A22B-Instruct-2507, GPT-4.1-mini, Qwen3-30B-A3B-Instruct-2507, and Devstral-Small-2505 (MistralAI, 2025).
-All models are evaluated in identical zero-shot conditions using standardized prompts and serialized layout inputs, as described in Section 3.
-Complete results disaggregated by question type, room type, and model are provided in Appendix E.
+We evaluate fifteen models on the full benchmark: 2,000 layouts (600 kitchens, 600 living rooms, 600 bedrooms, 200 HSSD) with eight questions each, yielding 16,000 layout–question pairs.
+The reasoning-oriented models include GPT-5 (OpenAI, 2025), gpt-oss-120B (Agarwal et al., 2025), DeepSeek-R1-0528 (DeepSeek-AI, 2025), GPT-5-mini, Gemini Flash 2.5 (Google DeepMind, 2025), gpt-oss-20B, and Qwen3-30B-A3B-Thinking-2507 (Yang et al., 2025).
+The general-purpose models include Claude Sonnet 4 (Anthropic, 2025), GPT-4.1 (OpenAI et al., 2023), Moonshot Kimi-K2-Instruct (Kimi Team, 2025), Qwen3-Coder-480B-A35B-Instruct, Qwen3-235B-A22B-Instruct-2507, GPT-4.1-mini, Qwen3-30B-A3B-Instruct-2507, and Devstral-Small-2505 (Mistral AI and All Hands AI, 2025).
+All models are evaluated in identical zero-shot conditions.
+The full taxonomy is shown in Table 1; complete per-model results are provided in Appendix E.
 4.1 Quantitative Results
-We begin by aggregating accuracy across models and question types for both reasoning and general model families.
-Figure 2 summarizes accuracy: the top row shows general models, and the bottom row shows reasoning models.
-In each figure, the left panel summarizes accuracy by model, and the right panel summarizes accuracy by question across room types.
-Kitchens lead across models because overlaps are rare, so most queries are straightforward.
-Scores on HSSD tend to lag behind the other room types; irregular, non-axis-aligned geometry and denser overlap make these layouts more demanding, see Appendix D for detailed statistics.
-Bedrooms and living rooms are mid-tier and nearly equal, lying between kitchens and HSSD.
-Comparing model families, general models struggle when many overlaps must be merged; they often treat object areas independently (no union), which hurts Free Space and Max Box and carries over into path planning.
-Reasoning models handle those cases with unions and rotations better, so they show gains on Free Space and Max Box.
-A practical limitation remains: Gemini Flash 2.5 and DeepSeek-R1 hit token limits on larger layouts, which drops scores, especially in HSSD.
-Task difficulty follows a consistent pattern in terms of accuracy.
-Metric-category questions, such as Pair Distance and View Angle, achieve the highest accuracies.
-Repositioning, Visibility, and Placement yield mid-range accuracies.
-Max Box and Free Space benefit most from reasoning-oriented models, yet their accuracies remain low and are comparable in difficulty to Shortest Path.
-Accuracy decreases as object count and overlap density increase, most notably on HSSD layouts.
-Detailed analysis on each question category is provided in Appendix G
+Room type difficulty.
+Figure 3 summarizes accuracy across models and question types, with general models in the top row and reasoning models in the bottom row.
+Kitchens yield the highest accuracy due to simpler geometry and fewer object overlaps: on average, each kitchen layout contains 0.52 overlapping object pairs (excluding rugs), compared to 1.52–1.82 overlapping pairs in bedrooms and living rooms (Appendix D, Table 4).
+HSSD layouts are more complex due to non-axis-aligned polygons and substantially more overlap (4.39 overlapping pairs per layout, excluding rugs), which increases the difficulty of union, clearance, and collision reasoning.
+Bedrooms and living rooms fall in between.
+Model rankings (left panel) mainly remain consistent across synthetic and HSSD layouts, suggesting that synthetic layouts test the same spatial reasoning skills as human-authored scenes.
 Figure 3: Accuracy of general (top) and reasoning (bottom) models.
-Left: by model.
-Right: by question.
-Each column corresponds to a specific room type represented in our dataset: Kitchens, Living Rooms, Bedrooms (synthetic subsets), plus HSSD layouts.
-4.2 Ablation
-To evaluate the robustness of layout interpretation under alternate encodings, we perform a format ablation that replaces the standard JSON layout with a semantically equivalent XML version.
-This substitution preserves all geometric and object-level content while modifying only the syntax and structural serialization.
-The prompt level and semantic ablations are described in Appendix, Section F.
-To avoid recomputing the full suite, we focus on the most variance-sensitive question types for HSSD layouts, as shown in Appendix F (Figs.
-5 and 6): View Angle, Visibility, and one task from a different category—Repositioning (Dynamic).
-We apply each input-format ablation to three representative models: two reasoning models (GPT-OSS-120B, GPT-OSS-20B) and one large general model (Qwen3-235B-A22B) for the same subset of HSSD layouts.
-As shown in Table 3, accuracy is largely stable in both formats for most categories; the GPT-OSS models change minimally, while Qwen shows modest fluctuations.
-This suggests that these models encode layout semantics in a manner that is relatively invariant to low-level representation details.
-Table 3: Accuracy using JSON vs XML layout encoding. Each cell shows performance on the original JSON representation and its equivalent XML rendering.
-Model Repositioning View Angle Visibility
-GPT-OSS-120B 60.5 → 59.0 74.0 → 74.0 70.0 → 72.0
-GPT-OSS-20B 40.0 → 39.0 37.5 → 38.5 45.5 → 43.5
-Qwen3-235B-A22B 39.0 → 42.0 50.5 → 54.0 70.5 → 65.5
+The left panel summarizes accuracy by model, averaged across all question types.
+The right panel summarizes accuracy by question, averaged across all models within the respective general or reasoning family.
+Each column corresponds to a specific subset in our dataset: HSSD layouts, plus our Kitchens, Living Rooms, and Bedrooms.
+Task difficulty hierarchy.
+Three distinct tiers emerge from analysing the right panel of Figure 3.
+Metric tasks (Pair Distance, View Angle) achieve 75–95% averaged accuracy on synthetic layouts, where objects are axis-aligned boxes.
+On HSSD layouts, accuracy drops to 35–60%: models frequently fail to compute centroids of complex polygons, making computational errors when applying the shoelace formula (see Appendix I.1, I.4), compared to simple averaging for 4-point boxes.
+Constraint reasoning tasks (Placement, Visibility, Repositioning) yield 60–85% averaged accuracy; performance degrades when multiple geometric constraints must be jointly verified.
+For Placement and Visibility, baseline-guessing is ruled out by balanced ground truth (49.9% True) and low empty-list frequency (¡20%), respectively (see Appendix I.5, I.7, I.2).
+Optimization and planning tasks (Free Space, Max Box, Shortest Path) achieve only 5–45% averaged accuracy—tasks requiring geometric unions, search, or multi-step collision-free reasoning remain largely unsolved.
+One exception is Free Space on Kitchens, yielding over 90% accuracy due to minimal object overlaps; in contrast, other layouts require computing unions of overlapping object pairs (e.g., TV and TV stand, nightstand and lamp) to correctly calculate non-occupied area (see Appendix I.3, I.8, I.6).
+Reasoning vs.
+general models.
+Reasoning-focused models show notable improvements on geometric union tasks, gaining +10–40% on Free Space and Max Box compared to general models on generated layouts.
+This suggests that extended inference computation helps with complex geometric operations.
+However, gains on path planning remain modest.
+A practical limitation affects some reasoning models: Gemini Flash 2.5 and DeepSeek-R1 frequently hit token limits on larger layouts (30–50% truncation; see Table 7), reducing their reported scores.
+Valid-only accuracy (Appendix K, Tables 13 and 15) shows these models perform competitively when responses are completed successfully.
+4.2 Ablations
+Input format.
+To test robustness to layout encoding, we replaced JSON with semantically equivalent XML for a subset of HSSD layouts across three tasks (View Angle, Visibility, Repositioning) and three models (gpt-oss-120B, gpt-oss-20B, Qwen3-235B).
+Accuracy remained stable across formats ($\pm$3 percentage points; see Appendix H.1, Table 9), suggesting models encode layout semantics independently of serialization syntax.
+Prompt and semantic variation.
+We tested robustness to object reference substitution (different object names, same template) and semantic perturbation (permuted object labels, identical geometry).
+Object reference substitution shows minimal effect on accuracy (Appendix H.2).
+Semantic perturbation reveals a task-dependent pattern: View Angle and Visibility remain stable when object labels are swapped, but Repositioning accuracy drops sharply (e.g., 60.5% → 40.0% for gpt-oss-120B).
+We hypothesize that for movement-related tasks, models rely on linguistic associations rather than geometric properties—for instance, when asked to move a ”sofa” whose label has been swapped with ”chair,” models sometimes search for furniture that semantically resembles a sofa rather than using the object’s actual geometry (Appendix H.3).
+Tool augmentation.
+To isolate whether failures stem from arithmetic errors or deeper reasoning limitations, we evaluated GPT-4.1 with an integrated Python Code Interpreter.
+Table 2 shows results on HSSD layouts.
+Additional results for generated layouts and GPT-4.1-mini model can be found in Appendix F, Table 8.
+Table 2: Accuracy on HSSD layouts with and without Python Code Interpreter (GPT-4.1).
+Task Raw Tools $\Delta$
+Pair Distance 56.0 99.0 +43.0
+View Angle 55.0 96.0 +41.0
+Visibility 46.5 86.5 +40.0
+Repositioning 47.0 83.5 +36.5
+Placement 64.5 95.0 +30.5
+Free Space 16.0 44.0 +28.0
+Max Box 7.5 3.0 -4.5
+Shortest Path 25.0 12.5 -12.5
+Tool augmentation substantially improves arithmetic-heavy tasks: Pair Distance (+43 points), View Angle (+41 points), Visibility (+40 points).
+However, performance on optimization and planning tasks decreases: Max Box (-4.5 points) and Shortest Path (-12.5 points).
+Inspection of failure cases (Figure 8) reveals that tool-augmented models generate Python code with incorrect algorithms—treating boundary contact as collision, or performing incomplete search over rectangle orientations.
+This indicates that the bottleneck for complex spatial tasks is algorithmic reasoning about spatial constraints, not numerical precision.
+Vision-language input.
+We also evaluated VLM inputs using rendered floorplan images alongside symbolic layouts.
+Results (Appendix F, Table 8) show selective improvements on synthetic layouts but no consistent advantage over symbolic input alone, suggesting structured representations already provide a strong baseline for these tasks.
 5 Conclusion
-We introduced FloorplanQA, a benchmark for spatial reasoning over symbolic 2D layouts aligned with architecture and robotics practice.
-We evaluated 15 language models (8 general, 7 reasoning) on 2,000 layouts (1,800 synthetic; 200 semi–real HSSD) across tasks ranging from metric queries to visibility, placement, and shortest path with clearance.
-Empirically, metric and simple visibility queries are reliable; kitchens score highest because overlaps are rare.
-HSSD layouts are more demanding—irregular, non–axis-aligned shapes and dense overlaps expose weaknesses such as centroid miscalculation and missed unions.
-Reasoning models improve notably on Free Space and Max Box by handling overlaps and rotations more consistently, while general models often subtract object areas independently and fail under heavy overlap.
-Shortest Path is sufficiently challenging because it requires multiple correct steps (clearance buffering, collision checks, and path search), where errors compound.
-These results indicate that current LLMs lack sufficiently robust internal geometric representations for complex spatial inference.
+We introduced FloorplanQA, a benchmark for evaluating spatial reasoning in LLMs over symbolic 2D layouts representative of architectural and planning contexts.
+We evaluated 15 language models (8 general, 7 reasoning) on 2,000 layouts (1,800 synthetic; 200 semi–real HSSD) across 8 types of questions spanning metric queries, visibility, placement, and collision-free path estimation.
+Key findings.
+Our evaluation (Section 4) reveals a consistent capability gap: while models reliably compute some geometric properties (distances, angles), they struggle with constraint satisfaction, geometric unions, and multi-step spatial planning.
+Three systematic failure modes emerge: (1) Models frequently behave as if object areas were independent rather than unioned (that is, they compute the sum of areas rather than the union), which aligns with low accuracy on Free Space and Max Box (Appendix D, Fig.
+17).
+(2) Performance drops on HSSD layouts with non-axis-aligned, high-vertex polygons, and qualitative analysis indicates that many Distance/Angle errors originate in centroid computation (Appendix I.1, I.4).
+(3) Shortest Path remains difficult, especially in cluttered scenes where buffering and obstacle interactions narrow corridors (Appendix I.6); sensitivity analysis shows this across Fréchet and clearance thresholds (Appendix G.2, Fig.
+12).
+Reasoning-focused models show notable improvements on geometric union tasks, demonstrating that extended inference computation helps with complex geometric operations.
+However, gains on planning tasks remain modest, suggesting that the bottleneck lies in spatial representation rather than purely computational depth.
+Extended capabilities.
+We also evaluated tool augmentation and vision-language input (Appendix F).
+Python Code Interpreter yields strong gains on arithmetic-heavy tasks (plus over 30–40 percentage points on Distance/Angle) but limited benefit on planning tasks, indicating failures stem from spatial reasoning rather than calculation errors.
+Rendered floorplan images provide selective improvements on synthetic layouts but do not consistently outperform symbolic input, suggesting symbolic representations already provide strong baselines.
+Future directions.
 Two complementary directions follow.
-Near term: hybridize with external geometric solvers or symbolic planning modules—set operations (unions/differences), centroid via shoelace, clearance buffering, oriented rectangle search, and A* path planning—to compensate for the models’ weaknesses in collision avoidance and clearance reasoning.
-Longer term: train with explicit spatial constraints and harder distributions (irregular, overlap–heavy layouts), and include constraint–violation exemplars and geometry-aware objectives so models learn to maintain coherence under rotation, clearance, and union operations in design-oriented tasks.
+Near term: hybridize with external geometric solvers or symbolic planning modules, including set operations (unions/differences), centroid via shoelace, clearance buffering, oriented rectangle search, and A* path planning, to compensate for models’ weaknesses in collision avoidance and clearance reasoning.
+Longer term: train with explicit spatial constraints and harder distributions (irregular, overlap-heavy layouts), and include constraint-violation exemplars and geometry-aware objectives so models learn to maintain coherence under rotation, clearance, and union operations.
+Beyond these directions, multi-step interaction (e.g., asking an agent to verify or revise its solution using a rendered floorplan) may further improve performance.
+FloorplanQA provides a standalone benchmark for measuring progress on spatial reasoning in layout-based tasks, with applications ranging from architectural design to indoor planning.
 Acknowledgments
-The work is supported by funding from King Abdullah University of Science and Technology (KAUST)—Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program
+The work is supported by funding from King Abdullah University of Science and Technology (KAUST)—Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program.
 Appendix A Technical Appendix and Supplementary Material for FloorplanQA
-This appendix details the dataset generation pipeline, dataset statistics, and the full evaluation of FloorplanQA across reasoning and general models.
-It also includes ablation studies, analysis of failure cases, and the exact prompts used for data generation and question asking.
+This appendix provides comprehensive technical details supporting the main paper.
+Section B describes synthetic layout generation and Section C covers HSSD extraction procedures.
+Section D provides detailed statistics on room complexity, object counts, and overlap distributions.
+Section E presents the complete evaluation breakdown by model, room type, and question category, including truncation analysis.
+Section F presents extended experiments with tool-augmented reasoning (Python Code Interpreter) and vision-language models using rendered floorplans.
+Section G provides sensitivity analyses for evaluation tolerances.
+Section H presents ablation studies examining input format, object reference substitution, and semantic perturbation.
+Section I contains qualitative failure analysis with annotated examples identifying common error patterns across question types.
+Section J includes full prompt templates for data generation and question asking.
+Section K provides supplementary accuracy tables.
 All artifacts (code, generation prompts, evaluation scripts, and visualization utilities) are included with the submission as supplementary material.
@@
 Table 3: Shape and size distribution across generated layouts for each room type.
-Room Type
-Shape Distribution
-(Rect / L-shaped / Open)
-Size Categories (in m2)
-(Small / Medium / Large)
+Room Type Shape Distribution (Rect / L-shaped / Open) Size Categories (in m2) (Small / Medium / Large)
 Kitchen 40% / 40% / 20% $\leq$7 / 7–18 / >>18
@@
 During 2D projection, we correct mislabeled windows and doors, snap nearly collinear edges, and resolve self-intersections.
-To smooth rectilinear artifacts and irregular edges in box-like footprints, we apply alpha-shape–based convex hulls, then recompute centroids and areas on the cleaned geometry.
+To smooth rectilinear artifacts and remove redundant vertices in box-like footprints, we apply Douglas-Peucker polygon simplification with tolerance $\epsilon=0.01$ m.
+This preserves the polygon’s overall shape while reducing token overhead.
+All ground-truth answers (centroids, areas, distances) are computed from this cleaned geometry.
 Figure 4: HSSD layout selection: comparison of the original (left) and simplified (right) 2D layouts.
@@
 Table 4 reports summary statistics for the layouts.
+Overlaps counts pairs of objects whose bounding boxes intersect.
+We report two versions: excluding rugs (rugs are ignored when counting overlaps) and including rugs (all overlaps counted).
+Object density is computed as the total number of objects divided by room area (objects/m²).
+Vertices per layout includes all polygon vertices from room objects.
 Kitchens are typically the smallest spaces, with relatively few objects and overlaps, but a high density due to their compact geometry.
@@
 Avg. # of Objects 10.35 10.76 11.69 12.20
-Avg. # of Overlaps 0.52 1.82 1.52 4.39
+Avg. # of Overlaps (excl. rugs) 0.52 1.82 1.52 4.39
+Avg. # of Overlaps (incl. rugs) 0.52 6.25 6.2 6.65
 Avg. Object Density 0.95 0.66 0.57 0.83
@@
 E.1 Detailed Accuracy by full dataset
-Tables 6 and 7 report the full per-model accuracy matrix for the base and reasoning model groups, respectively.
+Tables 5 and 6 report the full per-model accuracy matrix for the base and reasoning model groups, respectively.
 Each table covers nine question categories across four room types, resulting in a total of 36 rows.
-The path task is assessed using two complementary metrics: validity and Fréchet Distance.
+The Shortest path task is assessed using two complementary metrics: validity and Fréchet Distance.
 Together, these question types span a range of challenges, covering both reasoning-heavy and functionally grounded tasks.
@@
 For every model, question type accuracy is computed over a fixed set of 600 synthetic layouts (Kitchens, Living Rooms, Bedrooms) and 200 HSSD layouts, ensuring that results are directly comparable across models.
-E.2 Token-Limit Analysis and Valid-Only Accuracy
-In addition to overall accuracy, we analyze two complementary aspects of model performance.
-First, Tables 10 and 12 quantify the fraction of responses that were terminated due to the TOKEN LIMIT stop reason.
-This failure mode is particularly relevant for reasoning-oriented models, which often generate longer chain-of-thought outputs.
-Second, Tables 11 and 13 report accuracy over completed answers only, excluding truncated outputs (e.g., token-limit terminations).
-This metric isolates models’ reasoning performance on successfully produced, well-formed responses.
-Together, these analyses complement the full accuracy tables by disentangling reasoning failures from generation truncation and formatting issues.
 Table 5: Question-level accuracy on full dataset by room type for standard models.
-Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
+Question Room [Uncaptioned image] Claude
+Sonnet-4
+[Uncaptioned image] GPT-4.1
+[Uncaptioned image] Kimi-K2
+Instruct
+[Uncaptioned image] Qwen3
+Coder-480B
+[Uncaptioned image] Qwen3
+235B
+[Uncaptioned image] GPT-4.1
+mini
+[Uncaptioned image] Qwen3
+30B
+[Uncaptioned image] Devstral
+Small
 Pair distance K 99.8 96.5 95.7 96.8 99.2 90.7 89.5 58.8
@@
 HSSD 35.0 16.0 24.0 22.0 17.0 15.5 7.5 6.5
-Visibility K 87.7 63.2 54.5 67.0 98.3 90.2 91.5 18.5
-LR 81.7 52.7 43.3 52.2 98.3 86.8 88.7 10.0
-B 74.8 57.5 41.0 54.0 96.7 86.3 86.3 10.8
-HSSD 46.5 20.0 22.5 26.5 70.5 52.0 45.5 9.0
+Visibility K 63.2 87.7 54.5 67.0 98.3 90.2 91.5 18.5
+LR 52.7 81.7 43.3 52.2 98.3 86.8 88.7 10.0
+B 57.5 74.8 41.0 54.0 96.7 86.3 86.3 10.8
+HSSD 20.0 46.5 22.5 26.5 70.5 52.0 45.5 9.0
 View angle K 92.0 95.3 69.8 78.8 97.0 95.8 74.7 49.7
@@
 HSSD 15.5 22.5 18.0 8.0 20.0 12.5 2.5 11.5
+E.2 Token-Limit Analysis and Valid-Only Accuracy
+In addition to overall accuracy, we analyze two complementary aspects of model performance.
+First, Tables 12 and 14 quantify the fraction of responses that were terminated due to the TOKEN LIMIT stop reason.
+This failure mode is particularly relevant for reasoning-oriented models, which often generate longer chain-of-thought outputs.
+Second, Tables 13 and 15 report accuracy over completed answers only, excluding truncated outputs (e.g., token-limit terminations).
+This metric isolates models’ reasoning performance on successfully produced, well-formed responses.
+Together, these analyses complement the full accuracy tables by disentangling reasoning failures from generation truncation and formatting issues.
 Table 6: Question-level accuracy on full dataset by room type for reasoning models.
-Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
+Question Room [Uncaptioned image] GPT-5
+[Uncaptioned image] gpt-oss
+120b
+[Uncaptioned image] DeepSeek
+R1-0528
+[Uncaptioned image] Gemini
+Flash 2.5
+[Uncaptioned image] GPT-5
+mini-2025
+[Uncaptioned image] gpt-oss
+20b
+[Uncaptioned image] Qwen3
+30B Think.
 Pair distance K 99.8 99.3 98.0 96.3 100.0 94.2 97.7
@@
 HSSD 22.5 26.5 2.5 0.0 12.0 14.0 1.0
-Appendix F Additional Ablations and Analysis
-To further verify the robustness and generalizability of our evaluation setup, we conducted supplementary analyses beyond the main accuracy tables.
+E.3 Aggregated Truncation Rates
+Table 7 reports aggregated truncation and error-type statistics for each model, computed by averaging outcomes across all room types and all question categories in the full dataset.
+A response is counted as truncated when it terminates with the TOKEN LIMIT stop reason.
+We additionally report the fraction of invalid-format responses that could not be parsed into a valid answer under our evaluation protocol.
+For completeness, Table 7 also includes the proportions of wrong and correct responses, as well as an alternative accuracy (% alt), computed by extracting the last numeric value or the final list from each model’s output.
+This alternative metric serves purely as a robustness check on the parsing and evaluation pipeline.
+Table 7: Error-type distribution and accuracy on the full dataset.
+Values are aggregated for each model by averaging across all room types and all questions.
+Truncation (% trunc.) quantifies the fraction of responses that were terminated due to the TOKEN LIMIT stop reason.
+Invalid-format (% invalid) denotes responses that could not be parsed into a valid answer.
+Alternative accuracy (% alt) reports accuracy computed using the last numeric value or the last list in the model’s output.
+Model % trunc. % invalid % wrong % correct % alt
+gpt-5 3.86 0.27 30.42 65.45 65.45
+gpt-oss-120b 2.05 1.05 21.41 75.49 75.53
+DeepSeek-R1 31.31 0.78 17.26 50.65 50.65
+gpt-5-mini 15.50 1.34 12.54 70.62 70.62
+Gemini Flash 2.5 50.58 0.13 10.14 39.15 39.15
+gpt-oss-20b 16.94 1.18 21.70 60.18 60.31
+Qwen3-30B Think 38.95 0.09 16.52 44.44 44.44
+claude-sonnet-4 0.05 0.07 41.09 58.80 58.80
+gpt-4.1 0.14 0.09 41.52 58.25 58.25
+Kimi-K2 0.15 0.03 52.04 47.78 47.78
+Qwen Coder-480B 0.02 0.09 49.40 50.48 50.48
+Qwen-235B 13.64 0.13 20.31 65.92 65.98
+gpt-4.1-mini 0.67 0.10 42.18 57.05 57.05
+Qwen-30B Instr 6.85 3.09 38.47 51.59 51.84
+Devstral Small 2.75 1.56 65.64 30.05 30.07
+E.4 Radar summaries
+For readability, we also provide radar visualizations that summarize accuracy and variance across models and question types, complementing the main tables.
 General models.
@@
 By question type, Repositioning and Placement are handled reliably, whereas Max Box and Shortest Path remain the most difficult as well, with high variance across rooms.
-Figure 5: Radar plots for general models, showing mean and standard deviation of accuracy across (top) models and (bottom) question type.
-Figure 6: Radar plots for reasoning models, showing mean and standard deviation of accuracy across (top) models and (bottom) question type.
-Prompt Sensitivity Ablation
-To assess model robustness to prompt variation, we conducted an ablation in which each question was regenerated using the same template but with alternate object references.
-For example, a question originally referring to a ”sofa” might instead use a ”bookshelf” in the regenerated version.
-This allowed us to evaluate whether model performance remains stable under changes in object content while preserving linguistic structure.
-As shown in Table 9, accuracy under prompt regeneration is broadly stable; larger models change little, and smaller ones fluctuate modestly.
-We observe somewhat higher sensitivity for Repositioning: paraphrases can implicitly select different target objects or motion directions, occasionally introducing additional complexity or non-movable cases.
-Overall, the evaluation appears robust to prompt-level variations.
+Figure 6: Radar plots for general models.
+Top: mean accuracy (left) and standard deviation (right) across models, grouped by room type.
+Bottom: mean accuracy (left) and standard deviation (right) across question types.
+Figure 7: Radar plots for reasoning models.
+Top: mean accuracy (left) and standard deviation (right) across models, grouped by room type.
+Bottom: mean accuracy (left) and standard deviation (right) across question types.
+Appendix F Tools and VLM experiments
+To complement the text-only benchmark, we additionally evaluate (i) tool-augmented reasoning using an integrated Python interpreter and (ii) vision-language models (VLMs) using rendered floorplan images.
+These studies measure the potential benefit of external computation and visual input while keeping the task definitions and scoring identical to the main benchmark.
+Tool-augmented setting.
+We enable the Code Interpreter tool for GPT-4.1 and GPT-4.1-mini, allowing the model to invoke a Python sandbox during inference.
+The model is instructed to use Python whenever numeric computation is needed (e.g., distances, angles, polygon centroids or areas), and to output a final numeric answer in the same format as the raw setting.
+While the interpreter improves numeric precision, models can still fail on complex tasks either due to incorrect spatial reasoning or because the generated code is incomplete or erroneous.
+Results are reported in Table 8.
+VLM setting and renderings.
+We also evaluate VLM inputs by providing a top-down rendered floorplan image (720×720 pixels) alongside both the text question and the complete JSON layout.
+This multimodal setting allows models to leverage visual rendering while retaining access to precise symbolic coordinates.
+We use two controlled rendering styles derived from the same symbolic layout: (i) a minimal contour-based rendering, where room and object polygons are shown as boxes with text labels (similar to Fig.
+4, right), and (ii) an icon-based rendering, where objects are depicted with furniture icons instead of bare contours (similar to Fig.
+2).
+Because HSSD contains a wider variety of object categories than our current icon library, icon-based renderings are used only for synthetic layouts, while HSSD layouts use the contour style.
+VLM results under these input conditions are summarized in Table 8.
+Table 8 shows that Code Interpreter substantially improves metric-dominated scalar tasks (distance, angle, visibility, placement), but offers limited benefit on planning-heavy tasks (Max Box, Shortest Path), indicating that remaining errors are primarily spatial rather than arithmetic.
+VLM inputs provide selective gains, mainly on generated layouts and for object-fit or metric cues, while not consistently improving global performance.
+Table 8: Task accuracy for gpt-4.1-2025-04-14 and gpt-4.1-mini-2025-04-14 under three settings: raw text-only input (Raw), tool-augmented reasoning with a Python interpreter (Tools), and vision-language input using rendered floorplans (Img; Icons).
+HSSD columns report performance on 200 human-designed scenes.
+Generated columns report performance on 200 synthetic scenes (50 kitchens, 75 living rooms, and 75 bedrooms).
+Cell colors indicate gains (green) or drops (red) relative to the corresponding Raw setting for the same model and task.
+HSSD Generated
+Question GPT-4.1 GPT-4.1-mini GPT-4.1 GPT-4.1-mini
+Raw Tools Img Raw Tools Img Raw Tools Img Icons Raw Tools Img Icons
+Pair distance 56 99 60.5 37 98 51.5 95.5 99.5 99 99.5 86.5 98.5 94.5 99.5
+Placement 64.5 95 70.5 76.5 87.5 78.5 65 92.5 61.5 71.5 76 95.5 70.5 81
+Repositioning 47 83.5 44.5 40.5 68 37 56.5 48 42 39 70 59.5 45 50.5
+Free space 16 44 21.5 15.5 42 26 41 28.5 55 45 24.5 27.5 26 22.5
+Visibility 46.5 86.5 36.5 52 70 37 80 89 61.5 56.5 88 71.5 72.5 68
+View angle 55 96 63.5 46 93 54 92 99 93.5 95.5 91.5 97 92.5 92.5
+Max box 7.5 3 4 4.5 3.5 4 12 11.5 12 10.5 11 8 7.5 9.5
+Shortest path (valid) 25 12.5 22.5 15 14 16.5 51 34 46 52 46.5 24.5 44 45
+Shortest path (Fréchet) 22.5 12.5 23.5 12.5 9.5 13.5 40.5 31 44 45 24.5 21 33 27
+As illustrative failure cases, we also observe that tool augmentation does not guarantee correctness for tasks that require precise geometric reasoning.
+In Max Box (Fig.
+8(a)), the model-generated Python program performs an approximate rotated/grid search and returns a suboptimal rectangle (red, $5.57\,\mathrm{m}^{2}$), despite a larger valid axis-aligned solution existing (green, $7.57\,\mathrm{m}^{2}$).
+In a second example from Repositioning ( Fig.
+8(b)), the ground-truth maximum downward displacement of the dishwasher is $1.96$ m, but the model’s tool-based code predicts zero movement.
+Inspection shows that the program treats a shared boundary (the dishwasher touching another object or wall) as a collision, and therefore incorrectly concludes that no valid motion is possible for this concrete task.
+These cases highlight that tools reduce arithmetic imprecision, but complex tasks may still fail due to imperfect model-written algorithms and sensitivity to geometric edge cases (e.g., boundary contact vs.
+overlap).
+(a) Max Box tool failure.
+The model-written code returns a suboptimal rotated rectangle (red, $5.57\,\mathrm{m}^{2}$) despite a larger valid axis-aligned rectangle existing (green, $7.57\,\mathrm{m}^{2}$).
+(b) Repositioning tool failure. The true maximum downward move is $1.96$ m, but the model’s code outputs 0 by treating boundary contact as collision.
+Figure 8: Representative failure cases in the tool-augmented setting.
+While the python interpreter improves numeric precision, complex tasks can still fail due to imperfect model-written code and geometric edge cases.
+Appendix G Sensitivity studies for numeric tolerances and path thresholds
+G.1 Sensitivity studies for numeric tolerances.
+We examine the robustness of our numeric evaluation tolerances for both scalar and area-based question types.
+For clarity and cost control, we conduct these sensitivity analyses on a representative subset of six models (three reasoning-focused and three general-purpose).
+For scalar metrics (e.g., Pair distance, View angle, Repositioning, and Max Box), Figure 9 shows that relative errors are sharply concentrated near zero across models, with a clear knee before $\varepsilon_{\mathrm{rel}}=2\%$ (red dashed line).
+The accompanying tolerance sweep in Figure 11 (left) demonstrates smooth, monotone accuracy gains while changing tolerance from $0.5\%$ to $5\%$ without changing model rankings, indicating that the selection of $2\%$ lies in a stable regime rather than being outcome-sensitive.
+Figure 9: Aggregated relative-error distributions for scalar metrics across models.
+Left: CDF of relative error; the dashed line marks $\varepsilon_{\mathrm{rel}}=2\%$.
+Right: Histogram showing a strong peak near zero and a thin long tail.
+For free-space (area) questions, Figure 10 shows a broader, heavier-tailed error distribution; nevertheless, $5\%$ tolerance lies near saturation for higher-performing models while remaining strict for weaker ones.
+The sweep in Figure 11 (right) confirms that relaxing area tolerance from $1\%$ to $10\%$ yields gradual improvements and preserves qualitative conclusions.
+Figure 10: Aggregated relative-error distributions for free-space (area) questions across models.
+Left: CDF of relative error; the dashed line marks $\varepsilon_{\mathrm{rel}}=5\%$.
+Right: Histogram showing broader, heavier-tailed errors compared to scalar tasks.
+Overall, these results justify our use of a $2\%$ tolerance for scalar outputs and a $5\%$ tolerance for complex area computations.
+Figure 11: Tolerance sweeps for scalar (left) and free-space area (right) tasks.
+Aggregated accuracy increases smoothly as tolerance is relaxed (0.5–5% for scalars; 1–10% for areas), while model rankings remain unchanged.
+G.2 Sensitivity studies for path thresholds
+We analyze robustness of shortest-path scoring with respect to the Fréchet threshold $\tau$.
+Figure 12 shows that accuracy increases smoothly as $\tau$ is relaxed, and model rankings remain almost stable across the sweep.
+We therefore choose $\tau=0.6$ m as deviations within roughly 0.6 m correspond to paths that remain traversable for a person and allow minor alternate routes without accepting qualitatively different solutions.
+In addition, path validity requires collision-free traversal under a clearance buffer of 0.15 m.
+This value represents a minimal safety margin; larger buffers (e.g., 0.3 m) would incorrectly invalidate many feasible paths in compact rooms, particularly around $\sim$0.6 m-scale kitchen utilities, and would over-penalize narrow but realistic layouts.
+Figure 12: Shortest-path accuracy as a function of Fréchet tolerance $\tau$.
+Aggregated accuracy rises smoothly with increasing tolerance, while model ordering changes slightly.
+We use $\tau=0.6$ m as a moderate, human-scale tolerance.
+Appendix H Ablations
+H.1 Format Sensitivity Ablation
+To evaluate the robustness of layout interpretation under alternate encodings, we replaced the standard JSON layout with a semantically equivalent XML version.
+This preserves all geometric and object-level content while changing only syntax.
+We select three tasks that show variance across HSSD layouts: View Angle, Visibility, and Repositioning.
+These span different reasoning categories (Metric, Topology, Dynamic) and have moderate baseline accuracy, where format changes are more meaningful than on near-floor tasks like Max Box.
+We test three models: two reasoning (gpt-oss-120B, gpt-oss-20B) and one general (Qwen3-235B-A22B).
+As shown in Table 9, accuracy stays stable across formats ($\pm$3 points), suggesting models encode layout semantics independently of serialization syntax.
+Table 9: Accuracy using JSON vs XML layout encoding. Each cell shows performance on the original JSON representation and its equivalent XML rendering.
+Model Repositioning View Angle Visibility
+gpt-oss-120B 60.5 → 59.0 74.0 → 74.0 70.0 → 72.0
+gpt-oss-20B 40.0 → 39.0 37.5 → 38.5 45.5 → 43.5
+Qwen3-235B-A22B 39.0 → 42.0 50.5 → 54.0 70.5 → 65.5
+H.2 Object Reference Substitution Ablation
+To assess model robustness to object reference substitution, we conducted an ablation in which each question was regenerated using the same template but with different object references.
+For example, a question originally referring to a ”sofa” might instead refer to a ”bookshelf” in the regenerated version.
+This allowed us to evaluate whether model performance remains stable when object content changes while linguistic structure is preserved.
+As shown in Table 10, accuracy under object reference substitution is broadly stable; larger models change little, and smaller ones fluctuate modestly.
+We observe somewhat higher sensitivity for Repositioning: substitutions can implicitly select different target objects or motion directions, occasionally introducing additional complexity or non-movable cases.
+Overall, the evaluation appears robust to object reference substitution.
 Table 10: Accuracy under prompt variation. Each cell displays performance on the original prompt, followed by a regenerated version with alternative object references.
@@
 Qwen3-235B-A22B 39.0 → 49.0 50.5 → 48.0 70.5 → 74.0
-Semantic Ablation
+H.3 Semantic Ablation
 To evaluate the model’s reliance on semantic object labels rather than purely geometric reasoning, we conduct a semantic ablation experiment.
@@
 This suggests that large language models may entangle linguistic priors with geometric inference when tasks involve physical movement or interaction.
-Table 9: Accuracy under prompt variation. Each cell displays performance on the original prompt, followed by a regenerated version with alternative object references.
+Table 11: Accuracy under semantic ablation. Each cell displays performance on the original prompt, followed by a version with permuted object labels (same geometry, different semantics).
 Model Repositioning View Angle Visibility
@@
 Qwen3-235B-A22B 39.0 → 37.0 50.5 → 43.0 70.5 → 74.0
-Appendix G Case Studies by Question Type
+Appendix I Case Studies by Question Type
 To better understand the sources of model failure, we conducted a qualitative analysis of representative examples from the benchmark.
@@
 By examining both correct and incorrect outputs, we aim to identify common failure patterns and reasoning bottlenecks across different architectures.
-G.1 Pair Distance
+I.1 Pair Distance
 Task Definition
@@
 This error does not depend on polygon complexity (number of vertices).
-G.2 Repositioning
+I.2 Repositioning
 Task Definition
@@
 Models often fall back to axis-aligned bounding boxes (discarding shape orientation) or omit the obstacle-union step, which leads to systematic over- or underestimation of feasible travel distances in any direction (left, right, up, or down).
-G.3 Free Space
+I.3 Free Space
 Task Definition
@@
 Ground-Truth Computation
-We compute the free floor area using geometric operations provided by Shapely (Gillies et al., 2007).
+We compute the free floor area using geometric operations provided by Shapely (Gillies and others, 2007).
 All object polygons within the room are merged using unary_union to correctly handle overlaps.
@@
 Accuracy declines sharply as object count and overlap increase, reflecting the model’s difficulty in handling overlapping geometries.
-G.4 View Angle
+I.4 View Angle
 Task Definition
@@
 (2) Form the displacement vector $\mathbf{d}=\mathbf{c}_{v}-\mathbf{c}_{s}$ and its unit vector $\hat{\mathbf{d}}=\dfrac{\mathbf{d}}{\lVert\mathbf{d}\rVert}$.
-(3) Let the global north vector be $\mathbf{n}=(0,1)$ (already unit length). Compute the cosine via $\cos\theta=\mathrm{clip}\!\big(\hat{\mathbf{d}}\cdot\mathbf{n},\,-1,\,1\big)$.
-(4) Convert to degrees and take the smallest absolute angle: $\theta=\arccos(\cos\theta)\cdot\dfrac{180}{\pi}\ \in[0^{\circ},180^{\circ}]$.
+(3) Let the global north vector be $\mathbf{n}=(0,1)$ (already unit length). Compute the clipped dot product: $c=\mathrm{clip}\!\big(\hat{\mathbf{d}}\cdot\mathbf{n},\,-1,\,1\big)$.
+(4) Convert to degrees: $\theta=\arccos(c)\cdot\dfrac{180}{\pi}\ \in[0^{\circ},180^{\circ}]$.
 A prediction is correct if it is within $2\%$ of the ground-truth angle $\theta$.
@@
 On synthetic layouts (4-point, axis-aligned boxes), centroids are trivial, and this issue does not appear.
-G.5 Placement
+I.5 Placement
 Task Definition
@@
 Models often mis-handle overlap checks under rotation and falsely report feasibility/infeasibility due to incorrect intersection computations.
-G.6 Shortest Path
+I.6 Shortest Path
 Task Definition
@@
 Models often mishandle overlaps, producing paths that cut through obstacles or declaring no path when one exists.
-G.7 Visibility
+I.7 Visibility
 Task Definition
@@
 On HSSD layouts, accuracy is slightly worse due to the larger number of objects and overlaps; the increased number of polygons along the line increases intersection ambiguity and error rates.
-G.8 Max Box
+I.8 Max Box
 Task Definition
@@
 Compute free space $F=R\setminus O$.
-Search over orientations $\theta\in[0,\pi)$: rotate $F$ by $\theta$, find the largest axis-aligned empty rectangle inside the rotated $F$, record $(w_{\theta},h_{\theta})$ and area $A_{\theta}=w_{\theta}h_{\theta}$, then map back to get $(w^{*},h^{*},\theta^{*})$ with $A^{*}=\max_{\theta}A_{\theta}$.
+Search over orientations $\theta\in[0,\pi)$: rotate $F$ by $\theta$, find the largest axis-aligned empty rectangle inside the rotated $F$, record $(w_{\theta},h_{\theta})$ and area $A_{\theta}=w_{\theta}h_{\theta}$, then map back to get $(w^{*},h^{*},\theta^{*})$ with $A_{max}=\max_{\theta}A_{\theta}$.
 Main Issue
@@
 Models often (i) ignore rotation and return an axis-aligned box, or (ii) mis-handle overlaps in free space, leading to under- or over-estimated maxima.
-Appendix H Prompts
+Appendix J Prompts
 This section illustrates the design of prompt templates used in our benchmark.
@@
 Figure 24: Prompt for populating a bedroom layout with functionally and spatially valid object placements, following layout-specific design rules.
-Appendix I Supplementary Accuracy Analyses
+Appendix K Supplementary Accuracy Analyses
 Table 12: % token-limit stop reason for general models.
-Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
+Question Room [Uncaptioned image] Claude
+Sonnet-4
+[Uncaptioned image] GPT-4.1
+[Uncaptioned image] Kimi-K2
+Instruct
+[Uncaptioned image] Qwen3
+Coder-480B
+[Uncaptioned image] Qwen3
+235B
+[Uncaptioned image] GPT-4.1
+mini
+[Uncaptioned image] Qwen3
+30B
+[Uncaptioned image] Devstral
+Small
 Pair distance K 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.2
@@
 Table 13: Question-level accuracy on completed answers for general models.
-Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
+Question Room [Uncaptioned image] Claude
+Sonnet-4
+[Uncaptioned image] GPT-4.1
+[Uncaptioned image] Kimi-K2
+Instruct
+[Uncaptioned image] Qwen3
+Coder-480B
+[Uncaptioned image] Qwen3
+235B
+[Uncaptioned image] GPT-4.1
+mini
+[Uncaptioned image] Qwen3
+30B
+[Uncaptioned image] Devstral
+Small
 Pair distance K 99.8 96.5 95.7 96.8 99.2 90.7 89.5 58.9
@@
 Reposi- tioning K 73.8 63.9 48.7 45.3 88.5 66.3 75.4 14.9
-LR 79.3 60.4 56.7 64.5 92.9 76.8 72.9 25.3
-B 71.0 55.5 48.3 59.5 85.4 72.6 70.6 21.5
+LR 79.3 60.4 56.7 64.5 92.9 76.8 72.9 25.2
+B 71.0 55.5 48.3 59.5 85.4 72.6 70.6 21.4
 HSSD 42.0 47.0 28.0 34.0 73.6 41.3 46.5 10.5
 Free space K 97.8 93.2 83.1 84.2 95.2 95.2 84.1 67.9
-LR 0.2 14.2 2.8 1.8 3.7 1.2 0.5 3.0
-B 2.7 31.2 1.0 1.3 9.3 0.8 1.0 0.7
+LR 0.2 14.2 2.8 1.8 3.6 1.2 0.5 3.0
+B 2.7 31.2 1.0 1.3 9.2 0.8 1.0 0.7
 HSSD 35.0 16.3 24.0 22.0 35.8 15.7 12.3 9.1
-Visibility K 63.3 87.7 54.5 67.0 98.3 90.2 91.6 18.6
+Visibility K 63.3 87.7 54.5 67.0 98.3 90.2 91.7 18.6
 LR 52.8 81.7 43.3 52.2 98.5 87.0 89.1 10.0
@@
 LR 7.8 7.0 5.1 5.7 44.5 4.5 12.5 0.5
-B 5.8 6.8 7.4 5.0 54.5 4.9 9.0 1.7
+B 5.8 6.8 7.4 5.0 54.5 4.8 9.0 1.7
 HSSD 5.0 7.5 2.5 2.0 21.1 4.5 3.8 1.0
 Shortest path (valid) K 59.2 61.7 52.7 39.7 81.1 55.3 37.3 36.6
-LR 53.0 51.3 42.8 37.3 72.1 47.3 37.9 32.9
+LR 53.0 51.3 42.8 37.3 72.1 47.2 37.9 32.9
 B 48.6 52.2 40.7 34.7 73.8 45.8 36.3 36.9
@@
 Table 14: % token-limit stop reason for reasoning models.
-Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
+Question Room [Uncaptioned image] GPT-5
+[Uncaptioned image] gpt-oss
+120b
+[Uncaptioned image] DeepSeek
+R1-0528
+[Uncaptioned image] Gemini
+Flash 2.5
+[Uncaptioned image] GPT-5
+mini-2025
+[Uncaptioned image] gpt-oss
+20b
+[Uncaptioned image] Qwen3
+30B Think.
 Pair distance K 0.0 0.0 1.5 3.3 0.0 0.7 2.0
@@
 Table 15: Question-level accuracy on completed answers for reasoning models.
-Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
-Pair distance K 99.8 99.3 98.0 96.3 100.0 94.2 97.7
-LR 98.8 99.3 99.0 96.0 99.7 93.5 97.5
-B 98.3 99.5 96.8 95.5 99.7 93.8 95.3
-HSSD 69.0 78.5 25.5 12.5 32.5 40.5 18.0
-Placement K 84.7 92.0 89.0 59.7 90.8 85.7 68.2
-LR 75.5 89.0 82.2 53.3 86.3 78.5 46.5
-B 61.2 83.5 72.0 35.8 81.2 62.0 34.5
-HSSD 70.0 85.0 79.0 16.0 75.5 74.5 28.5
-Reposi- tioning K 83.0 85.5 79.2 90.5 84.5 70.8 61.5
-LR 85.5 89.8 86.2 91.2 92.8 87.8 77.0
-B 77.8 83.3 83.3 85.2 84.3 78.0 69.2
-HSSD 49.5 60.5 47.5 18.5 53.5 40.0 27.0
-Free space K 82.5 99.0 93.0 93.3 99.5 94.8 97.0
-LR 47.0 83.3 18.3 17.7 78.5 53.2 0.0
-B 50.5 87.5 34.8 33.3 82.2 74.0 1.0
-HSSD 19.5 31.0 6.5 1.0 5.0 9.0 1.0
-Visibility K 94.8 94.2 71.3 26.8 98.0 91.5 78.8
-LR 95.2 94.0 52.0 11.3 98.0 89.3 70.8
-B 94.2 92.5 53.5 11.2 95.5 89.2 64.2
-HSSD 57.0 70.0 10.0 0.5 39.0 45.5 3.0
-View Angle K 96.2 98.5 73.7 92.5 88.3 93.5 98.5
-LR 93.3 97.3 68.2 91.5 84.5 92.0 98.3
-B 95.2 98.2 75.2 93.8 86.8 91.3 98.3
-HSSD 59.5 74.0 13.5 20.0 25.5 37.5 26.0
-Max Box K 48.5 62.8 50.5 3.7 85.2 31.3 16.7
-LR 17.3 28.2 8.0 0.0 60.3 3.8 0.8
-B 13.0 30.3 11.0 0.0 61.2 5.8 0.5
-HSSD 5.0 9.5 2.5 0.0 17.0 0.5 0.0
-Shortest path (valid) K 64.7 64.2 12.3 3.2 52.3 43.0 14.2
-LR 21.2 66.0 12.5 1.5 53.7 37.7 16.7
-B 58.2 57.8 7.8 1.0 52.0 30.0 14.3
-HSSD 28.5 33.5 8.5 0.0 16.5 13.5 1.0
-Shortest path (Fréchet) K 56.3 55.5 11.5 3.2 50.5 42.2 14.0
-LR 12.3 40.5 9.3 1.3 47.5 28.5 16.3
-B 39.3 44.8 6.0 0.8 47.7 24.2 14.3
-HSSD 22.5 26.5 2.5 0.0 12.0 14.0 1.0
+Question Room [Uncaptioned image] GPT-5
+[Uncaptioned image] gpt-oss
+120b
+[Uncaptioned image] DeepSeek
+R1-0528
+[Uncaptioned image] Gemini
+Flash 2.5
+[Uncaptioned image] GPT-5
+mini-2025
+[Uncaptioned image] gpt-oss
+20b
+[Uncaptioned image] Qwen3
+30B Think.
+Pair distance K 99.8 98.0 99.5 99.7 100.0 94.8 99.7
+LR 98.8 99.3 99.8 100.0 99.8 94.1 99.8
+B 98.3 99.5 99.3 99.8 99.8 94.5 100.0
+HSSD 69.0 88.2 85.0 92.6 98.5 66.9 64.3
+Placement K 97.5 92.0 93.8 98.1 97.2 89.7 96.7
+LR 97.6 89.2 88.5 98.5 96.3 89.5 97.9
+B 95.8 83.6 78.4 97.7 93.1 83.4 97.2
+HSSD 95.9 85.0 84.0 100.0 90.4 87.7 96.6
+Reposi- tioning K 83.3 85.5 86.4 91.4 84.5 71.4 66.2
+LR 85.5 89.8 93.2 93.2 92.8 88.3 82.5
+B 78.0 83.3 86.1 86.6 84.3 78.7 75.5
+HSSD 50.5 60.8 71.4 77.1 74.8 51.6 73.0
+Free space K 83.5 99.0 98.4 95.2 99.5 95.6 99.3
+LR 47.1 83.5 33.0 30.3 79.8 61.6 0.0
+B 50.6 87.5 39.7 48.1 83.0 80.3 1.2
+HSSD 20.5 43.4 31.7 28.6 33.3 40.0 66.7
+Visibility K 94.8 95.6 97.3 98.8 98.0 91.8 97.5
+LR 95.2 95.6 97.8 95.8 98.3 90.2 96.2
+B 94.2 93.6 96.4 95.7 97.0 89.3 96.2
+HSSD 57.0 73.7 76.9 100.0 89.7 64.1 75.0
+View angle K 96.2 98.5 98.9 97.9 99.4 94.4 99.5
+LR 93.3 97.3 98.5 96.3 98.6 93.2 99.0
+B 95.2 98.2 98.7 97.4 99.4 92.4 99.2
+HSSD 59.5 80.0 87.1 81.6 94.4 60.5 76.5
+Max box K 48.5 62.8 72.3 84.6 86.9 44.0 44.6
+LR 17.3 28.2 21.9 0.0 64.0 9.8 41.7
+B 13.0 30.3 27.1 0.0 64.2 13.3 27.3
+HSSD 5.0 9.5 6.0 0.0 21.9 0.8 0.0
+Shortest path (valid) K 64.8 64.5 57.4 82.6 68.4 67.4 95.5
+LR 55.2 66.8 56.4 69.2 66.9 63.1 89.3
+B 58.5 58.1 38.2 50.0 63.0 60.4 87.8
+HSSD 28.5 34.5 47.2 0.0 47.8 20.0 50.0
+Shortest path (Fréchet) K 56.4 55.8 53.5 82.6 66.0 66.1 94.4
+LR 32.2 41.0 42.1 61.5 59.2 47.8 87.5
+B 39.5 45.1 29.3 41.7 57.8 48.7 87.8
+HSSD 22.5 27.3 13.9 0.0 34.8 20.7 50.0